### v1

# Hanke-Raus heuristic rule for iteratively regularized stochastic gradient descent

Abstract. In this work, we present a novel variant of the stochastic gradient descent method termed as iteratively regularized stochastic gradient descent (IRSGD) method to solve nonlinear ill-posed problems in Hilbert spaces. Under standard assumptions, we demonstrate that the mean square iteration error of the method converges to zero for exact data. In the presence of noisy data, we first propose a heuristic parameter choice rule (HPCR) based on the method suggested by Hanke and Raus, and then apply the IRSGD method in combination with HPCR. Precisely, HPCR selects the regularization parameter without requiring any a-priori knowledge of the noise level. We show that the method terminates in finitely many steps in case of noisy data and has regularizing features. Further, we discuss the convergence rates of the method using well-known source and other related conditions under HPCR as well as discrepancy principle. To the best of our knowledge, this is the first work that establishes both the regularization properties and convergence rates of a stochastic gradient method using a heuristic type rule in the setting of infinite-dimensional Hilbert spaces. Finally, we provide the numerical experiments to showcase the practical efficacy of the proposed method.
Keywords: Stochastic gradient descent; Iterative Landweber method; Convergence rates; Variational inequalities; Inverse problems; Nonlinear ill-posed problems; Regularization.
1. Introduction
This work focuses on deriving an approximate solution for a system of nonlinear ill-posed equations of the form
((1.1)) $F_{i}(u)=y^{\dagger}_{i},\hskip 5.69054pti=0,1,\ldots,P-1,$
[…]
((1.2)) $F(u)=y^{\dagger},$
where $F:U\rightarrow Y^{P}$ is defined as
$F(u)=\begin{pmatrix}F_{0}(u)\\ \vdots\\ F_{P-1}(u)\end{pmatrix}\ \ \text{and}\ \hskip 5.69054pty^{\dagger}=\begin{pmatrix}y^{\dagger}_{0}\\ \vdots\\ y^{\dagger}_{P-1}\end{pmatrix}.$
Usually, exact data $y^{\dagger}$ is never precisely available. Instead, a noisy data $y^{\delta}\in Y^{P}$ is available such that
((1.3)) $\|y^{\dagger}-y^{\delta}\|\leq\delta,$
for noise level $\delta\geq 0$. Here, we write $\delta=\sqrt{\delta_{0}^{2}+\delta_{1}^{2}+\cdots+\delta_{P-1}^{2}}$ and $y^{\delta}=(y^{\delta}_{0},y^{\delta}_{1},...,y^{\delta}_{P-1})$, where
((1.4)) $\|y_{i}^{\delta}-y_{i}^{\dagger}\|\leq\delta_{i}\ \ \text{for}\ \ i=0,1,2,\cdots,P-1.$
Due to ill-posedness of (1.1), its solution may not exist and even if it exists, it may not be unique. Moreover, the solution(s) can be highly sensitive to perturbations in the noisy data $y^{\delta}$ (see [6]). Consequently, regularization methods are needed for deducing the stable approximate solution of (1.1).
One of the prominent and highly effective classes of regularization methods is that of iterative regularization. This approach has achieved significant success in solving a wide variety of inverse problems (see [6, 16] and the references therein). Among these methods, the Landweber Iteration Method (LIM) is a well-studied and widely recognized classical iterative technique. For a Fréchet differentiable (or smooth) forward mapping $F:U\rightarrow Y$, LIM can be written as
((1.5)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta}),\hskip 28.45274ptk\geq 0.$
Here $u_{0}^{\delta}:=u_{0}$represents an initial guess that incorporates the prior knowledge about the solution to be reconstructed and $F^{\prime}(u_{k}^{\delta})^{*}$ is the adjoint of the Fréchet derivative of $F$ at $u_{k}^{\delta}$.
Scherzer [26] introduced a modification to the method (1.5) by introducing a damping term and named it as iteratively regularized Landweber iteration method. This method can be written as
((1.6)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta})-\lambda_{k}(u_{k}^{\delta}-u^{(0)}),\hskip 28.45274ptk\geq 0,$
where $\lambda_{k}\geq 0$. In comparison to the method (1.5), the convergence rates analysis of the method (1.6) needs fewer assumptions on the mapping $F$. Furthermore, the method (1.6) converges to a solution which is closed to $u^{(0)}$, however, the method (1.5) needs additional assumptions for the similar convergence behaviour (see [16] for nice details). This means that the inclusion of the additional damping term in (1.6) has several advantages over the method (1.5). When the noisy data is available, a proper stopping rule (a-posteriori) must be employed in order to show that the iterative schemes (1.5) and (1.6) are a regularization one. The commonly incorporated stopping rule is well known as discrepancy principle, i.e., the method is halted after $k_{\delta}=k_{\delta}(\delta,y^{\delta})$ steps, where
((1.7)) $\|F(u_{k_{\delta}}^{\delta})-y^{\delta}\|\leq\tau\delta<\|F(u_{k}^{\delta})-y^{\delta}\|,\hskip 28.45274pt0\leq k<k_{\delta},$
for some $\tau>1.$ In real-world applications, noise level is not always available or trustworthy, which makes the discrepancy principle (1.7) difficult to use. Consequently, while applying the discrepancy principle, overestimation or underestimation of the noise level may result in significant degradation of the reconstruction accuracy. Therefore, it is reasonable to work on heuristic stopping rules that do not rely on the knowledge of noise level. Real [25] proposed a heuristic rule for the Landweber iteration (1.5) in Banach space setting based on the discrepancy principle, by taking motivation from the method of Hanke and Raus [9]. For (1.5), heuristic rule determines an integer $k_{*}:=k_{*}(y^{\delta})$ by minimizing
((1.8)) $\Psi(k,y^{\delta}):=(k+a)\|F({u_{k}^{\delta}})-y^{\delta}\|^{p}$
over $k$, where $a\geq 1$ is a fixed number and $p>0$ is a constant.
It can be observed that for the conventional iterative regularization methods for solving inverse problems (such as the Landweber method [16], Levenberg-Marquardt method [5], conjugate gradient method [16], non-stationary Tikhonov iterative regularization method [20] and iteratively regularized Gauss-Newton method [16], among others), one of the common challenges is the significant computational burden at each iteration, mainly due to the need to process all the data, which can be excessively large and resource-intensive. To address this issue, a potential and effective approach is the utilization of stochastic gradient descent (SGD), a method introduced by Robbins et al. [24]. Further, Jin et al. [13] studied the convergence analysis of SGD method and demonstrated its regularizing nature when there is a noise in the data. The basic version of SGD can be written as (see [14])
((1.9)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\mu_{k}F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}),\hskip 28.45274ptk\geq 0,$
where $\mu_{k}$ is the corresponding step size and $i_{k}$ is the uniformly drawn index from the index set $\{0,1,...,P-1\}.$ It can be noted that the method (1.9) is a randomized version of (1.5). For (1.9), the stopping index $k(\delta)\in\mathbb{N}$ is defined using the following a-priori stopping rule
((1.10)) $\lim_{\delta\rightarrow 0^{+}}k(\delta)=\infty\hskip 14.22636pt\text{and}\hskip 14.22636pt\lim_{\delta\rightarrow 0^{+}}\delta^{2}\sum_{i=1}^{k(\delta)}\mu_{i}=0.$
We refer to [14, 17] for some recent literature on SGD method using a-priori stopping rule for linear as well nonlinear inverse problems and [12, 27] on SGD method using a-posteriori stopping rule for linear inverse problems. It is evident from these references that they exclusively use an a-priori or a-posteriori stopping rule in developing their convergence theories. From this observation, it is of significant interest to develop heuristic rules for stochastic gradient descent (SGD) type methods. Therefore, the primary objective of this paper is to explore the convergence analysis of the following iteratively regularized stochastic gradient descent (IRSGD) under heuristic stopping rules:
((1.11)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}),\hskip 14.22636ptk\geq 0,$
$\hskip 85.35826pt\text{with}\quad 0\leq\lambda_{k}\leq\lambda_{\text{max}}<\frac{1}{2},$
where $\omega_{k}$ is the step size, $\lambda_{k}$ is the weighted parameter, $F^{\prime}_{i}(u)$ is the Fréchet derivative of $F_{i}(u)$, $i_{k}$ is an index randomly drawn from the index set $\{0,1,\ldots,P-1\}$ through uniform distributation. This method differs from the SGD method in the sense that an extra damping term $-\lambda_{k}(u_{k}^{\delta}-u_{0})$ is considered in its formulation. A key advantage of studying IRSGD lies in its ability to reflect SGD by setting the damping factor $\lambda_{k}=0$. Our approach is novel in the realm of stochastic methods for solving ill-posed inverse problems as it provides convergence analysis without relying on an a-priori stopping rule. Additionally, it circumvents the need to know the noise level, which further enhances its practical applicability.
For SGD method, Jin et al. [14] established the convergence rate in the mean squared norm in the context of nonlinear ill-posed problems, assuming an appropriate source condition. Further, Jahn et al. [12] further extended these results using an a-posteriori stopping rule, however, precise estimates for the convergence rates are not provided in that work. Inspired by these findings, one of the main aims of this work is to develop convergence rates for the iterates generated by (1.11) under an a-posteriori stopping rule. In addition to that, we also derive convergence rates under heuristic stopping criteria. To the best of our knowledge, this is the first study that addresses convergence rates of a stochastic type method for solving nonlinear ill-posed problems under heuristic stopping rule.
The main contributions of this paper are outlined as follows:
  • Instead of focusing on the analysis of a single equation, we turn our attention to exploring stochastic methods for solving systems comprising a large number of equations (i.e., $P$ is large in (1.1)). Stochastic methods involve the random consideration of each equation within equation (1.1) independently, thereby significantly reducing memory requirements. This approach finds relevance in numerous real-world scenarios (see [7, 21, 22]).
  • The method is fully data-driven, as it is applicable without using a-priori or a-posteriori stopping rule. This is the first attempt of construction of a stochastic method using heuristic type stopping rule. We establish the (strong) convergence of IRSGD method, which improves the weak convergence (i.e., convergence in probability) results provided in [12].
  • We establish the convergence rates for the proposed scheme (1.11) using a step size $\omega_{k}\in[\omega,\Omega]$ and a decaying weight parameter, under two types of stopping criteria: the discrepancy principle and a heuristic rule. The rates are comparable to those provided in [14, Theorem 2.5], [8, Theorem 3.2], and [13, Theorem 2.2] (see Table 1).
  • This work presents experimental evidence showcasing the method’s enhanced efficiency and accelerated performance achieved through the inclusion of a damping factor in SGD. This modification yields significantly better results compared to the standard SGD approach discussed in [14].
The rest of the paper is structured as follows. In Section 2, we list the preliminaries needed in our work. Section 3 covers the convergence analysis of our method in the presence of noisy and exact data. Section 4 discusses the convergence rates of our method for a-posteriori and heuristic stopping rules under certain assumptions. Section 5 presents numerical experiments on a system of ill-posed problems, demonstrating the applicability of the proposed scheme with heuristic stopping rule. The final section offers concluding remarks. Additionally, the Appendix includes few auxiliary lemmas required for our analysis.
2. Preliminaries
In this section, we review certain basic definitions and notation related to our work, see [2, 4, 14, 15, 23] for more details.
The notation $u_{k}$ is used to denote the iterates of our method corresponding to the exact data $y^{\dagger}.$ Additionally, we refer $\mathcal{F}_{k}$ as the filtration generated by the random indices $\{i_{1},...,i_{k-1}\}$ up to the $(k-1)$th iteration. Because we randomly select the indices $i_{k}$, the iteration $u_{k}^{\delta}$ in the IRSGD method becomes a random process. To assess its convergence, multiple methods are available. In our case, we will use the mean squared norm defined by $\mathbb{E}[\|\cdot\|^{2}]$, where $\mathbb{E}[\cdot]$ represents the expectation taken with respect to the filtration $\mathcal{F}_{k}$. It is important to note that the iterates $u_{k}^{\delta}$ are measurable with respect to $\mathcal{F}_{k}$ (see [15] for more details).
The following well-known identities will be frequently utilized: for any random variables $\Theta,\Theta_{1},\Theta_{2}$ and $\Phi$, and real constants $c_{1},c_{2}$, we have
  • (a)
    $\mathbb{E}[\Theta]=\mathbb{E}[\mathbb{E}[\Theta|\mathcal{F}_{k}]].$
  • (b)
    Additive: $\mathbb{E}[\Theta_{1}+\Theta_{2}|\mathcal{F}_{k}]=\mathbb{E}[\Theta_{1}|\mathcal{F}_{k}]+\mathbb{E}[\Theta_{2}|\mathcal{F}_{k}].$
    Linearity: $\mathbb{E}[c_{1}\Theta+c_{2}|\mathcal{F}_{k}]=c_{1}\mathbb{E}[\Theta|\mathcal{F}_{k}]+c_{2}.$
  • (c)
    If $\Theta\leq\Phi$, then $\mathbb{E}[\Theta|\mathcal{F}_{k}]\leq\mathbb{E}[\Phi|\mathcal{F}_{k}]$ almost surely.
Here the expression $\mathbb{E}[\Theta|\mathcal{F}_{k}]$ represents the expected value of $\Theta$ given the information contained in $\mathcal{F}_{k}$.
Triangular inequality for expectation: Let $\Theta_{1}$ and $\Theta_{2}$ be random variables. Then, there holds
((2.1)) $\mathbb{E}[\|\Theta_{1}+\Theta_{2}\|]\leq\mathbb{E}[\|\Theta_{1}\|]+\mathbb{E}[\|\Theta_{2}\|].$
Jensen’s inequality: For any random variable $\Theta$ and for any convex function $f$, there holds
((2.2)) $f(\mathbb{E}[\|\Theta\|])\leq\mathbb{E}[f(\|\Theta\|)],$
provided $\mathbb{E}[\|\Theta\|]<\infty$ and $\mathbb{E}[f(\|\Theta\|)]<\infty$ (see [1] for more details).
Next, we introduce an important result that splits the mean squared error $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]$, where $e_{k}^{\delta}:=u_{k}^{\delta}-u^{\dagger}$ represents the error for any iteration $u_{k}^{\delta}$. This decomposition is a fundamental strategy in error analysis and plays a crucial role in deriving the convergence rates for iterative methods.
Bias-variance decomposition: The bias-variance decomposition for the iterative scheme is given by
((2.3)) $\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]=\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|^{2}+\mathbb{E}[\|B^{h}(e_{k}^{\delta}-\mathbb{E}[e_{k}^{\delta}])\|^{2}],$
where $B$ is defined as in (4.1) and $h\geq 0.$ In this decomposition:
  • The first term, $\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|^{2}$, represents the bias, which includes both the approximation error and the error due to noise in the data.
  • The second term, $\mathbb{E}[\|B^{h}(e_{k}^{\delta}-\mathbb{E}[e_{k}^{\delta}])\|^{2}]$, represents the variance, which accounts for the randomness introduced by the selection of the index $i_{k}$ during the iterations.
The bias reflects the deterministic error from the method’s approximation, while the variance captures the stochastic error due to the randomized nature of the iterative process.
We end this section by discussing few notations.
For some $u\in U$ and $\rho>0,$ we denote the open and closed balls in $U$ of radius $\rho$ centered at $u$, respectively by $B_{\rho}(u)$ and $\overline{B}_{\rho}(u).$ Let $\mathcal{D}_{\rho}(u^{\dagger})$ denote the set of all the solutions of (1.2) in $\overline{B}_{\rho}(u^{\dagger})$, i.e., $\mathcal{D}_{\rho}(u^{\dagger}):=\{u\in\overline{B}_{\rho}(u^{\dagger}):F(u)=y^{\dagger}\}$.
3. Iteratively regularized Stochastic gradient descent in Hilbert spaces
This section covers the convergence analysis of IRSGD method under certain assumptions. Precisely, this section is organised as follows. The subsection 3.1 aims to discuss the assumptions required for studying the convergence analysis. In subsection 3.2, we discuss a heuristic rule based on the modified discrepancy principle. Convergence of non-noisy iterates is discussed in subsection 3.3. Lastly, subsection 3.4 aims to demonstrate the convergence analysis of noisy iterates.
We recall that the IRSGD method can be expressed as
((3.1)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}),\hskip 14.22636ptk\geq 0,$
with $\text{with}\ 0\leq\lambda_{k}\leq\lambda_{\text{max}}<\frac{1}{2}$, where $F^{\prime}_{i_{k}}$ is the Frećhet derivative of $F_{i_{k}}$, $u_{0}^{\delta}$ is the initial guess of the exact solution, $\omega_{k}$ is the step size and the index $i_{k}$ is drawn randomly from the index set $\{0,1,\ldots,P-1\}$ at each iteration step.
3.1. Assumptions
In our work, we need the following assumptions.
Assumption 3.1.
For $i\in\{0,1,\ldots,P-1\}$ and $\rho>0$, there holds:
  • (i)
    The operator $F:U\rightarrow Y^{P}$ is continuous.
  • (ii)
    There exists a constant $L>0$ such that
    ((3.2)) $\|F_{i}^{\prime}(u)\|\leq L\ \text{for all}\ u\in B_{\rho+c(\rho)}(u_{0}),$
    where $L\leq\frac{1}{2\omega}$ (see Lemma 3.1 for the meaning of $c(\rho)$ and $\omega$). Moreover there is a constant $10<\eta<$ 1 such that
    ((3.3)) $\|F_{i}(\overline{u})-F_{i}(u)-F_{i}^{\prime}(u)(\overline{u}-u)\|\leq\eta\|F_{i}(\overline{u})-F_{i}(u)\|$
    for all $\overline{u},u\in B_{\rho+c(\rho)}(u_{0}).$
  • (iii)
    (1.2) has a solution $\hat{u}\in B_{\rho}(u_{0}).$
Motivated by [9], we consider the following heuristic rule for IRSGD in Hilbert spaces.
Rule 1.
Let $a\geq 1$ be a fixed number and let
$\Psi(k,y^{\delta}):=(k+a)\|F(u^{\delta}_{k})-y^{\delta}\|^{2}.$
We define $k_{*}:=k_{*}(y^{\delta})$ to be an integer such that
$k_{*}\in\arg\min\{\Psi(k,y^{\delta}):0\leq k\leq k_{\infty}\},$
where $k_{\infty}:=k_{\infty}(y^{\delta})$ is the largest integer such that $u^{\delta}_{k}\in\mathfrak{D}(F)$ for all $0\leq k<k_{\infty}$.
Remark 3.1.
Rule 1 is straightforward to implement in practice, as detailed in Section 5. During the iteration process, the value of $\Psi(k,y^{\delta})$ is monitored as a function of $k$. After a sufficient number of iterations, the process is terminated and $k_{*}$ is selected that minimizes $\Psi(k,y^{\delta})$. However, an important challenge associated with the nonlinear Landweber iteration [11] lies in determining an appropriate upper bound on the number of iterations. This is particularly critical due to its local convergence properties, which can lead to premature stopping or inaccurate identification of $k_{*}$. Therefore, the limit on the upper bound must be sufficiently large to avoid inaccurately estimating $k_{*}$ at the first local minimum, which is often not the global minimum [11]. To prevent $k_{*}$ from becoming too small, we have performed several numerical experiments on different values of $a$ in Rule 1.
Note that in order to prove Lemma 3.1 for IRSGD, it is essential to assume that $a\geq 1$. This Lemma will be further required to show the convergence of the scheme. Using the integer $k_{*}:=k_{*}(y^{\delta})$ determined by Rule 1, we call $u^{\delta}_{k_{*}(y^{\delta})}$ as an approximate solution.
A natural question arises concerning a family of noisy data $\{y^{\delta}\}$, where $y^{\delta}\to y^{\dagger}$ as $\delta\to 0$ is that how to guarantee that $u^{\delta}_{k_{*}(y^{\delta})}$ converges to a solution of (1.2) as $\delta\to 0\mathord{?}$
In general, the answer is no, as stated by Bakushinskii’s veto [3], which asserts that heuristic rules cannot guarantee the convergence in the worst-case scenario for any regularization method. Later, in [9], it is shown that the following assumption guarantee the convergence. We also need this condition in our work.
Assumption 3.2 (Noise condition).
The family of noisy data $\{y^{\delta}\}$ satisfies the condition $0<\|y^{\delta}-y^{\dagger}\|\rightarrow 0$ as $\delta\rightarrow 0$. Moreover, there exists a constant $\varkappa>0$ such that for every $y^{\delta}$ and for each $u\in S(y^{\delta})$, where $S(y^{\delta}):=\{u^{\delta}_{k}:0\leq k\leq k_{\infty}\}$ is constructed using (3.1), there holds
$\|y^{\delta}-F(u)\|\geq\varkappa\|y^{\delta}-y^{\dagger}\|.$
Under Assumption 3.2, it is immediate to see that Rule 1 comes out with a finite integer $k_{*}$. Indeed, if $k_{\infty}$ is finite, then there is nothing to prove. So we only consider that $k_{\infty}=\infty$. It then follows from Assumption 3.2 that
$\Psi(k,y^{\delta})=(k+a)\|F(u^{\delta}_{k})-y^{\delta}\|^{2}\geq(k+a)\varkappa\|y^{\dagger}-y^{\delta}\|^{2}\rightarrow\infty$
as $k\rightarrow\infty$. Consequently, we can conclude that a finite integer $k_{*}$ exists, which achieves the minimum of $\Psi(k,y^{\delta})$.
3.2. Analysis for Rule 1
This subsection aims to demonstrate that
((3.4)) $\Psi(k_{*}(y^{\delta}),y^{\delta})\rightarrow 0\quad\text{as}\quad\delta\rightarrow 0.$
Additionally, we demonstrate that IRSGD iterates $u_{k}^{\delta}$ remains in a ball around $u_{0}$ under certain conditions.
In order to show (3.4), we propose an auxiliary integer $\hat{k}_{\delta}$ defined by the stopping rule
((3.5)) $\|F(u^{\delta}_{\hat{k}_{\delta}})-y^{\delta}\|+\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y^{\delta}-y^{\dagger}\|<\|F(u^{\delta}_{k})-y^{\delta}\|+\frac{M}{k+a}\quad$
for $0\leq k<\hat{k}_{\delta}$, where $M=\frac{\varrho c(\rho)^{2}\tau}{\Omega(1+\varrho)(1+\eta)},\varrho>0$ is a constant, $\Omega>0,\eta$ is same as in (3.3), $c(\rho)$ is defined in (3.6) and $\tau>1$ is a real number (not to be confused with the one in (1.7)). Adding an extra term in the classical discrepancy principle produces a minor variation of the discrepancy principle (1.7). An essential component in establishing (3.4) is provided by this additional term, which ensures that $\hat{k}_{\delta}\rightarrow\infty$ as $\delta\rightarrow 0$.
Lemma 3.1.
Let Assumption 3.1 hold and let $\Omega,\omega$ be positive constants such that $\Omega\geq\omega$. Moreover, let $\kappa\in(0,1)$ and $\varrho>0$ be a sufficiently small constant and
((3.6)) $c(\rho):=\rho\lambda_{\textbf{max}}\frac{(1-\lambda_{\textbf{max}})+\sqrt{(1-\lambda_{\textbf{max}})^{2}+(\frac{1}{1+\varrho}-(1-\lambda_{\text{max}})^{2})(1+\kappa^{-2}\Omega^{2}L^{2})}}{(\frac{1}{1+\varrho}-(1-\lambda_{\text{max}})^{2})}.$
Then, under the following conditions
((3.7)) $D:=2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2}-\tau^{-1}\Omega(2+M)(1+\eta)>0,$
((3.8)) $\sum_{k=0}^{\infty}\lambda_{k}\leq\lambda<\infty,$
there holds
  • (a)
    the stopping index $\hat{k}_{\delta}$ defined in (3.5) is finite for $\delta>0$ and $\hat{k}_{\delta}\rightarrow\infty$ as $\delta\rightarrow 0.$
  • (b)
    for a solution $\hat{u}$ to (1.2) and for any $0\leq k<\hat{k}_{\delta}$, if $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$, then $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})\subset B_{\rho+c(\rho)}(u_{0}).$
Proof.
First we prove assertion (b). For $k=0$, result trivially follows from (iii) of Assumption 3.1. By assuming that $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<\hat{k}_{\delta}$, we demonstrate that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u}).$ From (3.1), we have
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $=\|u_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0})-\hat{u}\|^{2}$
$=(1-\lambda_{k})^{2}\|u^{\delta}_{k}-\hat{u}\|^{2}+\lambda_{k}^{2}\|u_{0}-\hat{u}\|^{2}+2\lambda_{k}(1-\lambda_{k})\langle u^{\delta}_{k}-\hat{u},u_{0}-\hat{u}\rangle$
((3.9)) $\hskip 14.22636pt+\omega_{k}^{2}\|F^{\prime}_{i_{k}}(u^{\delta}_{k})^{*}(y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}))\|^{2}+2\omega_{k}(1-\lambda_{k})\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F^{\prime}_{i_{k}}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle$
$\hskip 14.22636pt+2\omega_{k}\lambda_{k}\langle u_{0}-\hat{u},F^{\prime}_{i_{k}}(u^{\delta}_{k})^{*}(y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}))\rangle.$
To estimate 3.2, we can write the last two terms of the sum as
$\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F_{i_{k}}^{\prime}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle$ $=\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),y_{i_{k}}^{\delta}-y_{i_{k}}^{\dagger}\rangle-\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
((3.10)) $\hskip 5.69054pt-\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F_{i_{k}}(u^{\delta}_{k})-F_{i_{k}}(\hat{u})-F_{i_{k}}^{\prime}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle,$
$2\omega_{k}\lambda_{k}\langle u_{0}-\hat{u},F^{\prime}_{i_{k}}(u^{\delta}_{k})^{*}(y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}))\rangle$ $\leq 2\lambda_{k}\omega_{k}L\|u_{0}-\hat{u}\|\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|$
((3.11)) $\leq\kappa^{-2}\lambda_{k}^{2}\omega_{k}^{2}L^{2}\|u_{0}-\hat{u}\|^{2}+\kappa^{2}\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2},$
where we have used (3.2) and the inequality $2ab\leq a^{2}+b^{2}$ for $a,b\in\mathbb{R}$ in obtaining 3.2.
By incorporating 3.2, 3.2, (3.3) and (3.2) in 3.2, we attain that
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u^{\delta}_{k}-\hat{u}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\|u_{0}-\hat{u}\|^{2}$
$+2\lambda_{k}(1-\lambda_{k})\|u^{\delta}_{k}-\hat{u}\|\|u_{0}-\hat{u}\|$
((3.12)) $+\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|\big{(}2\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
$-\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\big{)}.$
Next, according to the definition of $\hat{k}_{\delta}$ in (3.5), for $k<\hat{k}_{\delta}$ we further derive that
((3.13)) $\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|\leq\frac{1}{\tau}\left(\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}+\frac{M}{k+a}\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|\right).$
Using the fact that $\omega_{k}\in[\omega,\Omega],u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ and $\hat{u}\in B_{\rho}(u_{0})$ along with inequality (3.13) in 3.2, we have
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq(1-\lambda_{k})^{2}c(\rho)^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\rho^{2}+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)$
((3.14)) $-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-2\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{2M\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|}{\tau(k+a)}.$
We note that
$(1+\varrho)\big{(}(1-\lambda_{k})^{2}c^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\rho^{2}+2\lambda_{k}(1-\lambda_{k})\rho c\big{)}\leq c^{2},$
provided
$c\bigg{(}\frac{1}{1+\varrho}-(1-\lambda_{k})^{2}\bigg{)}\geq\rho\lambda_{k}(1-\lambda_{k})+\sqrt{(1-\lambda_{k})^{2}+\bigg{(}\frac{1}{1+\varrho}-(1-\lambda_{k})^{2}\bigg{)}(1+\kappa^{-2}\Omega^{2}L^{2})},$
which is true for $c(\rho)$. From (3.6) along with the inequality $2ab\leq a^{2}+b^{2},\forall\hskip 5.69054pta,b\in\mathbb{R}$ and 3.2, we get
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-2\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau}\left(\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}+\frac{1}{(k+a)^{2}}\right).$
$\leq-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-(2+M)\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
As $\omega_{k}\in[\omega,\Omega],0\leq\lambda_{k}\leq\lambda_{\text{max}}<1$ and $1-\lambda_{k}\leq 1$, we further attain
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\Omega(1+\eta)}{\tau(k+a)^{2}}.$
Moreover, given the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$, we can write the last inequality as
$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}|\mathcal{F}_{k}]$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\|y^{\delta}-F(u^{\delta}_{k})\|^{2}$
$\hskip 28.45274pt+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\Omega(1+\eta)}{\tau(k+a)^{2}}.$
Finally, by taking into consideration the entire conditional, we arrive at
$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\mathbb{E}[\|y^{\delta}-F(u^{\delta}_{k})\|^{2}]$
$\hskip 28.45274pt+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\Omega(1+\eta)}{\tau(k+a)^{2}}.$
This with (3.7), (3.5) for $M=\frac{\varrho c(\rho)^{2}\tau}{\Omega(1+\varrho)(1+\eta)}$ and the inequality $k+a\geq 1$ implies that
$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]\leq\frac{c(\rho)^{2}}{1+\varrho}+\frac{\varrho c(\rho)^{2}}{1+\varrho}=c(\rho)^{2}.$
By incorporating Jensen’s inequality (2.2) with $f(x)=x^{2}$, we conclude that
$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|]\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]^{1/2}\leq c(\rho).$
This means that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})$. Further, by engaging the triangular inequality (2.1) and the fact that $\|u_{0}-\hat{u}\|\leq\rho$, we conclude that
$\mathbb{E}[\|u^{\delta}_{k+1}-u_{0}\|]$ $\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|]+\mathbb{E}[\|\hat{u}-u_{0}\|]\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]^{1/2}+\|\hat{u}-u_{0}\|\leq\rho+c(\rho),$
which implies that $u_{k+1}^{\delta}\in B_{\rho+c(\rho)}(u_{0})$.
Next, we will prove that $\hat{k}_{\delta}$ is finite. To establish this, we can utilize 3.2 with (3.13) and (3.7) along with the fact that $1-\lambda_{k}\leq 1$ for $\lambda\in[0,\frac{1}{2})$ to get
$\|u_{k+1}^{\delta}-\hat{u}\|^{2}+D\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\rho^{2}$
$+2\lambda_{k}(1-\lambda_{k})\|u_{k}^{\delta}-\hat{u}\|\rho+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
((3.15)) $\leq\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\rho^{2}+2\lambda_{k}\rho c(\rho)+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.\hskip 56.9055pt$
Using the inequality that $\lambda_{k}^{2}\leq\lambda_{k}$ and the definition of the measurability of $u_{k}^{\delta}$ along with the additivity of expected values with respect to $\mathcal{F}_{k}$ in 3.2 to get that
((3.16)) $\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}|\mathcal{F}_{k}]+D\|y^{\delta}-F(u_{k}^{\delta})\|^{2}$ $\leq\mathbb{E}[\|u_{k}^{\delta}-\hat{u}\|^{2}|\mathcal{F}_{k}]+\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}$
$\hskip 56.9055pt+2\rho c(\rho)\lambda_{k}+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
Taking the sum of (3.16) from $k=0$ to $k=\hat{k}_{\delta}-1$ to deduce that
((3.17)) $\sum_{k=0}^{\hat{k}_{\delta}-1}\|y^{\delta}-F(u^{\delta}_{k})\|^{2}\leq D^{-1}\left(\rho^{2}+\left(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)\right)\sum_{k=0}^{\hat{k}_{\delta}-1}\lambda_{k}+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1-\lambda_{k}}{(k+a)^{2}}\right).$
From (3.5) and inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})\hskip 5.69054pt\forall\hskip 5.69054pta,b\in\mathbb{R}$, we get
((3.18)) $\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\leq\left(\|F(u_{k}^{\delta})-y^{\delta}\|+\frac{M}{k+a}\right)^{2}\leq 2\left(\|F(u_{k}^{\delta})-y^{\delta}\|^{2}+\frac{M^{2}}{(k+a)^{2}}\right).$
By taking the sum of (3.18) from $k=0$ to $\hat{k}_{\delta}-1$, we get
((3.19)) $\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}\leq 2\left(\sum_{k=0}^{\hat{k}_{\delta}-1}\|F(u_{k}^{\delta})-y^{\delta}\|^{2}+M^{2}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right).$
Using (3.17) in (3.19) to get
$\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}$ $\leq 2\left[D^{-1}\left\{\rho^{2}+\left(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)\right)\sum_{k=0}^{\hat{k}_{\delta}-1}\lambda_{k}\right.\right.$
$\left.\left.\hskip 85.35826pt+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right\}+M^{2}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right].$
By applying (3.8) and the result $\sum_{k=0}^{\infty}\frac{1}{(k+a)^{2}}\leq\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$ for $a\geq 1$, we further get
$\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}<\infty.$
Hence $\hat{k}_{\delta}$ is finite. Finally, from the definition of $\hat{k}_{\delta}$ in (3.5), we get
$\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y^{\delta}-y^{\dagger}\|\to 0\text{\quad as\quad}\delta\to 0.$
Therefore, we must have $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$. ∎
Remark 3.2.
For the definition of $c(\rho)$ to be meaningful, there must exist a positive real root of the quadratic equation
$\left((1-\lambda_{k})^{2}-\frac{1}{1+\varrho}\right)c^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\rho^{2}+2\lambda_{k}(1-\lambda_{k})\rho c=0,$
which is satisfied if the condition $\left((1-\lambda_{k})^{2}-\frac{1}{1+\varrho}\right)<0$ holds. Therefore, a sufficiently small positive constant $\varrho$ is required in (3.6), due to the restriction that $\lambda_{k}\in[0,\frac{1}{2})$.
Lemma 3.2.
Let Assumption 3.1 hold and let the sequence $\{\lambda_{k}\}$ be chosen such that (3.7) and (3.8) hold. Additionally, let $\{y^{\delta}\}$ represent a family of noisy data that fulfill Assumption 3.2. Furthermore, assume that $k_{*}:=k_{*}(y^{\delta})$ is the integer determined by Rule 1. Then
$\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0\quad\text{as}\quad\delta\to 0.$
Consequently, we have
$\|F(u^{\delta}_{k_{*}(y^{\delta})})-y^{\delta}\|\to 0\quad\text{and}\quad k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\to 0\quad\text{as}\quad\delta\to 0.$
Proof.
By employing the condition (3.8) and the fact that $\sum_{k=0}^{\infty}\frac{1}{(k+a)^{2}}\leq\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$ for $a\geq 1$ in (3.17), we arrive at
((3.20)) $\sum_{k=0}^{\hat{k}_{\delta}-1}\|y^{\delta}-F(u^{\delta}_{k})\|^{2}\leq D^{-1}C_{1},$
where $C_{1}=\left(\rho^{2}+\left(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)\right)\lambda+\frac{M\Omega\pi^{2}(1+\eta)}{6\tau}\right)$. By minimality of $\Psi(k_{*},y^{\delta})$, it follows from Rule 1 that
$\|F(u_{k}^{\delta})-y^{\delta}\|^{2}=\frac{\Psi(k,y^{\delta})}{k+a}\geq\frac{\Psi(k_{*}(y^{\delta}),y^{\delta})}{k+a}.$
This and (3.20) further yields
((3.21)) $\Psi(k_{*}(y^{\delta}),y^{\delta})\sum^{\hat{k}_{\delta}-1}_{k=0}\frac{1}{k+a}\leq\sum_{k=0}^{\hat{k}_{\delta}-1}\|F(u^{\delta}_{k})-y^{\delta}\|^{2}\leq D^{-1}C_{1}.$
Note that
[…]
Plugging the last estimate in (3.21) yields
$\bigg{(}\log\frac{\hat{k}_{\delta}+a}{a}\bigg{)}\Psi(k_{*}(y^{\delta}),y^{\delta})\leq D^{-1}C_{1}.$
Since $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$, we must have $\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0$ as $\delta\to 0$. Further, we note that
$a\|F(u^{\delta}_{k^{*}(y^{\delta})})-y^{\delta}\|^{2}\leq\Psi(k_{*}(y^{\delta}),y^{\delta}).$
Also, Assumption 3.2 implies that
$(k_{*}(y^{\delta})+a)\varkappa^{2}\|y^{\delta}-y^{\dagger}\|^{2}\leq\Psi(k_{*}(y^{\delta}),y^{\delta}).$
Hence, we conclude that $\|F(u_{k_{*}(y^{\delta})}^{\delta})-y^{\delta}\|\to 0$ and $k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\to 0$ as $\delta\to 0,$ since $\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0$ as $\delta\to 0$. Thus, result. ∎
The next result demonstrates that IRSGD iterates $u_{k}^{\delta}$ remains in a ball around $u_{0}$ for any $0\leq k<k_{*}$, where $k_{*}$ is defined via Rule 1 (note that (b) of Lemma 3.1 shows the same for $0\leq k<\hat{k_{\delta}}$).
Lemma 3.3.
Let all the assumptions of Lemma 3.2 hold. Additionally, assume that
((3.22)) $D_{1}:=\kappa^{2}-\Omega(1-\lambda_{k})(1-3\eta)+\omega^{2}L^{2}\geq 0.$
Then, if $u^{\delta}_{k}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<k_{*}$, it follows that $u^{\delta}_{k+1}\in B_{c(\rho)}(\hat{u})\subset B_{\rho+c(\rho)}(u_{0})$, provided $\delta$ is sufficiently small.
Proof.
For $k=0$, result is trivial. By assuming $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<k^{*},$ we show that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})$. From 3.2 and the inequalities $\|u_{k}^{\delta}-\hat{u}\|\leq c(\rho)$ and $\|u_{0}-\hat{u}\|\leq\rho$, it follows that
$\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}$ $-\|u_{k}^{\delta}\ -\hat{u}\|^{2}$
((3.23)) $\leq\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}^{2}+2\omega_{k}(1-\lambda_{k})(1+\eta)\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
$+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}$
$\leq\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}^{2}+\omega_{k}(1-\lambda_{k})(1+\eta)(\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}+\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|^{2})$
$+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}$
$\leq\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}^{2}+\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|^{2}+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)$
$+(\omega_{k}(1-\lambda_{k})(1+\eta)-\left(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\right)\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}.$
Using the fact that $1-\lambda_{k}\leq 1$ and $\lambda_{k}^{2}\leq\lambda_{k}$ for $\lambda\in[0,\frac{1}{2})$, we get
((3.24)) $\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}$ $-\|u_{k}^{\delta}\ -\hat{u}\|^{2}\leq(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda_{k}+\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|^{2}$
$+(\omega_{k}(1-\lambda_{k})(1+\eta)-\left(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\right)\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}.$
Further, the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$ implies that
$\mathbb{E}[\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}-\|u_{k}^{\delta}\ -\hat{u}\|^{2}|\mathcal{F}_{k}]$ $\leq(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda_{k}+\Omega(1-\lambda_{k})(1+\eta)\|y^{\dagger}-y^{\delta}\|^{2}$
$\hskip 28.45274pt+\left(\Omega(1-\lambda_{k})(1-3\eta)-\omega^{2}L^{2}-\kappa^{2})\right)\|F(u_{k}^{\delta})-y^{\delta}\|^{2}.$
By considering the conditional in its whole and utilizing $1-\lambda_{k}\leq 1$, we further reach at
$\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}]-\mathbb{E}[\|u_{k}^{\delta}-\hat{u}\|^{2}]$ $\leq(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda_{k}+\Omega(1+\eta)\mathbb{E}[\|y^{\dagger}-y^{\delta}\|^{2}]$
$\hskip 28.45274pt-D_{1}\mathbb{E}[\|F(u_{k}^{\delta})-y^{\delta}\|^{2}].$
The last inequality derives that
((3.25)) $\begin{split}\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}]+D_{1}\sum_{l=0}^{k}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]&\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k}\lambda_{l}\\ &\hskip 28.45274pt+\Omega(1+\eta)(k+1)\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}]\end{split}$
and hence by using (3.8) and the fact that $k<k_{*}$, where $k_{*}$ is a finite integer, we get
$\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\Omega(1+\eta)k_{*}\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}].$
From Lemma 3.2, one can note that for sufficiently small $\delta$, there exist a constant $\xi$ such that
((3.26)) $\Omega(1+\eta)k_{*}(y^{\delta})\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}]\leq\xi,\ \text{with}\ \rho^{2}+\ (\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\xi\leq c(\rho)^{2}.$
Hence as a result, we get
$\mathbb{E}[\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}]\leq c(\rho)^{2},$
which completes the proof. ∎
Corollary 3.1.
Let the assumptions of Lemma 3.1 be fulfilled. Moreover assume that IRSGD method (3.1) is stopped according to Rule 1, then for sufficiently small $\delta$, we have
$\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq D_{1}^{-1}\left((\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{k}+\rho^{2}+\xi\right)<\infty.$
In particular, for the case $y^{\delta}=y^{\dagger}$ ((((i.e., if $\delta=0)$, we have
$\sum_{l=0}^{\infty}\mathbb{E}[\|F(u_{l})-y^{\dagger}\|^{2}]\leq D_{1}^{-1}\left(\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{\infty}\lambda_{k}\right)<\infty,$
where $D_{1}$ is defined in (3.22).
Proof.
From (3.25), we get
((3.27)) $D_{1}\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{l}+\Omega(1+\eta)k_{*}\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}].$
Using (3.26), we note that
((3.28)) $\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq D_{1}^{-1}\left((\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{k}+\rho^{2}+\xi\right)<\infty.$
Next, for the case $\delta=0$, it follows directly from (3.27) that
((3.29)) $\sum_{l=0}^{\infty}\mathbb{E}[\|F(u_{l})-y^{\dagger}\|^{2}]\leq D_{1}^{-1}\left(\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{\infty}\lambda_{k}\right)<\infty.$
Here we have used the fact that $\sum_{l=0}^{\infty}\lambda_{k}<\infty.$ Thus, (3.28) and (3.29) complete the proof. ∎
Remark 3.3.
If $\lambda_{\text{max}}=0$, the parameter $\kappa$ in conditions (3.7) and (3.22) can be set to zero. This adjustment mirrors the conditions required to establish the same results for the stochastic gradient descent (SGD) method (1.9).
3.3. Convergence for non noisy case
The following result illustrate that the sequence of IRSGD iterates $\{u_{k}\}_{k\geq 1}$ in (3.1) with exact data $y$ forms a Cauchy sequence.
Lemma 3.4.
Let all the assumptions of Lemma 3.1 hold. Additionally assume that
((3.30)) $D_{2}:=(2\omega(1-\lambda_{k})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})\geq 0.$
Then the IRSGD method (3.1) for $\delta=0$ produces a sequence of iterates $\{u_{k}\}_{k\geq 1}$ which is a Cauchy sequence.
Proof.
For a solution $u^{\dagger}$ of (1.2), we define $e_{k}:=u_{k}-u^{\dagger}$. As indicated in part (b) of Lemma 3.1, $\mathbb{E}[\|e_{k}\|^{2}]$ is a bounded sequence and hence it has a convergent subsequence $\mathbb{E}[\|e_{k_{n}}\|^{2}]$. Let it converge to some $\epsilon\geq 0$. It follows directly from 3.2 that, for $\delta=0$ we have
$\|u_{k+1}-u^{\dagger}\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u_{k}-u^{\dagger}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
$+2\lambda_{k}(1-\lambda_{k})\|u_{k}-u^{\dagger}\|\|u_{0}-u^{\dagger}\|$
$-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|y_{i_{k}}-F_{i_{k}}(u_{k})\|^{2}.$
From the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$, we deduce that
$\mathbb{E}[\|u_{k+1}-u^{\dagger}\|^{2}|\mathcal{F}_{k}]$ $\leq\mathbb{E}[(1-\lambda_{k})^{2}\|u_{k}-u^{\dagger}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
$+2\lambda_{k}(1-\lambda_{k})\|u_{k}-u^{\dagger}\|\|u_{0}-u^{\dagger}\||\mathcal{F}_{k}]$
$-\|y-F(u_{k})\|^{2}(2\omega(1-\lambda_{k})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2}).$
By using (3.30) and considering the full conditional, we reach at
$\mathbb{E}[\|u_{k+1}-u^{\dagger}\|^{2}]+D_{2}\mathbb{E}[\|y-F(u_{k})\|^{2}]$ $\leq(1-\lambda_{k})^{2}\mathbb{E}[\|u_{k}-u^{\dagger}\|^{2}]+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
$+2\lambda_{k}(1-\lambda_{k})\mathbb{E}[\|u_{k}-u^{\dagger}\|]\|u_{0}-u^{\dagger}\|.$
This means that
$\mathbb{E}[\|u_{k+1}-u^{\dagger}\|^{2}]^{1/2}\leq(1-\lambda_{k})\mathbb{E}[\|u_{k}-u^{\dagger}\|^{2}]^{1/2}+\rho\lambda_{k}\sqrt{1+\kappa^{-2}\Omega^{2}L^{2}}.$
Hence by induction, we further get that
((3.31)) $\mathbb{E}[\|e_{k}\|^{2}]^{1/2}\leq\mathbb{E}[\|e_{m}\|^{2}]^{1/2}\prod_{l=m}^{k-1}(1-\lambda_{l})+\rho\sqrt{1+\kappa^{-2}\Omega^{2}L^{2}}(1-\prod_{l=m}^{k-1}(1-\lambda_{l}))$
for $m<k,$ where we have used the formula
$1-\prod_{l=m}^{n}(1-\lambda_{l})=\sum_{j=m}^{n}\lambda_{j}\prod_{l=j+1}^{n}(1-\lambda_{l})$
which, under the standard condition that $\prod_{l=n+1}^{n}(1-\lambda_{l})=1$, is satisfied for $n\geq m$.
If $k_{n-1}<k<k_{n}$, then according to (3.31) it follows that
$\mathbb{E}[\|e_{k}\|^{2}]^{1/2}$ $\leq\mathbb{E}[\|e_{k_{n-1}}\|^{2}]^{1/2}+(\rho\sqrt{1+\kappa^{-2}\Omega^{2}L^{2}}-\mathbb{E}[\|e_{k_{n-1}}\|^{2}]^{1/2})\left(1-\prod_{l=k_{n-1}}^{k-1}(1-\lambda_{l})\right),$
$\mathbb{E}[\|e_{k}\|^{2}]^{1/2}$ $\geq\mathbb{E}[\|e_{k_{n}}\|^{2}]^{1/2}-(\rho\sqrt{1+\kappa^{-2}\Omega^{2}L^{2}}-\mathbb{E}[\|e_{k_{n}}\|^{2}]^{1/2})\left(1-\prod_{l=k}^{k_{n-1}}(1-\lambda_{l})\right).$
These estimates, along with the convergence of $\mathbb{E}[\|e_{k_{n}}\|^{2}]$, immediately imply
$\lim_{k\to\infty}\mathbb{E}[\|e_{k}\|^{2}]^{1/2}=\epsilon^{1/2},$
since (3.8) implies that
((3.32)) $\prod_{l=m}^{k-1}(1-\lambda_{l})\to 1\text{ as }m\to\infty,\;k>m.$
Further, to show the convergence of the iterates $u_{k}$, it is crucial to note that $\mathbb{E}[\langle\cdot,\cdot\rangle]$ serves as an inner product. For any $n\geq k$, we select an index $m$ satisfying $n\geq m\geq k$ such that
((3.33)) $\mathbb{E}[\|y-F(u_{m})\|^{2}]\leq\mathbb{E}[\|y-F(u_{i})\|^{2}]\quad\forall k\leq i\leq n.$
Using inequality $\mathbb{E}[\|e_{n}-e_{k}\|^{2}]^{1/2}\leq\mathbb{E}[\|e_{n}-e_{m}\|^{2}]^{1/2}+\mathbb{E}[\|e_{m}-e_{k}\|^{2}]|^{1/2}$ and the identities
$\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{j},e_{m}\rangle]+\mathbb{E}[\|e_{n}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
$\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]+\mathbb{E}[\|e_{k}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
it is sufficient to establish that $\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ and $\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ tend to zero as $k\rightarrow\infty$. As $k\rightarrow\infty$, the last two terms on the right-hand sides of the preceding two identities tend to $\epsilon-\epsilon=0$. This is due to the convergence of $\mathbb{E}[\|e_{k}\|^{2}]$ to $\epsilon$, as discussed above. Next, we proceed to demonstrate that the term $\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]$ also approaches to zero as $k\rightarrow\infty$. In fact, by virtue of the definition of $u_{k}$, we have the following representation
$e_{k}=e_{m}\prod_{l=m}^{k-1}(1-\lambda_{l})+\sum_{j=m}^{k-1}\omega_{j}F_{i_{j}}^{\prime}(u_{j})^{\ast}(y_{i_{j}}-F_{i_{j}}(u_{j}))\prod_{l=j+1}^{k-1}(1-\lambda_{l})+(u_{0}-u^{\dagger})\prod_{l=m}^{k-1}(1-\lambda_{l}).$
for $m<k$. Using this formula, we obtain the estimate
$|\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]|$ $\leq\left(1-\prod_{l=m}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$
$+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y_{i_{j}}-F_{i_{j}}(u_{j}),F_{i_{j}}^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle],$
$|\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]|$ $\leq\left(1-\prod_{l=m}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$
$+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y-F(u_{j}),F^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle]|.$
Since $|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$ is bounded, the first term on the right-hand side tends to zero due to (3.32). Since $1-\lambda_{l}\leq 1$, the second term on the right-hand side can be estimated by
((3.34)) $\Omega(1+3\eta)\sum_{j=m}^{n-1}\mathbb{E}[\|y-F(u_{j})\|^{2}].$
Due to (3.8) and Corollary 3.1, it follows that (3.34) tends to zero as $m\to\infty$. This means $\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]\to 0$. Analogously, we have $\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]\to 0$. Thus, result holds. ∎
In the following theorem, we establish the convergence of non-noisy iterates.
Theorem 3.1.
Suppose that all the assumptions of Lemma 3.1 and (3.30) hold. Then, the IRSGD method (3.1) corresponding to $\delta=0$ generates a sequence of iterates $\{u_{k}\}_{k\geq 1}$ that converges to an element of $\mathcal{D}_{\rho+c(\rho)}(u^{\dagger})$ in $U$, i.e.,
$\lim_{k\to 0}\mathbb{E}[\|u_{k}-\hat{u}\|^{2}]=0,\quad\hat{u}\in\mathcal{D}_{\rho+c(\rho)}(u^{\dagger}).$
Proof.
The proof is similar to [2, Theorem 1]. ∎
3.4. Convergence for noisy case
The next result demonstrates the pathwise stability of the IRSGD iterates $u_{k}^{\delta}$ as $\delta\to 0$. Specifically, for any fixed integer $k\geq 0$ and any sample path $(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}$, it holds that $\|u_{k}^{\delta}-u_{k}\|^{2}\to 0$ as $\delta\to 0$. As a result, the expectation $\mathbb{E}[\|u_{k}^{\delta}-u_{k}\|^{2}]\to 0$ as $\delta\to 0$, confirming the pathwise stability of the iterates for each realization within the filtration $\mathcal{F}_{k}$.
Lemma 3.5.
Let all the assumptions of Lemma 3.1 be satisfied. For any fixed $k\in\mathbb{N}$ and any path $(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}$, let $u_{k}$ and $u_{k}^{\delta}$ be the IRSGD iterates along the path for exact data $y$ and noisy data $y^{\delta}$, respectively. Then
$\lim_{\delta\to 0^{+}}\mathbb{E}[\|u_{k}^{\delta}-u_{k}\|^{2}]=0.$
Proof.
We use induction to prove the assertion. For $k=0$ it is trivial. Suppose that it holds for all indices up to $k$ and any path in $\mathcal{F}_{k}$. From (3.1), for any fixed path $(i_{1},\ldots,i_{k})$, we have
$u_{k+1}^{\delta}-u_{k+1}$ $=(1-\lambda_{k})(u_{k}^{\delta}-u_{k})-\omega_{k}((F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}-F_{i_{k}}^{\prime}(u_{k})^{*})(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})$
$+F_{i_{k}}^{\prime}(u_{k})^{*}((F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}))).$
Thus, by the triangle inequality, we get
$\|u_{k+1}^{\delta}-u_{k+1}\|$ $\leq(1-\lambda_{k})\|u_{k}^{\delta}-u_{k}\|+\omega_{k}\|F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}-F_{i_{k}}^{\prime}(u_{k})^{*}\|\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|$
((3.35)) $+\omega_{k}\|F_{i_{k}}^{\prime}(u_{k})^{*}\|\|(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger})\|.$
Next, we claim that for any specific value of $k$, (i) $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}\|$ is bounded, and (ii) $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}^{\delta}-u_{k}\|$ is uniformly bounded. Indeed, by (3.2) and (3.3), we derive that
$\|u_{k+1}-\hat{u}\|$ $\leq\|u_{k}-\hat{u}\|+\omega_{k}\|F_{i_{k}}^{\prime}(u_{k})^{*}\|\|F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}\|+\lambda_{k}\|u_{0}-u_{k}\|$
$\leq\left(1+\lambda_{k}+\omega_{k}\frac{L^{2}}{1-\eta}\right)\|u_{k}-\hat{u}\|+\lambda_{k}\|u_{0}-\hat{u}\|.$
This, together with an induction argument, establishes the claim (i). Similarly, we note that
$\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|$ $\leq\|F_{i_{k}}(u_{k}^{\delta})-F_{i_{k}}(u_{k})\|+\|F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}\|+\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
((3.36)) $\leq\frac{L}{1-\eta}\left(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|\right)+\|y^{\delta}-y^{\dagger}\|,$
and consequently 3.4 and 3.4 derive that
$\|u_{k+1}^{\delta}-u_{k+1}\|$ $\leq\omega_{k}\left(\frac{L}{1-\eta}\left(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|\right)+\|y^{\delta}-y^{\dagger}\|\right)\|F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}-F_{i_{k}}^{\prime}(u_{k})^{*}\|$
$+\omega_{k}L\|(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger})\|+(1-\lambda_{k})\|u_{k}^{\delta}-u_{k}\|.$
This, along with mathematical induction establishes claim (ii). Let $c=\frac{L}{1-\eta}\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}^{k}}(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|)+\|y^{\delta}-y^{\dagger}\|$. Then, it follows from 3.4, 3.4 and (3.2) that
$\lim_{\delta\to 0^{+}}\|u_{k+1}^{\delta}-u_{k+1}\|$ $\leq(1-\lambda_{k})\lim_{\delta\to 0^{+}}\|u_{k}^{\delta}-u_{k}\|+c\omega_{k}\lim_{\delta\to 0^{+}}\|F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}-F_{i_{k}}^{\prime}(u_{k})^{*}\|$
$+\omega_{k}L\lim_{\delta\to 0^{+}}\|(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger})\|.$
Finally, the result follows from the continuity of the operators $F_{i}$ and $F_{i}^{\prime}$, i.e.,
$\lim_{\delta\to 0^{+}}\|u_{k}^{\delta}-u_{k}\|=0\implies\lim_{\delta\to 0^{+}}\mathbb{E}[\|u_{k}^{\delta}-u_{k}\|^{2}]=0.$
This completes the proof. ∎
The final result of this section establishes the regularizing nature of IRSGD method.
Theorem 3.2.
Let all the assumptions of Lemma 3.1 hold. Let $\{y^{\delta}\}$ be the sequence of noisy data such that Assumption 3.2 hold and let IRSGD iteration corresponding to noisy data is stopped at $k_{*}(y^{\delta})$ defined via Rule 1. Then there exist a solution $\hat{u}$ of (1.2) such that
$\mathbb{E}[\|u_{k_{*}(y^{\delta})}^{\delta}-\hat{u}\|^{2}]\to 0\quad\text{ as}\quad\delta\to 0.$
Proof.
Let $\{y^{\delta_{m}}\}$ be a subsequene of $\{y^{\delta}\}$, and let $K:=\lim\inf_{m\to\infty}k_{*}(y^{\delta_{m}}).$ By taking a subsequence of $\{y^{\delta_{m}}\}$ if necessary, we may assume that $K=\lim_{m\to\infty}k_{*}(y^{\delta_{m}})$ and, according to Lemma 3.5, we have
((3.37)) $\mathbb{E}[\|u_{k}^{\delta_{m}}-u_{k}\|^{2}]\to 0\quad\text{as}\quad m\to\infty.$
for all $0\leq k\leq K.$ Let $\hat{u}$ be the limit of $u_{k}$, which is a solution of (1.2) in $B_{\rho+c(\rho)}(u_{0}).$ First, we assume that $K<\infty,$ then we have $k_{*}(y^{\delta_{m}})=K$ for large $m.$ Thus, from Lemma 3.2 it follows that
$\|F(u_{K}^{\delta_{m}})-y^{\delta_{m}}\|\rightarrow 0\quad\text{as}\quad m\rightarrow\infty.$
By engaging (3.37) and the continuity of $F$, we have $F(u_{K})=y^{\dagger}.$ In other words, $K$th iteration of IRSGD method with exact data is a solution of $F(u)=y^{\dagger}.$ Hence, the iteration terminates with $\hat{u}=u_{K}$ and $u_{K}^{\delta_{m}}\to\hat{u}$ as $m\to\infty.$
If $K=\infty,$ then Theorem 3.1 and Lemma 3.5 guarantees that, for each $\epsilon>0,$ there exist $\hat{k}\in\mathbb{N}$ and $\overline{m}\in\mathbb{N}$ such that
((3.38)) $\mathbb{E}[\|u_{\hat{k}}-\hat{u}\|^{2}]<\frac{\epsilon}{8},\quad\text{and}\quad\mathbb{E}[\|u_{\hat{k}}^{\delta_{m}}-u_{\hat{k}}\|^{2}]<\frac{\epsilon}{8}\quad\text{for all}\quad m\geq\overline{m}.$
At this stage, we apply the full conditional on (3.16) to arrive at
((3.39)) $\mathbb{E}[\|u_{K}^{\delta_{m}}-\hat{u}\|^{2}]\leq 2\mathbb{E}[\|u_{\hat{k}}^{\delta_{m}}-u_{\hat{k}}\|^{2}+2\mathbb{E}[\|u_{\hat{k}}-\hat{u}\|^{2}+C_{2}\sum_{l=\hat{k}}^{K-1}\lambda_{l}+C_{3}\sum_{l=\hat{k}}^{K-1}\frac{1}{(k+a)^{2}},$
where $C_{2}=2\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})$ and $C_{3}=\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau}.$ By choosing $\hat{k}$ such that
$\sum_{l=\hat{k}}^{K-1}\lambda_{l}\leq\frac{\epsilon}{4C_{2}}\quad\text{and}\quad\sum_{l=\hat{k}}^{K-1}\frac{1}{(k+a)^{2}}\leq\frac{\epsilon}{4C_{3}},$
and plugging these estimates along with (3.38) in (3.39), we finally get
$\mathbb{E}[\|u_{K}^{\delta_{m}}-\hat{u}\|^{2}]\leq\epsilon\quad\text{for all}\quad m\geq\overline{m}.$
This completes the proof. ∎
4. Convergence rates
The main results of this section are presented in Theorem 4.3 and Theorem 4.4, which respectively address the rate of convergence under the discrepancy principle (1.7) and the rate of convergence under the heuristic rule 1. As a by-product, convergence rates for SGD (1.9) under these two stopping criteria can also be derived by selecting $\lambda_{k}=0$ in the iterative scheme (3.1). These outcomes signify the second major contribution of the work.
Organisation: This section is organised as follows. At the outset, we introduce additional assumptions necessary for the analysis of convergence rates. Subsections 4.1 and 4.2 are dedicated to Theorems 4.1, 4.2, respectively and these theorems play an important role in deriving the convergence rates. In subsection 4.3, we establish convergence rates under a-posteriori stopping criteria, while subsection 4.4 focuses on deriving convergence rates using heuristic stopping criteria.
We employ the following shorthand notation. Let
((4.1)) $K_{i}=F_{i}^{\prime}(u^{\dagger}),\quad K=\begin{pmatrix}K_{0}\\ \vdots\\ K_{P-1}\end{pmatrix}\ \ \text{and}\quad B=K^{*}K=\sum_{i=0}^{P-1}K_{i}^{*}K_{i}.$
Additionally, we commonly use the abbreviated notation
((4.2)) $\prod_{j}^{k-1}(B_{\lambda})=\prod_{i=j}^{k-1}(I-\lambda_{i}-\omega_{i}B),$
with the convention $\prod_{j}^{k}(B_{\lambda})=I$ for $j>k$. Also, for $h\geq 0,$ and $j\in\mathbb{N},$ we define
((4.3)) $\tilde{h}=h+\frac{1}{2}\quad\text{and}\quad\Theta_{j}^{h}=\|B^{h}\prod_{j+1}^{k-1}(B_{\lambda})\|.$
Note that our primary focus is on the cases where $h=0$ and $h=\frac{1}{2}$.
Following the approach in [14], we assume the following conditions.
Assumption 4.1 (Range invariance condition).
There is a family of bounded operators $\{R_{u}^{i}:u\in B_{\rho+c(\rho)}(u_{0})\}$ for all $i\in\{0,1,\ldots,P-1\}$ such that
$F_{i}^{\prime}(u)=R_{u}^{i}F_{i}^{\prime}(u^{\dagger})\quad\text{and}\quad R_{u}=\text{diag}(R_{u}^{i}):Y^{P}\to Y^{P},$
with
$\|R_{u}-I\|\leq c_{R}\|u-u^{\dagger}\|,\quad u\in B_{\rho+c(\rho)}(u_{0})$
for some positive constant $c_{R}>0$ and $R_{u}=\text{diag}(R_{u}^{i})$ (diagonal matrix).
This assumption has been validated for a range of nonlinear inverse problems, such as parameter identification and nonlinear integral equations (see [8]).
Assumption 4.2 (Source condition [16]).
There exist $\gamma\in(0,1/2)$ and $z\in U$ such that
((4.4)) $u^{\dagger}-u_{0}=(F^{\prime}(u^{\dagger})^{*}F^{\prime}(u^{\dagger}))^{\gamma}z.$
This assumption reflects a particular smoothness requirement for the exact solution $u^{\dagger}$. The restriction on $\gamma$ arises due to technical considerations (see, for example, (7.9)).
Assumption 4.3 (Stochastic range invariance condition).
Under the notations defined in Assumption 4.1, there exist $\theta\in(0,1]$ and $c_{R}>0$ such that for the function $Q(u)=F(u)-y^{\delta}$ or $Q(u)=F^{\prime}(u^{\dagger})(u-u^{\dagger}),u_{k}^{\delta}\in B_{\rho+c(\rho)}(u_{0})$ and $w_{s}=su_{k}^{\delta}+(1-s)u^{\dagger},s\in[0,1],$ there holds
$\mathbb{E}\left[\|(I-R_{w_{s}})Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}}$ $\leq c_{R}\mathbb{E}\left[\|u^{\delta}_{k}-u^{\dagger}\|^{2}\right]^{\frac{\theta}{2}}\mathbb{E}\left[\|Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}},$
$\mathbb{E}\left[\|(I-R^{*}_{w_{s}})Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}}$ $\leq c_{R}\mathbb{E}\left[\|u^{\delta}_{k}-u^{\dagger}\|^{2}\right]^{\frac{\theta}{2}}\mathbb{E}\left[\|Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}}.$
This assumption plays a crucial role in the convergence rates analysis (see [14, Lemma 4.6]) and allows for bounding the terms that involves conditional dependence.
4.1. Recursion on the bias
In this subsection, we derive a recursive formula for the bias of the IRSGD iterate $u_{k}^{\delta}.$ The following lemma gives a useful representation of the mean $\mathbb{E}[e_{k}^{\delta}]$.
Lemma 4.1.
Under the conditions of Assumption 4.1, the error $e_{k}^{\delta}:=u_{k}^{\delta}-u^{\dagger}$ exhibits
((4.5)) $\mathbb{E}[e_{k}^{\delta}]=\Pi^{k-1}_{j=0}(B_{\lambda})e_{0}+\sum_{j=0}^{k-1}\omega_{j}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)-\sum_{j=0}^{k-1}\lambda_{j}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0}),$
with the vector $v_{k}\in Y^{P}$ is given by
((4.6)) $v_{k}=\underbrace{-\left(F(u^{\delta}_{k})-F(u^{\dagger})-K(u^{\delta}_{k}-u^{\dagger})\right)}_{\text{Linearization error}}+\underbrace{(I-R_{u^{\delta}_{k}}^{*})(F(u^{\delta}_{k})-y^{\delta})}_{\text{Range invariance}}.$
Proof.
From the definition of IRSGD iterates $u_{k}^{\delta}$ in (3.1) we have
$e_{k+1}^{\delta}$ $=(1-\lambda_{k})e_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u^{\dagger}-u_{0}).$
The relation $F^{\prime}_{i_{k}}(u^{\delta}_{k})^{\ast}=(R^{i_{k}}_{u^{\delta}_{k}}F^{\prime}_{i_{k}}(u^{\dagger}))^{\ast}=K^{\ast}_{i_{k}}({R^{i_{k}}_{u^{\delta}_{k}}})^{*}$ from Assumption 4.1 and (4.1) leads to
$e^{\delta}_{k+1}=(1-\lambda_{k})e^{\delta}_{k}-\omega_{k}\{K^{\ast}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})+K^{\ast}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})-K^{\ast}_{i_{k}}v_{k,i_{k}}\}-\lambda_{k}(u^{\dagger}-u_{0}),$
with the random variable $v_{k,i_{k}}$ defined by
((4.7)) $v_{k,i_{k}}=-\left(F_{i_{k}}(u^{\delta}_{k})-F_{i_{k}}(u^{\dagger})-K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})\right)+(I-({R^{i_{k}}_{u^{\delta}_{k}}})^{*})(F_{i_{k}}(u^{\delta}_{k})-y^{\delta}_{i_{k}}).$
Consequently, by the measurability of $u^{\delta}_{k}$ (and hence $e^{\delta}_{k}$) with regard to $\mathcal{F}_{k}$, we write
$\mathbb{E}[e^{\delta}_{k+1}\mid\mathcal{F}_{k}]$ $=(I-\lambda_{k}-\omega_{k}B)e^{\delta}_{k}-\omega_{k}K^{\ast}(y^{\dagger}-y^{\delta})+\omega_{k}K^{\ast}v_{k}-\lambda_{k}(u^{\dagger}-u_{0}).$
Then taking the full conditional, using (4.2) and applying the recursion repeatedly yields
$\mathbb{E}[e_{k}^{\delta}]=\Pi^{k-1}_{j=0}(B_{\lambda})e_{0}+\sum_{j=0}^{k-1}\omega_{j}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)-\sum_{j=0}^{k-1}\lambda_{j}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0}).$
This brings the proof to a conclusion. ∎
Remark 4.1.
The term $v_{k}$ in equation (4.6) accounts for both the linearization error of the nonlinear operator $F$, expressed as $F(u^{\delta}_{k})-F(u^{\dagger})-K(u^{\delta}_{k}-u^{\dagger})$, and the range invariance of the derivative $F^{\prime}(u)$ as specified in (ii) of Assumption 3.1 and Assumption 4.1.
The following conclusion provides a reasonable bound on the mean $\mathbb{E}[v_{j}]$.
Lemma 4.2.
Under Assumption 3.1, for $v_{j}$ defined in (4.6), the following inequality holds
((4.8)) $\|\mathbb{E}[v_{j}]\|\leq\frac{(3-\eta)c_{R}}{2(1-\eta)}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{\frac{1}{2}}\mathbb{E}[\|B^{\frac{1}{2}}e^{\delta}_{j}\|^{2}]^{\frac{1}{2}}+c_{R}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{\frac{1}{2}}\delta.$
Proof.
See [14, Lemma 4.3] for the proof, where $c_{R}>0$ denotes the same constant as specified in Assumption 4.1. ∎
Finally, in the next theorem, we establish a bound for $e_{k}^{\delta}$ in terms of a weighted norm.
Theorem 4.1.
Under Assumptions 3.1, 4.1 and 4.2, for any $h\geq 0,$ there holds
((4.9)) $\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|\leq\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}(\delta+\|\mathbb{E}[v_{j}]\|),$
where $\|\mathbb{E}[v_{j}]\|$ is same as given in Lemma 4.2.
Proof.
As $B^{h}$ is a bounded linear operator for $h\geq 0$, by Lemma 4.1, we have
$B^{h}\mathbb{E}[e_{k}^{\delta}]$ $=B^{h}\Pi^{k-1}_{j=0}(B_{\lambda})e_{0}+\sum_{j=0}^{k-1}\omega_{j}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)$
((4.10)) $\hskip 28.45274pt-\sum_{j=0}^{k-1}\lambda_{j}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0}):=I+II+III.$
To establish the validity of the theorem, it is sufficient to bound all three terms $I,II$ and $III$. We will address each of these terms individually. From (4.4), we have
((4.11)) $\|I\|=\|B^{h}\Pi^{k-1}_{j=0}(B_{\lambda})B^{\gamma}z\|\leq\|\Pi^{k-1}_{j=0}(B_{\lambda})B^{\gamma+h}\|\|z\|.$
Next, we estimate the term $II$ as
((4.12)) $\Omega\left\|\sum_{j=0}^{k-1}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)\right\|\leq\Omega\sum_{j=0}^{k-1}\|B^{\tilde{h}}\Pi^{k-1}_{j+1}(B_{\lambda})\|\left(\delta+\|\mathbb{E}[v_{j}]\|\right).$
Likewise, we can also estimate the term $III$ using (4.4) as
$\sum_{j=0}^{k-1}\|\lambda_{j}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0})\|\leq\sum_{j=0}^{k-1}\lambda_{j}\|B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})B^{\gamma}z\|\leq\sum_{j=0}^{k-1}\lambda_{j}\|\Pi^{k-1}_{j+1}(B_{\lambda})B^{\gamma+h}\|\|z\|.$
Plugging this along with (4.11) and (4.12) in 4.1 together with the notation of $\Theta_{j}^{h}$ concludes the proof. ∎
Remark 4.2.
The weighted norm $\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|$ is valuable because the upper bound presented in Theorem 4.1 incorporates $\mathbb{E}[\|B^{\frac{1}{2}}e_{k}^{\delta}\|^{2}]$ (due to $\|\mathbb{E}[v_{j}]\|$) with $h=\frac{1}{2}$. In the context of linear inverse problems where $R_{u}=I$, $c_{R}=0$ and $\|\mathbb{E}[v_{j}]\|=0$ (due to Lemma 4.2), the recursion simplifies to
$\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|\leq\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\delta,$
where these terms correspond to the approximation error and the data error, respectively.
4.2. Recursion on variance
In this subsection, we address the computational variance $\mathbb{E}[\|B^{h}(u^{\delta}_{k}-\mathbb{E}[u^{\delta}_{k}])\|^{2}]$, which arises from the random index $i_{k}$. The variance was first constrained in terms of the iteration noises $H_{j,1}$ and $H_{j,2}$, which are specified below in (4.14).
Lemma 4.3.
Under Assumption 4.1, the IRSGD iterate $u^{\delta}_{k}$ as outlined in (3.1), satisfies
$\mathbb{E}\left[\|B^{h}(u^{\delta}_{k}-\mathbb{E}[u^{\delta}_{k}])\|^{2}\right]$ $\leq\Omega^{2}\left(\sum_{j=0}^{k-1}(\Theta^{\tilde{h}}_{j})^{2}\mathbb{E}[\|H_{j,1}\|^{2}]+2\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta^{\tilde{h}}_{i}\Theta^{\tilde{h}}_{j}\mathbb{E}[\|H_{i,1}\|\|H_{j,2}\|]\right.$
((4.13)) $\left.\hskip 142.26378pt+\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\Theta^{\tilde{h}}_{i}\Theta^{\tilde{h}}_{j}\mathbb{E}[\|H_{i,2}\|\|H_{j,2}\|]\right),$
where the random variables $H_{j,1}$ and $H_{j,2}$, respectively, defined as
((4.14)) $H_{j,1}$ $=(K(u^{\delta}_{j}-u^{\dagger})-K_{ij}(u^{\delta}_{j}-u^{\dagger})\varphi_{ij})+((y^{\dagger}-y^{\delta})-(y^{\dagger}_{i}-y^{\delta}_{i})\varphi_{i_{j}}),$
$H_{j,2}$ $=-\mathbb{E}[v_{j}]+v_{j,i_{j}}\varphi_{i_{j}}.$
Here $v_{k}$ and $v_{k,i}$ are defined in (4.6) and (4.7),respectively, and $\varphi_{i}=(0,\ldots,0,1,0,\ldots,0)$ represents the canonical $i$th basis vector in $\mathbb{R}^{P}$.
Proof.
From the IRSGD iterates as defined in (3.1) along with the Assumption 4.1 that $F^{\prime}_{i_{k}}(u^{\delta}_{k})^{\ast}=(R^{i_{k}}_{u^{\delta}_{k}}F^{\prime}_{i_{k}}(u^{\dagger}))^{\ast}=K^{\ast}_{i_{k}}({R^{i_{k}}_{u^{\delta}_{k}}})^{*}$, we have
((4.15)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}K^{\ast}_{i_{k}}({R^{i_{k}}_{u^{\delta}_{k}}})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
Rearranging the terms of (4.15) to write it as
((4.16)) $u_{k+1}^{\delta}$ $=u_{k}^{\delta}-\omega_{k}\left(K^{*}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})+K^{*}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})+K_{i_{k}}^{*}\left(F_{i_{k}}(u^{\delta}_{k})-F_{i_{k}}(u^{\dagger})\right.\right.$
$\left.\left.-K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})-(I-({R^{i_{k}}_{u^{\delta}_{k}}})^{*})(F_{i_{k}}(u^{\delta}_{k})-y^{\delta}_{i_{k}})\right)\right)-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
Further, from the definition of $v_{k,i}$ as in (4.7), we arrive at
((4.17)) $u^{\delta}_{k+1}=u^{\delta}_{k}-\omega_{k}K^{*}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})-\omega_{k}K^{*}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})+\omega_{k}K^{*}_{i_{k}}v_{k,i_{k}}-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
By the definition of $v_{k}$ in (4.6) and the measurability of $u^{\delta}_{k}$ with respect to $\mathcal{F}_{k}$, we obtain
$\mathbb{E}[u^{\delta}_{k+1}|\mathcal{F}_{k}]=u^{\delta}_{k}-\omega_{k}B(u^{\delta}_{k}-u^{\dagger})-\omega_{k}K^{*}(y^{\dagger}-y^{\delta})+\omega_{k}K^{*}v_{k}-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
Taking the full conditional yields
((4.18)) $\mathbb{E}[u^{\delta}_{k+1}]=\mathbb{E}[u^{\delta}_{k}]-\omega_{k}B\mathbb{E}[u^{\delta}_{k}-u^{\dagger}]-\omega_{k}K^{*}(y^{\dagger}-y^{\delta})+\omega_{k}K^{*}\mathbb{E}[v_{k}]-\lambda_{k}\mathbb{E}[u_{k}^{\delta}-u_{0}].$
Thus, subtracting (4.18) from (4.17) shows that $x_{k}:=u^{\delta}_{k}-\mathbb{E}[u^{\delta}_{k}]$ satisfies
((4.19)) $x_{k+1}=(I-\lambda_{k}-\omega_{k}B)x_{k}+\omega_{k}G_{k},$
with $x_{0}=0$ and the iteration noise $G_{k}$ given by $G_{k}=G_{k,1}+G_{k,2}$, where
$G_{k,1}$ $=(B(u^{\delta}_{k}-u^{\dagger})-K^{*}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger}))+(K^{*}(y^{\dagger}-y^{\delta})-K^{*}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})),$
$\ G_{k,2}$ $=-(K^{*}\mathbb{E}[v_{k}]-K^{*}_{i_{k}}v_{k,i_{k}}).$
Repeatedly applying the recursion (4.19) with $x_{0}=0$ leads to
$x_{k}=\sum_{j=0}^{k-1}\omega_{j}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j}.$
With the decomposition of $G_{j}=G_{j,1}+G_{j,2}$, we directly obtain
((4.20)) $\mathbb{E}[\|B^{h}x_{k}\|^{2}]$ $=\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\mathbb{E}[\langle B^{h}\Pi^{k-1}_{i+1}(B_{\lambda})G_{i,1},B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j,1}\rangle]$
$+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\mathbb{E}[\langle B^{h}\Pi^{k-1}_{i+1}(B_{\lambda})G_{i,1},B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j,2}\rangle]$
$+\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\mathbb{E}[\langle B^{h}\Pi^{k-1}_{i+1}(B_{\lambda})G_{i,2},B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j,2}\rangle]:=B_{I}+B_{II}+B_{III}.$
Next, we proceed to simplify the three terms. Noting that $u^{\delta}_{j}$ is measurable with respect to $\mathcal{F}_{j}$, we have $\mathbb{E}[G_{j,1}|\mathcal{F}_{j}]=0.$ Due to this, we claim that
((4.21)) $\mathbb{E}[\langle B^{h}G_{i,1},B^{h}G_{j,1}\rangle]=0,\quad i\neq j.$
Indeed, for $i>j$, there holds
$\mathbb{E}[\langle B^{h}G_{i,1},B^{h}G_{j,1}\rangle|\mathcal{F}_{i}]=\langle B^{h}\mathbb{E}[G_{i,1}|\mathcal{F}_{i}],B^{h}G_{j,1}\rangle=0,$
and taking the full conditional therefore establishes (4.21). Consequently, $B_{I}$ simplifies to
((4.22)) $B_{I}=\Omega^{2}\sum_{j=1}^{k}\mathbb{E}[\|B^{h}\Pi^{k}_{j+1}(B_{\lambda})G_{j,1}\|^{2}].$
Moreover, for $i>j$, a similar reasoning leads to $\mathbb{E}[\langle B^{h}G_{i,1},B^{h}G_{j,2}\rangle]=0,$ and thus
((4.23)) $B_{II}=2\Omega^{2}\sum_{i=1}^{k}\sum_{j=i}^{k}\mathbb{E}[\langle B^{h}\Pi^{k}_{i+1}(B_{\lambda})G_{i,1},B^{h}\Pi^{k}_{j+1}(B_{\lambda})G_{j,2}\rangle].$
Next, we proceed to further simplify $G_{j,1}$ and $G_{j,2}$. From the definitions of $H_{j,1}$ and $H_{j,2}$, and noting that $(K^{*})^{\dagger}$ represents the pseudoinverse of $K^{*}$, we obtain $(K^{*})^{\dagger}G_{j}=H_{j,1}+H_{j,2}$. Thus, by using triangle inequality along with (4.22) and (4.23) in (4.20), we arrive at
$\mathbb{E}[\|B^{h}x_{k}\|^{2}]$ $\leq\Omega^{2}\sum_{j=0}^{k-1}\mathbb{E}[\|B^{h+1/2}\Pi^{k-1}_{j+1}(B_{\lambda})\|^{2}\|H_{j,1}\|^{2}]$
$+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\|B^{h+1/2}\Pi^{k-1}_{i+1}(B_{\lambda})\|\|B^{h+1/2}\Pi^{k-1}_{j+1}(B_{\lambda})\|\mathbb{E}[\|H_{i,1}\|\|H_{j,2}\|]$
$+\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\|B^{h+1/2}\Pi^{k-1}_{i+1}(B_{\lambda})\|\|B^{h+1/2}\Pi^{k-1}_{j+1}(B_{\lambda})\|\mathbb{E}[\|H_{i,2}\|\|H_{j,2}\|].$
This together with the notation of $\Theta_{j}^{h}$ completes the proof. ∎
The next result provides the bounds on $H_{j,1}$ and $H_{j,2}$ defined in (4.14).
Theorem 4.2.
Let Assumptions 3.1 and Assumption 4.3 be fulfilled. Then for any $h\in[0,\frac{1}{2}]$, there holds
$\mathbb{E}$ $\left[\|B^{h}(\mathbb{E}[x^{\delta}_{k}]-x^{\delta}_{k})\|^{2}\right]\leq\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{\tilde{h}})^{2}\left(\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+\delta\right)^{2}$
$+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{\tilde{h}}\Theta_{j}^{\tilde{h}}(\mathbb{E}\left[\|B^{1/2}e_{i}^{\delta}\|^{2}\right]^{1/2}+\delta)\left(\frac{(2+\theta-\eta)c_{R}}{(1+\theta)(1-\eta)}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}$
$+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\left(\frac{(2+\theta-\eta)c_{R}}{(1+\theta)(1-\eta)}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}\right)^{2}.$
Proof.
We recall from [14, Lemma 4.6], with the modification being the exclusion of the scaling factor $n^{1/2}$, that
((4.24)) $\mathbb{E}[\|H_{j,1}\|^{2}]^{1/2}$ $\leq\left(\mathbb{E}[\|B^{1/2}e^{\delta}_{j}\|^{2}]^{1/2}+\delta\right),$
$\mathbb{E}[\|H_{j,2}\|^{2}]^{1/2}$ $\leq\left(c_{R}(2+\theta-\eta)\frac{(1+\theta)}{(1-\eta)}\mathbb{E}[\|B^{1/2}e^{\delta}_{j}\|^{2}]^{1/2}+c_{R}\delta\right)\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{\theta/2}.$
The conclusion directly follows by using (4.24) in Lemma 4.3. ∎
4.3. Convergence rates using a-posteriori stopping criteria
We recall that our primary objective is to establish the convergence rates using the heuristic stopping criteria (i.e., Rule 1). However, before doing so, in this subsection, we first derive the convergence rates based on the a-posteriori stopping rule (1.7). We emphasize that the convergence analysis for the method with the stopping rule (1.7) follows a similar approach to the one we developed for the heuristic rule in Section 3 with some modification being required due to the change in the stopping rule.
In this subsection, we begin by deriving bounds on the stopping index $k_{\delta}$ and the corresponding error $e_{k}^{\delta}:=u_{k}^{\delta}-u^{\dagger}$.
Lemma 4.4.
Suppose that Assumptions 3.1, 4.1 and 4.2 hold with $\|z\|$ sufficiently small, and let $k_{\delta}$ be the integer specified by (1.7). Also, suppose $y^{\delta}$ satisfies (1.3) and let the weighted parameter sequence $\{\lambda_{k}\}_{k\geq 0}$ be chosen as $\lambda_{k}=(k+1)^{-n}$ with $n>1$, then
((4.25)) $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{1/2}\leq C_{\gamma}\|z\|k^{-\gamma}\quad\text{and}\quad\mathbb{E}[\|B^{1/2}e_{k}^{\delta}\|^{2}]^{1/2}\leq C_{\gamma}\|z\|k^{-\gamma-1/2},$
for $0\leq k<k_{\delta},$ $C_{\gamma}$ is a positive constant depending only on $\gamma,\eta,n,\tau$ and $\theta$. In case of exact data $(\delta=0)$, (4.25) holds for all $k\geq 0.$
Proof.
Using Bias-variance decomposition, we know that
((4.26)) $\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]=\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|^{2}+\mathbb{E}[\|B^{h}(e_{k}^{\delta}-\mathbb{E}[e_{k}^{\delta}])\|^{2}].$
Using assertions of Theorem 4.1 and Theorem 4.2 in (4.26) with $p_{0}=\frac{(2+\theta-\eta)c_{R}}{(1+\theta)(1-\eta)}$, we have
$\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]\leq\left(\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}(\delta+\|\mathbb{E}[v_{j}]\|)\right)^{2}\hskip 85.35826pt$
((4.27)) $+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{\tilde{h}})^{2}\left(\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+\delta\right)^{2}+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{\tilde{h}}\Theta_{j}^{\tilde{h}}\left(\mathbb{E}\left[\|B^{1/2}e_{i}^{\delta}\|^{2}\right]^{1/2}+\delta\right)$
$\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}\right)^{2}.$
By plugging $\|\mathbb{E}[v_{j}]\|$ from Lemma 4.2 in (4.27) and the fact that $\frac{(3-\eta)c_{R}}{2(1-\eta)}<p_{0}$, we get
$\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]\leq\bigg{(}\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}(\delta+p_{0}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{1/2}\mathbb{E}[\|B^{1/2}e^{\delta}_{j}\|^{2}]^{1/2}\hskip 28.45274pt$
$+c_{R}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{1/2}\delta)\bigg{)}^{2}+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{\tilde{h}})^{2}\left(\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+\delta\right)^{2}+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{\tilde{h}}\Theta_{j}^{\tilde{h}}\left(\mathbb{E}\left[\|B^{1/2}e_{i}^{\delta}\|^{2}\right]^{1/2}+\delta\right)$
$\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}\right)^{2}.$
Let $s_{j}=\mathbb{E}[\|e_{j}^{\delta}\|^{2}]$ and $t_{j}=\mathbb{E}[\|B^{1/2}e_{j}^{\delta}\|^{2}]$. Then, by first setting $h=0$ and then $h=1/2$ in the last inequality gives
((4.28)) $s_{k}$ $\leq\left(\Theta_{-1}^{\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1/2}(\delta+p_{0}s_{j}^{1/2}t_{j}^{1/2}+c_{R}\delta s_{j}^{1/2})\right)^{2}++2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{1/2}\Theta_{j}^{1/2}$
$\left(t_{j}^{1/2}+\delta\right)\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1/2})^{2}\left(t_{j}^{1/2}+\delta\right)^{2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1/2}\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}\right)^{2},$
((4.29)) $t_{k}$ $\leq\left(\Theta_{-1}^{1/2+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{1/2+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1}(\delta+p_{0}s_{j}^{1/2}t_{j}^{1/2}+c_{R}s_{j}^{1/2}\delta)\right)^{2}+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{1}\Theta_{j}^{1}$
$\left(t_{j}^{1/2}+\delta\right)\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1})^{2}\left(t_{j}^{1/2}+\delta\right)^{2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1}\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}\right)^{2}.$
Because of (3.3) and (1.7), we have
$\tau\delta<\|y^{\delta}-F(u_{k}^{\delta})\|\leq\delta+\frac{1}{1-\eta}\|Ke_{j}^{\delta}\|.$
From the measuribility of $e_{j}^{\delta}$ with respect to $\mathcal{F}_{j}$ and the inequality $\|Ke_{j}^{\delta}\|=\|B^{1/2}e_{j}^{\delta}\|,$ we get
((4.30)) $\delta<\frac{1}{(\tau-1)(1-\eta)}\mathbb{E}[\|B^{1/2}e_{j}^{\delta}\|^{2}]^{1/2}=c_{\tau\eta}t_{j}^{1/2},\quad\text{where}\quad c_{\tau\eta}=\frac{1}{(\tau-1)(1-\eta)}.$
For the sequence of weighted parameter $\{\lambda_{k}\}_{k\geq 0}$, where $\lambda_{k}=(k+1)^{-n}$ and using (4.30) in (4.28) and (4.29) with the notations $c^{\prime}=p_{0}+c_{R}c_{\tau\eta}$ and $c^{\prime\prime}=1+c_{\tau\eta}$ , we have
((4.31)) $s_{k}\leq\left(\Theta_{-1}^{\gamma}\|z\|+\sum_{j=0}^{k-1}\Theta_{j}^{\gamma}(j+1)^{-n}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1/2}[c_{\tau\eta}+c^{\prime}s_{j}^{1/2}]t_{j}^{1/2}\right)^{2}\hskip 71.13188pt$
$+2\Omega^{2}c^{\prime}c^{\prime\prime}\sum_{i=0}^{k-1}\Theta_{i}^{1/2}t_{i}^{1/2}\sum_{j=0}^{k-1}\Theta_{j}^{1/2}t_{j}^{1/2}s_{j}^{\theta/2}+\Omega^{2}(c^{\prime\prime})^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1/2})^{2}t_{j}+\Omega^{2}(c^{\prime})^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1/2}t_{j}^{1/2}s_{j}^{\theta/2}\right)^{2},$
((4.32)) $t_{k}\leq\left(\Theta_{-1}^{1/2+\gamma}\|z\|+\sum_{j=0}^{k-1}\Theta_{j}^{1/2+\gamma}(j+1)^{-n}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1}[c_{\tau\eta}+c^{\prime}s_{j}^{1/2}]t_{j}^{1/2}\right)^{2}\hskip 42.67912pt$
$+2\Omega^{2}c^{\prime}c^{\prime\prime}\sum_{i=0}^{k-1}\Theta_{i}^{1}t_{i}^{1/2}\sum_{j=0}^{k-1}\Theta_{j}^{1}t_{j}^{1/2}s_{j}^{\theta/2}+\Omega^{2}(c^{\prime\prime})^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1})^{2}t_{j}+\Omega^{2}(p_{0}+c_{R}c_{\tau\eta})^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1}t_{j}^{1/2}s_{j}^{\theta/2}\right)^{2}.$
We claim that for all $0\leq j<k_{\delta}$, for a certain constant $C_{\gamma}$, there holds
((4.33)) $s_{j}^{1/2}\leq C_{\gamma}\|z\|j^{-\gamma}\quad\text{and}\quad t_{j}^{1/2}\leq C_{\gamma}\|z\|j^{-\gamma-1/2}.$
For $j=0$, (4.33) holds trivially. This is because Assumption 4.2 implies that
((4.34)) $\|e_{0}\|\leq\|(F^{\prime}(u^{\dagger})^{*}F^{\prime}(u^{\dagger}))^{\gamma}z\|\leq\|F^{\prime}(u^{\dagger})^{*}F^{\prime}(u^{\dagger})\|^{\gamma}\|z\|\leq\|F^{\prime}(u^{\dagger})\|^{2\gamma}\|z\|.$
Using Assumption 3.1 $(ii)$ in (4.34), we conclude that
$\|e_{0}\|\leq L^{2\gamma}\|z\|\leq C_{\gamma}\|z\|.$
for sufficiently large $C_{\gamma}$. For $j>0$, the proof proceeds by induction. We assume that (4.33) is true for all $0\leq j<k$ for some $k<k_{\delta}$, and then verify it for $j=k$. So using the induction hypothesis in (4.31) and (4.32) along with Lemma A1 and A3 with appropriate choices of $a,b$ and $c$, we have
((4.35)) $s_{k}\leq\bigg{\{}\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}k^{-\gamma}+\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}\left(\frac{1}{1-\gamma}+\frac{1}{1-n}\right)+2^{1-\gamma-n}\|B^{a}\|\right)(k/2)^{1-\gamma-n}\right)\|z\|$
$+\Omega c_{\tau\eta}\hslash\widetilde{C}_{1/2,1/2,0}(k/2)^{-\gamma}+\Omega c^{\prime}\hslash^{2}\widetilde{C}_{1/2,1/2,1/2}(k/2)^{-2\gamma}\bigg{\}}^{2}+2\Omega^{2}c^{\prime}c^{\prime\prime}\hslash^{2+\theta}\widetilde{C}_{1/2,1/2,0}\widetilde{C}_{1/2,1/2,\theta/2}(k/2)^{-\gamma(2+\theta)}$
$+\frac{\Omega^{2}}{2}(c^{\prime\prime})^{2}\hslash^{2}\left(\frac{1}{e\omega}(l_{1,0}+2\ln k)+2^{-2\gamma-1}\|B\|\right)(k/2)^{-2\gamma-1}+\Omega^{2}(c^{\prime})^{2}\hslash^{2(1+\theta)}\widetilde{C}_{1/2,1/2,\theta/2}^{2}(k/2)^{-2\gamma(1+\theta)}.$
Further, using the fact that for any $r>0,$ there holds
((4.36)) $s^{-r}\ln{s}\leq(er)^{-1}\quad\forall s\geq 0,$
and thus $k^{-1}\ln{k}\leq e^{-1},$ we further obtain from (4.35) using the notation $\hslash=C_{\gamma}\|z\|$ that
((4.37)) $s_{k}\leq\hslash^{2}\bigg{\{}\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}2^{-\gamma}+\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}\left(\frac{1}{1-\gamma}+\frac{1}{1-n}\right)+2^{1-\gamma-n}\|B^{a}\|\right)(k/2)^{1-n}\right)C_{\gamma}^{-1}$
$+\Omega c_{\tau\eta}\widetilde{C}_{1/2,1/2,0}+\Omega c^{\prime}\hslash\widetilde{C}_{1/2,1/2,1/2}(k/2)^{-\gamma}\bigg{\}}^{2}+2\Omega^{2}c^{\prime}c^{\prime\prime}\hslash^{\theta}\widetilde{C}_{1/2,1/2,0}\widetilde{C}_{1/2,1/2,\theta/2}(k/2)^{-\gamma\theta}$
$+\frac{\Omega^{2}}{2}(c^{\prime\prime})^{2}\bigg{\{}\left(\frac{l_{1,0}}{e\omega}+2^{-2\gamma-1}\|B\|\right)(k/2)^{-1}+\frac{4}{e^{2}\omega}\bigg{\}}+\Omega^{2}(c^{\prime})^{2}\hslash^{2\theta}\widetilde{C}_{1/2,1/2,\theta/2}^{2}(k/2)^{-2\gamma\theta}\Bigg{]}(k/2)^{-2\gamma}$
$:=\hslash^{2}B_{\gamma,k}(k/2)^{-2\gamma}.\hskip 256.0748pt$
At this point, let us assume that $B_{\gamma,k}>0$ be such that $B_{\gamma,k}\leq 2^{-2\gamma}$ (for sufficiently small $\|z\|$ and sufficiently large $C_{\gamma}$, this condition is satisfied), which means that
((4.38)) $s_{k}\leq\hslash^{2}2^{-2\gamma}\left(\frac{k}{2}\right)^{-2\gamma}\leq\hslash^{2}k^{-2\gamma}=C_{\gamma}^{2}\|z\|^{2}k^{-2\gamma}.$
This completes the induction step and the proof for $s_{k}$. The rest part of the proof is moved to Appendix as it follows on the lines of that of $s_{k}$. ∎
Theorem 4.3.
Under the assumptions of Lemma 4.4, we have
((4.39)) $k_{\delta}\leq C_{\gamma,\eta,n,\tau,\theta}\|z\|^{\frac{2}{2\gamma+1}}\delta^{-\frac{2}{2\gamma+1}}$
where $C_{\gamma,\eta,n,\tau,\theta}$ is a positive constant depending only on $\gamma,\eta,n,\tau,\theta$. In addition to that
((4.40)) $\mathbb{E}[\|e_{k_{\delta}}^{\delta}\|^{2}]^{1/2}\leq C_{\gamma}C_{\gamma,\eta,n,\tau,\theta}\|z\|^{1/(2\gamma+1)}\delta^{2\gamma/(2\gamma+1)},$
[…]
Proof.
We apply (4.30) along with (4.33) with $k=k_{\delta}$ to obtain
$k_{\delta}^{\gamma+1/2}\leq c_{\tau\eta}C_{\gamma}\|z\|/\delta$
with $C_{\gamma}$ as in Lemma 4.4. Hence, we obtain
$k_{\delta}\leq C_{\gamma,\eta,n,\tau,\theta}\|z\|^{\frac{2}{2\gamma+1}}\delta^{-\frac{2}{2\gamma+1}},$
where $C_{\gamma,\eta,n,\tau,\theta}=(c_{\tau\eta}C\gamma)^{\frac{2}{2\gamma+1}}.$ The error estimate (4.40) originates from (4.25). ∎
4.4. Convergence rates using heuristic type stopping criteria
In this section, we will derive an error estimate for the expression $u^{\delta}_{k_{*}(y^{\delta})}-u^{\dagger}$ in mean squared norm.
Lemma 4.5.
Let Assumption 3.1 hold and $y^{\delta}$ be a noisy data satisfying Assumption 3.2, also let $k_{*}:=k_{*}(y^{\delta})$ be the integer defined by Rule 1. Additionally, assume that
((4.41)) $\|F(u_{0})-y^{\delta}\|>\tau\|y^{\dagger}-y^{\delta}\|.$
If Assumption 4.1 holds with $20<\gamma<1/$ 2 for sufficiently small $\|z\|,$ then
((4.42)) $(k_{*}+a)\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}\|z\|^{\frac{2}{2\gamma+1}}\delta^{-\frac{2}{2\gamma+1}},$
where $C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}=\frac{\tau^{2}(1+a)C_{\gamma,\eta,n,\tau,\theta}}{\varkappa^{2}}$ is a constant depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$.
Proof.
By Assumption 3.2 and the definition of $k_{*}$ and $k_{\delta}$, we have
$\varkappa^{2}(k_{*}+a)\delta^{2}\leq\Psi(k_{*},y^{\delta})\leq\Psi(k_{\delta},y^{\delta})=(k_{\delta}+a)\|F(u_{k_{\delta}}^{\delta})-y^{\delta}\|^{2}\leq\tau^{2}(k_{\delta}+a)\delta^{2}.$
This implies that
$\varkappa^{2}(k_{*}+a)\leq\tau^{2}(k_{\delta}+a)\leq\tau^{2}(1+a)k_{\delta},\ \text{as}\ k_{\delta}\geq 1$
since $\|F(u_{0})-y^{\delta}\|>\tau\delta$. Therefore, by virtue of Theorem 4.3, the proof is complete. ∎
Lemma 4.6.
Let Assumption 3.1 hold and let $k_{*}$ be the integer defined by Rule 1 and let $y^{\delta}$ satisfy (1.3). If Assumption 4.1 holds with $20<\gamma<1/$ 2 for sufficiently small $\|z\|,$ then
((4.43)) $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{1/2}\leq\tilde{C}_{\gamma}\|z\|k^{-\gamma},\ \ \ \text{and}\ \ \ \mathbb{E}[\|B^{1/2}e_{k}^{\delta}\|^{2}]^{1/2}\leq\tilde{C}_{\gamma}\|z\|k^{-\gamma-1/2}$
for $0\leq k\leq k_{*},$ $\tilde{C}_{\gamma}>0$ is a constant depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$. In case of exact data $(\delta=0)$, (4.25) holds for all $k\geq 0.$
Proof.
The proof follows on the similar lines as given in Lemma 4.4 with a small modification in (4.30) due to heuristic type stopping rule $k_{*}$. By virtue of Lemma 4.5, we have
((4.44)) $\delta\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma+1/2}\|z\|(k_{*}+a)^{-\gamma-1/2}\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma+1/2}\|z\|j^{-\gamma-1/2}.$
Using (4.44) in (4.28) and (4.29) and following the same procedure as in Lemma 4.4 we get the desired result. ∎
Remark 4.3.
In comparison to Lemma 4.4, the only difference in Lemma 4.6 is a constant factor. This variation arises due to (4.44), which is applied in the proof in place of (4.30).
Theorem 4.4.
Under the assumptions of Lemma 4.6 we have
((4.45)) $\mathbb{E}[\|e_{k_{*}}^{\delta}\|^{2}]^{1/2}\leq\frac{\tilde{C}_{\gamma}}{C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}}\|z\|^{\frac{1}{2\gamma+1}}\delta^{\frac{2\gamma}{2\gamma+1}},$
with some constants $\tilde{C}_{\gamma},C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}>0,$ depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$.
Proof.
We apply (4.42) along with (4.43) with $k=k_{*}$ as
$\mathbb{E}[\|e_{k_{*}}^{\delta}\|^{2}]^{1/2}$ $\leq\tilde{C}_{\gamma}\|z\|C^{-\gamma}_{\gamma,\eta,n,\tau,\theta,\varkappa,a}\left(\frac{\|z\|}{\delta}\right)^{\frac{-2\gamma}{2\gamma+1}}\leq\frac{\tilde{C}_{\gamma}}{C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma}}\|z\|^{\frac{1}{2\gamma+1}}\delta^{\frac{2\gamma}{2\gamma+1}}.$
Hence we get the desired result.
4.5. Comparison of convergence rates
In this subsection, we present a detailed comparison of various iterative methods. As shown in Table 1, the convergence rates achieved by our approach are comparable to those reported in the existing literature. Furthermore, we selected the step size $\omega_{k}\in[\omega,\Omega]$ in such a way that it can either remain constant or decay monotonically, depending on the specific method used.
Method Step-size Stopping criteria Convergence Rate
LM [8] Constant a posteriori (1.7) $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma>0$
IRLM [16] Constant a-priori $O(\delta^{1/2})$
SGD [14] Decaying $k^{*}=\left[(\frac{\delta}{\|w\|})^{-\frac{2}{(2\gamma+1)(1-\alpha)}}\right]$ $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma\in(0,\frac{1}{2})$ and $\alpha$ close to $1$
SGD [12] Decaying a posteriori (1.7) Not available
IRSGD (This paper) Any a posteriori (1.7) $O(\delta^{\frac{2\gamma}{(2\gamma+1)}}),\gamma\in(0,\frac{1}{2})$
IRSGD (This paper) Any heuristic type (Rule 1) $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma\in(0,\frac{1}{2})$
Table 1. Convergence rates of various iterative methods.
5. Numerical Experiments
The objective of this section is to discuss the applicability of proposed scheme using Rule 1 on linear as well as nonlinear ill-posed inverse problems.
5.1. Linear Case
Consider the following problem of linear ill-posed system (see [15] for details on this problem)
[…]
derived from first-kind linear integral equations on $[a,b]$ by sampling at $t_{i}\in[a,b]$ for $i=0,\ldots,P-1,$ where $t_{i}=a+\frac{i(b-a)}{P-1}$ for $i=0,\ldots,P-1,$ and the kernel $k(\cdot,\cdot)$ is continuous on $[a,b]\times[a,b]$. In this case, the iterative scheme (3.1) becomes
[…]
where $i_{k}$ is the randomly chosen index at the $k$th iteration step from the index set $\{0,\ldots,P-1\}$. We consider the system (5.1) with $[a,b]=[-6,6],P=1000,$ and $k(t,s)=\varphi(t-s),$ where $\varphi(t)=(1+\cos(\frac{\pi t}{3}))_{\chi_{\{|t|<3\}}}$ to test the effectiveness of technique (5.2). We consider the following solution to be reconstructed
$u^{\dagger}(s)=\sin\left(\frac{\pi s}{12}\right)+\sin\left(\frac{\pi s}{3}\right)+\frac{1}{200}s^{2}(1-s).$
Instead of the exact data $y_{i}=F_{i}(u^{\dagger})$, we use the noisy data
((5.3)) $y^{\delta}_{i}=y_{i}+\delta_{\text{rel}}\times|y_{i}|\times{\epsilon_{i}},\quad i=0,\ldots,P-1,$
where $\epsilon_{i}$ is a random noise that follows a standard Gaussian distribution $N(0,1)$, and $\delta_{\text{rel}}$ is the relative noise level. Based on the division of $[-6,6]$ into $P-1$ subintervals of similar length, the trapezoidal rule approximates the integrals involved in the procedure and discretized into a linear system of size $n=1000$ with the forward operator $F_{n\times n}$. We execute the method (5.2) with the numerical settings as follows:
  • (a)
    Initial guess $u_{0}=0.$
  • (b)
    The step size is chosen as $\omega_{k}=\omega=10^{-3}.$
  • (c)
    The sequence of weighted parameters $\{\lambda_{k}\}_{k\geq 1}$ is chosen with $\lambda_{k}=\frac{1}{k^{2}}.$
  • (d)
    The maximum number of iterations is set to $k_{\text{max}}=1000$.
  • (e)
    We employ noisy data, where $\delta_{\text{rel}}=10^{-1},10^{-2},10^{-3}$ and $10^{-4}$ represent four different relative noise levels.
We mainly perform the three experiments as:
5.1.1. Convergence behavior of $\Psi(k,y^{\delta})$
In order to investigate the convergence behavior of $\Psi(k,y^{\delta})$ as $\delta\to 0$, we present the plots of the values of $\Psi(k,y^{\delta})$ versus $k$ in Figure 1 using the fixed constant $a=100$ (as in Rule 1) and for different levels of noise.
Figure 1. Convergence behavior of $\Psi(k,y^{\delta})$ as defined in Rule 1, alongside the increase in the stopping index $k_{*}$ as the noise level decreases.
Figure 1 clearly shows that the noise level in the input data has a significant effect on the convergence behavior of $\Psi(k,y^{\delta})$. As $\delta_{\text{rel}}$ decreases from $10^{-1}$ to $10^{-4}$, convergence becomes both faster and more stable. Additionally, it demonstrates that $\Psi(k_{*},y^{\delta})\to 0$ as $\delta\to 0$, which supports the assertion in Lemma 3.2. Furthermore, it is also evident that as the noise level decreases, the stopping index $k_{*}$ increases.
5.1.2. Reconstructed solutions
We reconstructed the solution using (5.2) at various noise levels, as shown in Figures LABEL:M=0.1-LABEL:M=0.001. These results illustrate the impact of the constant $a$ (as defined in Rule 1) on reconstruction accuracy. At a noise level of $10^{-1}$, smaller values of $a$ lead to early stopping of the algorithm, resulting in an inaccurate reconstruction. As $a$ increases (e.g., $a=100$ and $a=1000$), the reconstruction accuracy improves significantly.
From Figures LABEL:M=0.01 and LABEL:M=0.001 observe that, as the noise level decreases, the inaccuracy decreases for small values of $a$ (e.g., $a=1$ and $a=10$). However, for $a=1000$, the number of iterations increases significantly compared to the other cases, leading to a higher computational burden. Therefore, for lower noise levels, smaller values of $a$ should be chosen to reduce computational cost and achieve faster results.
Tables 2 provide detailed numerical results (stopping index $k_{*}$ and the relative errors $E^{\delta}_{k}:=\frac{\|u^{\delta}_{k}-u^{\dagger}\|_{L^{2}(\Omega)}}{\|u^{\dagger}\|_{L^{2}(\Omega)}}$) for Rule 1 at various noise levels.
$\delta_{\text{rel}}$ $a$ in Rule 1
[…]
$k_{*}$ $E_{k_{*}}^{\delta}$ $k_{*}$ $E_{k_{*}}^{\delta}$ $k_{*}$ $E_{k_{*}}^{\delta}$ $k_{*}$ $E_{k_{*}}^{\delta}$
$1\text{e}^{-1}$ 18 $3.550\text{e}^{-2}$ 32 $1.565\text{e}^{-2}$ 49 $1.441\text{e}^{-2}$ 249 $4.213\text{e}^{-3}$
$1\text{e}^{-2}$ 133 $4.817\text{e}^{-3}$ 80 $8.626\text{e}^{-3}$ 506 $1.300\text{e}^{-3}$ 565 $1.664\text{e}^{-3}$
$1\text{e}^{-3}$ 172 $2.293\text{e}^{-3}$ 260 $4.982\text{e}^{-3}$ 2044 $5.845\text{e}^{-4}$ 3404 $4.982\text{e}^{-4}$
$1\text{e}^{-4}$ 3304 $5.341\text{e}^{-4}$ 3167 $5.134\text{e}^{-4}$ 867 $6.111\text{e}^{-4}$ 4874 $3.185\text{e}^{-4}$
Table 2. Results of IRSGD with different values of $a$ in Rule 1.
5.1.3. Comparison between SGD and IRSGD
We compare the results of IRSGD with SGD. For this comparison, we set the constant $a$ in the rule (referred to as Rule 1) to $a=100$. The numerical results for the comparison are presented in Table 3.
Methods $\delta_{\text{rel}}$ $k_{*}$ $E_{k_{*}}^{\delta}$ $\Psi(k_{*},y^{\delta})$
IRSGD $1e^{-1}$ 68 $1.58e^{-2}$ 139237.06
SGD $1e^{-1}$ 66 $3.34e^{-2}$ 153135.32
IRSGD $1e^{-2}$ 180 $2.49e^{-3}$ 24514.94
SGD $1e^{-2}$ 192 $1.44e^{-3}$ 27692.34
IRSGD $1e^{-3}$ 258 $5.53e^{-3}$ 20785.89
SGD $1e^{-3}$ 772 $3.40e^{-4}$ 8772.18
IRSGD $1e^{-4}$ 3492 $7.93e^{-4}$ 21809.47
SGD $1e^{-4}$ 4122 $4.59e^{-4}$ 12432.96
IRSGD $1e^{-5}$ 3468 $6.24e^{-4}$ 17168.44
SGD $1e^{-5}$ 3529 $7.71e^{-4}$ 17952.29
Table 3. Comparison of IRSGD and SGD with Rule 1.
We observe that, as noise level decreases (moving down the table), both methods generally require more iterations ($k_{*}$) to achieve an acceptable solution. Lower noise levels yield lower relative errors $E_{k_{*}}^{\delta}$, which is expected as there is less perturbation in the data, allowing both methods to approximate the true solution more accurately. For different noise levels, IRSGD tends to converge in fewer iterations than SGD, indicating that the additional damping term in IRSGD can speed up convergence. However, for very low noise ($\delta_{\text{rel}}=1\times 10^{-5}$), the iteration counts for IRSGD and SGD are similar, suggesting that both methods stabilize in a similar time frame.
5.2. Nonlinear case
This subsection presents the mathematical model that illustrates the Schlieren tomography and validates the proposed methods. The Schlieren optical system produces a light intensity pattern that is proportional to the square of the line integral of pressure along the light path as it passes through the tank (see [10] for more details).
For each $i$ between 00 and $P-1$, let $\sigma_{i}\in S^{1}$ be a set of recording directions such that $\zeta_{i}=\zeta(\theta_{i})=(\cos\theta_{i},\sin\theta_{i})$. The problem is to reconstruct a function $u$ supported on a bounded domain $D:=[-1,1]\times[-1,1]\subset\mathbb{R}^{2}$ from the system of operator equations
$F_{i}(u)=y_{i}^{\delta},\quad\text{for }i=0,\ldots,P-1,$
where the Schlieren transform $F_{i}=R_{i}^{2}$, with the operator $R_{i}$ denoting the Radon transform along the direction $\zeta_{i}$, satisfying $R_{i}:C_{0}^{\infty}(D)\to C_{0}^{\infty}(I)$:
[…]
From [10], we know that each operator $F_{i}:H_{0}^{1}(D)\to L^{2}(I)$ with $I=[-1,1]$ is continuous and Fréchet differentiable with
$F_{i}^{\prime}(u)h=2R_{i}(u)R_{i}(h),\quad\forall h\in H_{0}^{1}(D).$
In addition, the adjoint $F_{i}^{\prime}(u)^{*}:L^{2}(I)\to H_{1}^{0}(D)$ is given by
$F_{i}^{\prime}(u)^{*}g=(I-\Delta)^{-1}\left(2R_{i}^{*}(R_{i}(u)g)\right),\quad\forall g\in L^{2}(I),$
where $R_{i}^{*}:L^{2}(I)\to L^{2}(D)$ is the adjoint of $R_{i}$ given by $(R_{i}^{*}g)(u)=g(u\cdot\sigma_{i}),$ $I$ refers to the identity operator and $\Delta$ represents the Laplace operator acting on $H_{0}^{1}(D)$. The computation of the adjoint operator requires solving an elliptic boundary value problem, which is approximated using a multigrid method implemented through finite difference discretization.
In this numerical experiment, we set the parameters as follows: $\delta_{\text{rel}}=10^{-2}$, $k_{\text{max}}=750$, $\omega_{k}=\omega=5\times 10^{-3}$, and $\lambda_{k}=\frac{1}{k^{3}}$. The true image and initial guess are shown in Figure LABEL:initial_and_true_figures. Using the proposed method, we reconstructed the image with a stopping criterion based on Rule 1. We evaluated the reconstruction quality by tracking both the relative errors and the progression of $\Psi(k,y^{\delta})$ for different values of $a$: specifically, $a=10^{3}$ and $10^{4}$. The results for each case are displayed in Figures LABEL:RE=1000-LABEL:RE=_10000, respectively.
6. Conclusion
In this paper, we present an Iteratively Regularized Stochastic Gradient Descent (IRSGD) method specifically developed to address systems of ill-posed inverse problems, utilizing the Hanke-Raus heuristic rule. Notably, the proposed IRSGD method is fully data-driven, as the Hanke-Raus rule operates independently of any prior knowledge of the noise level. Our convergence analysis is based on the tangential cone condition and assumes Lipschitz continuity of the operators involved.
Furthermore, we have conducted a thorough analysis of the convergence of our method in the mean squared norm. Additionally, we have established convergence rates under both a-posteriori and heuristic rules by leveraging the source condition, range invariance condition, and stochastic range invariance condition for arbitrary step sizes. To demonstrate the practical effectiveness of our approach, we have provided extensive numerical experiments on both linear and nonlinear inverse problems.
Lastly, we outline several directions for future research. First, it is essential to validate our assumptions and theoretical analysis in the context of specific nonlinear inverse problems to further demonstrate the practical applicability of the IRSGD method. Second, following the work in [2, 4, 5, 19], the IRSGD approach can be extended to handle nonsmooth nonlinear problems, thereby broadening its range of applications. Finally, it is important to investigate adaptive step size rules to provide practical insights and guidelines for real-world implementation. The heuristic rule (based on the modified discrepancy principle), which is used in this analysis can be further modified to study the stochastic analysis of regularization techniques presented in [5, 16, 18, 19, 20].
Acknowledgments
Harshit Bajpai would like to express gratitude to the Science and Engineering Research Board (SERB), India, for providing his PhD fellowship through grant CRG/2022/005491. Ankik Kumar Giri also extends his thanks to SERB for supporting his research under the project “Fast and efficient iterative methods for solving nonlinear inverse problems” funded by grant CRG/2022/005491.

### v3

@@
 Abstract.
-In this work, we present a novel variant of the stochastic gradient descent method termed as iteratively regularized stochastic gradient descent (IRSGD) method to solve nonlinear ill-posed problems in Hilbert spaces.
-Under standard assumptions, we demonstrate that the mean square iteration error of the method converges to zero for exact data.
-In the presence of noisy data, we first propose a heuristic parameter choice rule (HPCR) based on the method suggested by Hanke and Raus, and then apply the IRSGD method in combination with HPCR.
-Precisely, HPCR selects the regularization parameter without requiring any a-priori knowledge of the noise level.
+Over the past decade, stochastic algorithms have emerged as scalable and efficient tools for solving large-scale ill-posed inverse problems by randomly selecting subsets of equations at each iteration.
+However, due to the ill-posedness and measurement noise, these methods often suffer from oscillations and semi-convergence behavior, posing challenges in achieving stable and accurate reconstructions.
+This study proposes a novel variant of the stochastic gradient descent (SGD) approach, the iteratively regularized stochastic gradient descent (IRSGD) to address nonlinear ill-posed problems in Hilbert spaces.
+Under standard assumptions, we demonstrate that the mean square iteration error of the method tends to zero for exact data.
+In the presence of noisy data, we first propose a heuristic parameter choice rule (HPCR) and then apply the IRSGD method in combination with HPCR.
+Precisely, HPCR selects the regularization parameter without requiring any a-priori information of the noise level.
 We show that the method terminates in finitely many steps in case of noisy data and has regularizing features.
-Further, we discuss the convergence rates of the method using well-known source and other related conditions under HPCR as well as discrepancy principle.
-To the best of our knowledge, this is the first work that establishes both the regularization properties and convergence rates of a stochastic gradient method using a heuristic type rule in the setting of infinite-dimensional Hilbert spaces.
-Finally, we provide the numerical experiments to showcase the practical efficacy of the proposed method.
-Keywords: Stochastic gradient descent; Iterative Landweber method; Convergence rates; Variational inequalities; Inverse problems; Nonlinear ill-posed problems; Regularization.
+Finally, some numerical experiments are performed to demonstrate the practical efficacy of the method.
+Keywords: Unknown noise level; Stochastic gradient descent; Heuristic stopping rules; Nonlinear ill-posed problems; Inverse problems; Schlieren tomography.
+MSC codes: 65J22, 65J20, 47J06
 1. Introduction
-This work focuses on deriving an approximate solution for a system of nonlinear ill-posed equations of the form
-((1.1)) $F_{i}(u)=y^{\dagger}_{i},\hskip 5.69054pti=0,1,\ldots,P-1,$
-[…]
+This work focuses on deriving an approximate solution for a system of nonlinear ill-posed equations
+((1.1)) $F_{i}(u)=y^{\dagger}_{i},$
+where $i=0,1,\ldots,P-1,$ and for each $i$, the mapping $F_{i}:\mathfrak{D}(F_{i})\subset U\rightarrow Y_{i}$ represents a nonlinear operator between the Hilbert spaces $U$ and $Y_{i}$ with domain $\mathfrak{D}(F_{i})$.
+In this context, $y_{i}^{\dagger}\in Y_{i}$ represents the exact data, $P\in\mathbb{N}$, $U$ and $Y_{i}$ are endowed with their standard inner products and norms, respectively.
+Systems of the form (1.1) commonly appear in various real-life applications.
+For instance, they are fundamental in various tomography methods used in fields such as medical imaging and industrial diagnostics [7, 20, 29].
+Conventionally, for the space $Y^{P}=Y_{0}\times Y_{1}\times\cdots\times Y_{P-1}$ equipped with the natural inner product derived from those of $Y_{i}$, (1.1) can be interpreted as
 ((1.2)) $F(u)=y^{\dagger},$
-where $F:U\rightarrow Y^{P}$ is defined as
-$F(u)=\begin{pmatrix}F_{0}(u)\\ \vdots\\ F_{P-1}(u)\end{pmatrix}\ \ \text{and}\ \hskip 5.69054pty^{\dagger}=\begin{pmatrix}y^{\dagger}_{0}\\ \vdots\\ y^{\dagger}_{P-1}\end{pmatrix}.$
-Usually, exact data $y^{\dagger}$ is never precisely available. Instead, a noisy data $y^{\delta}\in Y^{P}$ is available such that
-((1.3)) $\|y^{\dagger}-y^{\delta}\|\leq\delta,$
-for noise level $\delta\geq 0$. Here, we write $\delta=\sqrt{\delta_{0}^{2}+\delta_{1}^{2}+\cdots+\delta_{P-1}^{2}}$ and $y^{\delta}=(y^{\delta}_{0},y^{\delta}_{1},...,y^{\delta}_{P-1})$, where
-((1.4)) $\|y_{i}^{\delta}-y_{i}^{\dagger}\|\leq\delta_{i}\ \ \text{for}\ \ i=0,1,2,\cdots,P-1.$
-Due to ill-posedness of (1.1), its solution may not exist and even if it exists, it may not be unique.
-Moreover, the solution(s) can be highly sensitive to perturbations in the noisy data $y^{\delta}$ (see [6]).
-Consequently, regularization methods are needed for deducing the stable approximate solution of (1.1).
-One of the prominent and highly effective classes of regularization methods is that of iterative regularization.
-This approach has achieved significant success in solving a wide variety of inverse problems (see [6, 16] and the references therein).
-Among these methods, the Landweber Iteration Method (LIM) is a well-studied and widely recognized classical iterative technique.
-For a Fréchet differentiable (or smooth) forward mapping $F:U\rightarrow Y$, LIM can be written as
-((1.5)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta}),\hskip 28.45274ptk\geq 0.$
-Here $u_{0}^{\delta}:=u_{0}$represents an initial guess that incorporates the prior knowledge about the solution to be reconstructed and $F^{\prime}(u_{k}^{\delta})^{*}$ is the adjoint of the Fréchet derivative of $F$ at $u_{k}^{\delta}$.
-Scherzer [26] introduced a modification to the method (1.5) by introducing a damping term and named it as iteratively regularized Landweber iteration method. This method can be written as
-((1.6)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta})-\lambda_{k}(u_{k}^{\delta}-u^{(0)}),\hskip 28.45274ptk\geq 0,$
-where $\lambda_{k}\geq 0$.
-In comparison to the method (1.5), the convergence rates analysis of the method (1.6) needs fewer assumptions on the mapping $F$.
-Furthermore, the method (1.6) converges to a solution which is closed to $u^{(0)}$, however, the method (1.5) needs additional assumptions for the similar convergence behaviour (see [16] for nice details).
-This means that the inclusion of the additional damping term in (1.6) has several advantages over the method (1.5).
-When the noisy data is available, a proper stopping rule (a-posteriori) must be employed in order to show that the iterative schemes (1.5) and (1.6) are a regularization one.
-The commonly incorporated stopping rule is well known as discrepancy principle, i.e., the method is halted after $k_{\delta}=k_{\delta}(\delta,y^{\delta})$ steps, where
+where the mapping $F:U\rightarrow Y^{P}$ is given by
+$F(u)=\begin{pmatrix}F_{0}(u)\\ \vdots\\ F_{P-1}(u)\end{pmatrix},\ \ \text{and}\ \hskip 5.69054pty^{\dagger}=\begin{pmatrix}y^{\dagger}_{0}\\ \vdots\\ y^{\dagger}_{P-1}\end{pmatrix}$.
+The exact data $y^{\dagger}$ is typically never available. Rather, a noisy data $y^{\delta}\in Y^{P}$ with the noise level $\delta\geq 0$ is accessible such that
+((1.3)) $\|y^{\dagger}-y^{\delta}\|\leq\delta.$
+Here, we write $\delta=\sqrt{\delta_{0}^{2}+\delta_{1}^{2}+\cdots+\delta_{P-1}^{2}}$ and $y^{\delta}=(y^{\delta}_{0},y^{\delta}_{1},...,y^{\delta}_{P-1})$, where
+((1.4)) $\|y_{i}^{\delta}-y_{i}^{\dagger}\|\leq\delta_{i}\ \ \text{for}\ \ i=0,1,\ldots,P-1.$
+The solution to (1.1) might not exist because of its ill-posedness, and even if it does, it is unlikely to be unique.
+Moreover, the solution(s) can be highly sensitive to perturbations in $y^{\delta}$ (see [5]).
+Therefore, regularization techniques are required to find the stable approximate solution of (1.1).
+Iterative regularization is a well-established and highly effective class of regularization techniques.
+This approach has achieved significant success in solving a wide variety of inverse problems, (see [5, 22]).
+Among these methods, the Landweber Iteration Method (LIM) is a widely recognized classical iterative technique.
+For the mapping $F:U\rightarrow Y^{P}$ that is Fréchet differentiable, LIM is
+((1.5)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta}),\hskip 28.45274ptk\geq 0,$
+where $u_{0}^{\delta}:=u_{0}$ is an initial guess and $F^{\prime}(u_{k}^{\delta})^{*}$ represents the adjoint of the Fréchet derivative of $F(u_{k}^{\delta})$.
+The method (1.5) was modified by Scherzer [33], who included a damping term and called it the iteratively regularized Landweber iteration method. For $\lambda_{k}\geq 0$, the iteration takes the form
+((1.6)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta})-\lambda_{k}(u_{k}^{\delta}-u^{(0)}),$
+where $k\geq 0$. The $k$-th iteration of ((1.6)) can be viewed as performing a gradient descent step on the functional
+$J(u)=\frac{1}{2}\left(\|F(u)-y^{\delta}\|^{2}+\lambda_{k}\|u-u^{(0)}\|^{2}\right)=\frac{1}{2}\left(\sum_{i=0}^{P-1}\|F_{i}(u)-y_{i}^{\delta}\|^{2}+\lambda_{k}\|u-u^{(0)}\|^{2}\right).$
+Compared to the method (1.5), the analysis of deriving convergence rates for the method (1.6) requires less assumptions on $F$.
+Moreover, the method (1.6) converges to a solution that remains close to $u^{(0)}$, whereas achieving a similar convergence behavior with the method (1.5) needs additional assumptions (see [22]).
+This highlights the advantage of incorporating the extra damping term in (1.6) as compared to (1.5).
+When the noisy data is available, it is necessary to use an appropriate stopping rule (a-posteriori) to illustrate that the iterative schemes (1.5) and (1.6) are regularization methods.
+The prevalent stopping rule is referred to as discrepancy principle [28], which states that the method is stopped after $k_{\delta}=k_{\delta}(\delta,y^{\delta})$ steps, where
 ((1.7)) $\|F(u_{k_{\delta}}^{\delta})-y^{\delta}\|\leq\tau\delta<\|F(u_{k}^{\delta})-y^{\delta}\|,\hskip 28.45274pt0\leq k<k_{\delta},$
-for some $\tau>1.$ In real-world applications, noise level is not always available or trustworthy, which makes the discrepancy principle (1.7) difficult to use.
-Consequently, while applying the discrepancy principle, overestimation or underestimation of the noise level may result in significant degradation of the reconstruction accuracy.
-Therefore, it is reasonable to work on heuristic stopping rules that do not rely on the knowledge of noise level.
-Real [25] proposed a heuristic rule for the Landweber iteration (1.5) in Banach space setting based on the discrepancy principle, by taking motivation from the method of Hanke and Raus [9].
+for some $\tau>1.$ However, it is challenging to apply the discrepancy principle (1.7) in practicality since the noise level is not always reliable or available.
+Consequently, while applying the discrepancy principle, overestimation or underestimation of the noise level could lead to serious deterioration of the reconstruction accuracy.
+Therefore, it is reasonable to work on heuristic stopping rules that do not depend on the information of noise level.
+Real [32] proposed a heuristic rule for the Landweber iteration (1.5) based on the discrepancy principle, which is further motivated by the method of Hanke and Raus [9].
 For (1.5), heuristic rule determines an integer $k_{*}:=k_{*}(y^{\delta})$ by minimizing
 ((1.8)) $\Psi(k,y^{\delta}):=(k+a)\|F({u_{k}^{\delta}})-y^{\delta}\|^{p}$
-over $k$, where $a\geq 1$ is a fixed number and $p>0$ is a constant.
-It can be observed that for the conventional iterative regularization methods for solving inverse problems (such as the Landweber method [16], Levenberg-Marquardt method [5], conjugate gradient method [16], non-stationary Tikhonov iterative regularization method [20] and iteratively regularized Gauss-Newton method [16], among others), one of the common challenges is the significant computational burden at each iteration, mainly due to the need to process all the data, which can be excessively large and resource-intensive.
-To address this issue, a potential and effective approach is the utilization of stochastic gradient descent (SGD), a method introduced by Robbins et al.
-[24].
-Further, Jin et al.
-[13] studied the convergence analysis of SGD method and demonstrated its regularizing nature when there is a noise in the data.
-The basic version of SGD can be written as (see [14])
+over $k$, where $p>0$ is a constant and $a\geq 1$ is a fixed number.
+It may be observed that the significant computational cost per iteration is a common problem with traditional iterative regularization techniques mainly due to the need to process complete dataset.
+Applying stochastic gradient descent (SGD) [15, 16, 17, 31] is a viable and successful way to deal with this problem.
+Using [16], the SGD method can be expressed as
 ((1.9)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\mu_{k}F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}),\hskip 28.45274ptk\geq 0,$
-where $\mu_{k}$ is the corresponding step size and $i_{k}$ is the uniformly drawn index from the index set $\{0,1,...,P-1\}.$ It can be noted that the method (1.9) is a randomized version of (1.5).
-For (1.9), the stopping index $k(\delta)\in\mathbb{N}$ is defined using the following a-priori stopping rule
+where $i_{k}$ is a uniformly drawn random index from $\{0,1,\ldots,P-1\}$ and $\mu_{k}\geq 0$ is the step size. For the SGD method (1.9), the stopping index $k(\delta)\in\mathbb{N}$ satisfies the rule
 ((1.10)) $\lim_{\delta\rightarrow 0^{+}}k(\delta)=\infty\hskip 14.22636pt\text{and}\hskip 14.22636pt\lim_{\delta\rightarrow 0^{+}}\delta^{2}\sum_{i=1}^{k(\delta)}\mu_{i}=0.$
-We refer to [14, 17] for some recent literature on SGD method using a-priori stopping rule for linear as well nonlinear inverse problems and [12, 27] on SGD method using a-posteriori stopping rule for linear inverse problems.
-It is evident from these references that they exclusively use an a-priori or a-posteriori stopping rule in developing their convergence theories.
-From this observation, it is of significant interest to develop heuristic rules for stochastic gradient descent (SGD) type methods.
-Therefore, the primary objective of this paper is to explore the convergence analysis of the following iteratively regularized stochastic gradient descent (IRSGD) under heuristic stopping rules:
+Notably, the aforementioned rule is an a-priori rule.
+We refer to [16, 24] for some recent work on the SGD method using a-priori stopping rule and [13, 35] for the SGD method using a-posteriori stopping rule.
+To the best of our knowledge, convergence results of SGD method via heuristic stopping rules are not yet available in the literature.
+Accordingly, the primary objective of this study is to explore the convergence analysis of the following iteratively regularized stochastic gradient descent (IRSGD) under heuristic stopping rules:
 ((1.11)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}),\hskip 14.22636ptk\geq 0,$
 $\hskip 85.35826pt\text{with}\quad 0\leq\lambda_{k}\leq\lambda_{\text{max}}<\frac{1}{2},$
-where $\omega_{k}$ is the step size, $\lambda_{k}$ is the weighted parameter, $F^{\prime}_{i}(u)$ is the Fréchet derivative of $F_{i}(u)$, $i_{k}$ is an index randomly drawn from the index set $\{0,1,\ldots,P-1\}$ through uniform distributation.
-This method differs from the SGD method in the sense that an extra damping term $-\lambda_{k}(u_{k}^{\delta}-u_{0})$ is considered in its formulation.
+where $\lambda_{k}$ is the weighted parameter, $\omega_{k}$ is the step size.
+This method differs from the SGD method in the sense that an extra damping term $\lambda_{k}(u_{k}^{\delta}-u_{0})$ is considered in its formulation.
 A key advantage of studying IRSGD lies in its ability to reflect SGD by setting the damping factor $\lambda_{k}=0$.
-Our approach is novel in the realm of stochastic methods for solving ill-posed inverse problems as it provides convergence analysis without relying on an a-priori stopping rule.
+Our approach is novel in the realm of stochastic methods for solving ill-posed problems as it provides convergence analysis without relying on an a-priori stopping rule.
 Additionally, it circumvents the need to know the noise level, which further enhances its practical applicability.
-For SGD method, Jin et al.
-[14] established the convergence rate in the mean squared norm in the context of nonlinear ill-posed problems, assuming an appropriate source condition.
-Further, Jahn et al.
-[12] further extended these results using an a-posteriori stopping rule, however, precise estimates for the convergence rates are not provided in that work.
-Inspired by these findings, one of the main aims of this work is to develop convergence rates for the iterates generated by (1.11) under an a-posteriori stopping rule.
-In addition to that, we also derive convergence rates under heuristic stopping criteria.
-To the best of our knowledge, this is the first study that addresses convergence rates of a stochastic type method for solving nonlinear ill-posed problems under heuristic stopping rule.
-The main contributions of this paper are outlined as follows:
-  • Instead of focusing on the analysis of a single equation, we turn our attention to exploring stochastic methods for solving systems comprising a large number of equations (i.e., $P$ is large in (1.1)).
-Stochastic methods involve the random consideration of each equation within equation (1.1) independently, thereby significantly reducing memory requirements.
-This approach finds relevance in numerous real-world scenarios (see [7, 21, 22]).
-  • The method is fully data-driven, as it is applicable without using a-priori or a-posteriori stopping rule.
-This is the first attempt of construction of a stochastic method using heuristic type stopping rule.
-We establish the (strong) convergence of IRSGD method, which improves the weak convergence (i.e., convergence in probability) results provided in [12].
-  • We establish the convergence rates for the proposed scheme (1.11) using a step size $\omega_{k}\in[\omega,\Omega]$ and a decaying weight parameter, under two types of stopping criteria: the discrepancy principle and a heuristic rule.
-The rates are comparable to those provided in [14, Theorem 2.5], [8, Theorem 3.2], and [13, Theorem 2.2] (see Table 1).
-  • This work presents experimental evidence showcasing the method’s enhanced efficiency and accelerated performance achieved through the inclusion of a damping factor in SGD.
-This modification yields significantly better results compared to the standard SGD approach discussed in [14].
-The rest of the paper is structured as follows.
-In Section 2, we list the preliminaries needed in our work.
-Section 3 covers the convergence analysis of our method in the presence of noisy and exact data.
-Section 4 discusses the convergence rates of our method for a-posteriori and heuristic stopping rules under certain assumptions.
-Section 5 presents numerical experiments on a system of ill-posed problems, demonstrating the applicability of the proposed scheme with heuristic stopping rule.
-The final section offers concluding remarks.
-Additionally, the Appendix includes few auxiliary lemmas required for our analysis.
-2. Preliminaries
-In this section, we review certain basic definitions and notation related to our work, see [2, 4, 14, 15, 23] for more details.
-The notation $u_{k}$ is used to denote the iterates of our method corresponding to the exact data $y^{\dagger}.$ Additionally, we refer $\mathcal{F}_{k}$ as the filtration generated by the random indices $\{i_{1},...,i_{k-1}\}$ up to the $(k-1)$th iteration.
-Because we randomly select the indices $i_{k}$, the iteration $u_{k}^{\delta}$ in the IRSGD method becomes a random process.
-To assess its convergence, multiple methods are available.
-In our case, we will use the mean squared norm defined by $\mathbb{E}[\|\cdot\|^{2}]$, where $\mathbb{E}[\cdot]$ represents the expectation taken with respect to the filtration $\mathcal{F}_{k}$.
-It is important to note that the iterates $u_{k}^{\delta}$ are measurable with respect to $\mathcal{F}_{k}$ (see [15] for more details).
-The following well-known identities will be frequently utilized: for any random variables $\Theta,\Theta_{1},\Theta_{2}$ and $\Phi$, and real constants $c_{1},c_{2}$, we have
-  • (a)
-    $\mathbb{E}[\Theta]=\mathbb{E}[\mathbb{E}[\Theta|\mathcal{F}_{k}]].$
-  • (b)
-    Additive: $\mathbb{E}[\Theta_{1}+\Theta_{2}|\mathcal{F}_{k}]=\mathbb{E}[\Theta_{1}|\mathcal{F}_{k}]+\mathbb{E}[\Theta_{2}|\mathcal{F}_{k}].$
-    Linearity: $\mathbb{E}[c_{1}\Theta+c_{2}|\mathcal{F}_{k}]=c_{1}\mathbb{E}[\Theta|\mathcal{F}_{k}]+c_{2}.$
-  • (c)
-    If $\Theta\leq\Phi$, then $\mathbb{E}[\Theta|\mathcal{F}_{k}]\leq\mathbb{E}[\Phi|\mathcal{F}_{k}]$ almost surely.
-Here the expression $\mathbb{E}[\Theta|\mathcal{F}_{k}]$ represents the expected value of $\Theta$ given the information contained in $\mathcal{F}_{k}$.
-Triangular inequality for expectation: Let $\Theta_{1}$ and $\Theta_{2}$ be random variables. Then, there holds
-((2.1)) $\mathbb{E}[\|\Theta_{1}+\Theta_{2}\|]\leq\mathbb{E}[\|\Theta_{1}\|]+\mathbb{E}[\|\Theta_{2}\|].$
-Jensen’s inequality: For any random variable $\Theta$ and for any convex function $f$, there holds
-((2.2)) $f(\mathbb{E}[\|\Theta\|])\leq\mathbb{E}[f(\|\Theta\|)],$
-provided $\mathbb{E}[\|\Theta\|]<\infty$ and $\mathbb{E}[f(\|\Theta\|)]<\infty$ (see [1] for more details).
-Next, we introduce an important result that splits the mean squared error $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]$, where $e_{k}^{\delta}:=u_{k}^{\delta}-u^{\dagger}$ represents the error for any iteration $u_{k}^{\delta}$.
-This decomposition is a fundamental strategy in error analysis and plays a crucial role in deriving the convergence rates for iterative methods.
-Bias-variance decomposition: The bias-variance decomposition for the iterative scheme is given by
-((2.3)) $\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]=\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|^{2}+\mathbb{E}[\|B^{h}(e_{k}^{\delta}-\mathbb{E}[e_{k}^{\delta}])\|^{2}],$
-where $B$ is defined as in (4.1) and $h\geq 0.$ In this decomposition:
-  • The first term, $\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|^{2}$, represents the bias, which includes both the approximation error and the error due to noise in the data.
-  • The second term, $\mathbb{E}[\|B^{h}(e_{k}^{\delta}-\mathbb{E}[e_{k}^{\delta}])\|^{2}]$, represents the variance, which accounts for the randomness introduced by the selection of the index $i_{k}$ during the iterations.
-The bias reflects the deterministic error from the method’s approximation, while the variance captures the stochastic error due to the randomized nature of the iterative process.
-We end this section by discussing few notations.
-For some $u\in U$ and $\rho>0,$ we denote the open and closed balls in $U$ of radius $\rho$ centered at $u$, respectively by $B_{\rho}(u)$ and $\overline{B}_{\rho}(u).$ Let $\mathcal{D}_{\rho}(u^{\dagger})$ denote the set of all the solutions of (1.2) in $\overline{B}_{\rho}(u^{\dagger})$, i.e., $\mathcal{D}_{\rho}(u^{\dagger}):=\{u\in\overline{B}_{\rho}(u^{\dagger}):F(u)=y^{\dagger}\}$.
+We perform the detailed convergence analysis of the proposed method with early stopping, which is guaranteed to terminate after a finite number of iterations.
+Unlike the a priori stopping rules, which require a deterministic stopping index, our method yields a stopping index that is inherently random due to its dependence on the sample paths.
+This randomness introduces analytical challenges that cannot be addressed using the techniques in [16, 20, 24].
+To overcome these, we develop a convergence framework grounded in Hilbert space theory and probability.
+Our convergence framework, which addresses the random stopping index defined by a heuristic rule, is inspired by the recent work [11], where random stopping based on the discrepancy principle is studied in the context of stochastic mirror descent.
+By constructing a probability-one event, pathwise convergence is established for the exact data case, and, together with a stability result, this leads to almost sure convergence in the noisy case as the noise level approaches zero.
+The main outcomes of this study are listed as follows:
+  • The proposed method is fully data-driven, requiring neither a-priori nor a-posteriori stopping rules.
+To the best of our knowledge, this is the first work to formulate and analyze a stochastic method with a heuristic-type stopping criterion.
+  • Unlike prior works [13, 15, 16, 17] that employ a polynomially decaying step size, we allow the step size $\omega_{k}$ to vary arbitrarily within a fixed interval $\omega,\Omega]$.
+We establish convergence of the proposed IRSGD scheme ((1.11)) in Hilbert spaces under a heuristic stopping rule.
+The analysis is technically challenging due to the randomness of the stopping index and the inclusion of a damping term.
+  • We extend the scope from analyzing a single equation to developing and studying stochastic methods for solving large-scale systems of equations (i.e., with large $P$ in ((1.1))).
+By randomly selecting and processing individual equations from ((1.1)) independently, our approach drastically reduces memory requirements, making it particularly relevant to large-scale inverse problems arising in practical applications [7, 29, 30].
+  • Through extensive numerical experiments, we demonstrate that incorporating a damping factor into the SGD framework along with the heuristic stopping significantly improves both reconstruction quality and convergence speed compared to standard SGD variants, confirming the practical benefits of the proposed approach.
+Outline.
+Section 2 introduces the required assumptions, presents the IRSGD algorithm, and develops the analysis for the heuristic stopping rule, including a result ensuring the iterates remain within a certain ball.
+Section 3 provides convergence results for both noisy and exact data.
+Section 4 reports numerical experiments illustrating the effectiveness of the method, and the final section concludes the paper.
 2. Iteratively regularized Stochastic gradient descent in Hilbert spaces
-This section covers the convergence analysis of IRSGD method under certain assumptions.
-Precisely, this section is organised as follows.
-The subsection 3.1 aims to discuss the assumptions required for studying the convergence analysis.
-In subsection 3.2, we discuss a heuristic rule based on the modified discrepancy principle.
-Convergence of non-noisy iterates is discussed in subsection 3.3.
-Lastly, subsection 3.4 aims to demonstrate the convergence analysis of noisy iterates.
-We recall that the IRSGD method can be expressed as
-((3.1)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}),\hskip 14.22636ptk\geq 0,$
-with $\text{with}\ 0\leq\lambda_{k}\leq\lambda_{\text{max}}<\frac{1}{2}$, where $F^{\prime}_{i_{k}}$ is the Frećhet derivative of $F_{i_{k}}$, $u_{0}^{\delta}$ is the initial guess of the exact solution, $\omega_{k}$ is the step size and the index $i_{k}$ is drawn randomly from the index set $\{0,1,\ldots,P-1\}$ at each iteration step.
+To investigate the convergence behavior of the iterates $u_{k}^{\delta}$ in the IRSGD method, we consider the mean squared norm $\mathbb{E}[\|\cdot\|^{2}]$.
+In this context, the expectation $\mathbb{E}[\cdot]$ is taken with respect to the filtration $\mathcal{F}_{k}$, which is determined by the sequence of random indices $\{i_{1},\dots,i_{k-1}\}$ (see [20]).
+We denote by $B_{\rho}(u)$ the closed ball in $U$ of radius $\rho$ centered at $u$, where $u\in U$ and $\rho>0$.
+Furthermore, let $\mathcal{D}_{\rho}(u^{\dagger})$ represent the set of all solutions to (1.2) within the closed ball $B_{\rho}(u^{\dagger})$, which is formally given by
+$\mathcal{D}_{\rho}(u^{\dagger}):=\{u\in B_{\rho}(u^{\dagger}):F(u)=y^{\dagger}\}.$
+The remainder of this section is structured as follows: Subsection 2.1 introduces the foundational assumptions required for the convergence analysis and provides a detailed exposition of the IRSGD method, including a heuristic stopping criterion as presented in Algorithm 1.
+Subsection 2.2 discusses a heuristic stopping rule based on a modified discrepancy principle.
+In this context, we establish the almost sure finiteness of the stopping index defined by this principle and present several key theoretical results related to the proposed scheme.
 2.1. Assumptions
-In our work, we need the following assumptions.
+Our analysis demands the following assumptions.
 Assumption 2.1.
-For $i\in\{0,1,\ldots,P-1\}$ and $\rho>0$, there holds:
+The following is true for $i\in\{0,1,\ldots,P-1\}$, and $\rho>0$:
   • (i)
-    The operator $F:U\rightarrow Y^{P}$ is continuous.
+    $F:U\rightarrow Y^{P}$ is a continuous operator.
   • (ii)
-    There exists a constant $L>0$ such that
-    ((3.2)) $\|F_{i}^{\prime}(u)\|\leq L\ \text{for all}\ u\in B_{\rho+c(\rho)}(u_{0}),$
-    where $L\leq\frac{1}{2\omega}$ (see Lemma 3.1 for the meaning of $c(\rho)$ and $\omega$). Moreover there is a constant $10<\eta<$ 1 such that
-    ((3.3)) $\|F_{i}(\overline{u})-F_{i}(u)-F_{i}^{\prime}(u)(\overline{u}-u)\|\leq\eta\|F_{i}(\overline{u})-F_{i}(u)\|$
+    There exists a constant $0<L\leq\frac{1}{2\omega}$ such that
+    ((2.1)) $\|F_{i}^{\prime}(u)\|\leq L\ \text{for all}\ u\in B_{\rho+c(\rho)}(u_{0}).$
+    (see Lemma 2.1 for the meaning of $c(\rho)$ and $\omega$.) Moreover, constant $\eta$ exists such that
+    ((2.2)) $\|F_{i}(\overline{u})-F_{i}(u)-F_{i}^{\prime}(u)(\overline{u}-u)\|\leq\eta\|F_{i}(\overline{u})-F_{i}(u)\|,\ \ 0<\eta<1$
     for all $\overline{u},u\in B_{\rho+c(\rho)}(u_{0}).$
@@
     (1.2) has a solution $\hat{u}\in B_{\rho}(u_{0}).$
-Motivated by [9], we consider the following heuristic rule for IRSGD in Hilbert spaces.
-Rule 1.
-Let $a\geq 1$ be a fixed number and let
-$\Psi(k,y^{\delta}):=(k+a)\|F(u^{\delta}_{k})-y^{\delta}\|^{2}.$
-We define $k_{*}:=k_{*}(y^{\delta})$ to be an integer such that
+The conditions in Assumption 2.1, including the tangential cone condition ((2.2)) that quantifies the nonlinearity of the forward operator $F_{i}$, are standard in convergence analyses of iterative regularization for nonlinear inverse problems [8, 16, 22] and have been verified for various settings such as nonlinear integral equations and parameter identification in PDEs [8].
+Motivated by [9], we consider the following heuristic stopping rule for the IRSGD method.
+Rule 1 (Heuristic).
+Let $a\geq 1$ be a fixed constant such that
+$\Psi(k,y^{\delta}):=(k+a)\sum_{i=0}^{P-1}\|F_{i}(u^{\delta}_{k})-y_{i}^{\delta}\|^{2}.$
+We define $k_{*}:=k_{*}(y^{\delta})$ as the integer satisfying
 $k_{*}\in\arg\min\{\Psi(k,y^{\delta}):0\leq k\leq k_{\infty}\},$
-where $k_{\infty}:=k_{\infty}(y^{\delta})$ is the largest integer such that $u^{\delta}_{k}\in\mathfrak{D}(F)$ for all $0\leq k<k_{\infty}$.
+where $k_{\infty}:=k_{\infty}(y^{\delta})$ denotes the largest integer for which $u^{\delta}_{k}\in\mathfrak{D}(F)$ holds for all $0\leq k<k_{\infty}$.
+Using the integer $k_{*}:=k_{*}(y^{\delta})$ defined by Rule 1, we call $u^{\delta}_{k_{*}(y^{\delta})}$ as an approximate solution.
+This gives rise to the main algorithm of the paper, presented as follows.
+Algorithm 1 IRSGD with heuristic stopping
+1:Given: $F_{i},y_{i}^{\delta}$ for all $i\in\{0,1,\ldots,P-1\}$ and a constant $a\geq 1$
+2:Initialize: $u_{0}^{\delta}=u_{0}$, set $k=0$
+3:repeat
+4:  Choose $\omega_{k}\in[\omega,\Omega]$ and the weight parameter $\lambda_{k}$ such that
+$0\leq\lambda_{k}\leq\lambda_{\text{max}}<\frac{1}{2}.$
+5:  Randomly pick an index $i_{k}$ from $\{0,1,\ldots,P-1\}$ uniformly.
+6:  Update
+$u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}),$
+where $F^{\prime}_{i_{k}}$ is the Fréchet derivative of $F_{i_{k}}$, and $u_{0}$ is the initial guess.
+7:  Set $k\leftarrow k+1$
+8:until
+$\Psi(k,y^{\delta}):=(k+a)\sum_{i=0}^{P-1}\|F_{i}(u^{\delta}_{k})-y_{i}^{\delta}\|^{2}$
+does not achieve its minimum (for the details see heuristic rule 1)
+9:return  $u^{\delta}_{k_{*}(y^{\delta})}$
 Remark 2.1.
 Rule 1 is straightforward to implement in practice, as detailed in Section 4.
-During the iteration process, the value of $\Psi(k,y^{\delta})$ is monitored as a function of $k$.
+Throughout the iterative procedure, the value of $\Psi(k,y^{\delta})$ is monitored as a function of $k$.
 After a sufficient number of iterations, the process is terminated and $k_{*}$ is selected that minimizes $\Psi(k,y^{\delta})$.
@@
 Therefore, the limit on the upper bound must be sufficiently large to avoid inaccurately estimating $k_{*}$ at the first local minimum, which is often not the global minimum [12].
-To prevent $k_{*}$ from becoming too small, we have performed several numerical experiments on different values of $a$ in Rule 1.
-Note that in order to prove Lemma 3.1 for IRSGD, it is essential to assume that $a\geq 1$.
-This Lemma will be further required to show the convergence of the scheme.
-Using the integer $k_{*}:=k_{*}(y^{\delta})$ determined by Rule 1, we call $u^{\delta}_{k_{*}(y^{\delta})}$ as an approximate solution.
-A natural question arises concerning a family of noisy data $\{y^{\delta}\}$, where $y^{\delta}\to y^{\dagger}$ as $\delta\to 0$ is that how to guarantee that $u^{\delta}_{k_{*}(y^{\delta})}$ converges to a solution of (1.2) as $\delta\to 0\mathord{?}$
-In general, the answer is no, as stated by Bakushinskii’s veto [3], which asserts that heuristic rules cannot guarantee the convergence in the worst-case scenario for any regularization method.
-Later, in [9], it is shown that the following assumption guarantee the convergence.
-We also need this condition in our work.
+Remark 2.2.
+Evaluating the residual $\|F(u_{k}^{\delta})-y^{\delta}\|^{2}=\sum_{i=0}^{P-1}\|F_{i}(u_{k}^{\delta})-y_{i}^{\delta}\|^{2}$ at every iteration of Algorithm 1 for heuristic rule 1 can be computationally expensive.
+A practical alternative is to perform this computation at prescribed intervals, following the approach used in stochastic gradient methods [13, 21], or to exploit randomized SVD techniques in problems with intrinsic low-rank structure [23].
+Similar computational overheads have been reported for the discrepancy principle [6, 11].
+Mitigating these costs in the heuristic setting, potentially by adapting the strategy in [11], is a promising direction for future research.
+Algorithm 1 leads to the question of how to ensure that $u^{\delta}_{k_{*}(y^{\delta})}$ converges to a solution of (1.2) as $\delta\to 0.$ Specifically, we seek to determine the conditions under which this convergence holds for a sequence of noisy data $\{y^{\delta}\}$, where $y^{\delta}\to y^{\dagger}$ as $\delta\to 0$.
+In general, the answer is inadequate, as stated by Bakushinskii’s veto [2], which asserts that heuristic rules cannot guarantee convergence in the worst-case scenario for any regularization method.
+Subsequently, in [9], a noise condition was introduced for linear problems in Hilbert spaces.
+This condition was later generalized under the Hanke-Raus rule (see Assumption 2.2) to examine the convergence of nonlinear regularization methods as well as iterative regularization methods in Banach spaces [18, 19, 32, 36].
+We also utilize this condition in our work.
 Assumption 2.2 (Noise condition).
-The family of noisy data $\{y^{\delta}\}$ satisfies the condition $0<\|y^{\delta}-y^{\dagger}\|\rightarrow 0$ as $\delta\rightarrow 0$.
+The family of noisy data $\{y^{\delta}\}$ satisfies the condition $\|y^{\delta}-y^{\dagger}\|\rightarrow 0$ as $\delta\rightarrow 0$.
 Moreover, there exists a constant $\varkappa>0$ such that for every $y^{\delta}$ and for each $u\in S(y^{\delta})$, where $S(y^{\delta}):=\{u^{\delta}_{k}:0\leq k\leq k_{\infty}\}$ is constructed using (1.11), there holds
 $\|y^{\delta}-F(u)\|\geq\varkappa\|y^{\delta}-y^{\dagger}\|.$
-Under Assumption 3.2, it is immediate to see that Rule 1 comes out with a finite integer $k_{*}$.
-Indeed, if $k_{\infty}$ is finite, then there is nothing to prove.
-So we only consider that $k_{\infty}=\infty$.
-It then follows from Assumption 3.2 that
+It directly follows from Assumption 2.2 that Rule 1 yields a finite integer $k_{*}$.
+Indeed, if $k_{\infty}$ is finite, then $k_{*}$ is trivially finite.
+Therefore, we assume $k_{\infty}=\infty$.
+From Assumption 2.2 there holds
 $\Psi(k,y^{\delta})=(k+a)\|F(u^{\delta}_{k})-y^{\delta}\|^{2}\geq(k+a)\varkappa\|y^{\dagger}-y^{\delta}\|^{2}\rightarrow\infty$
-as $k\rightarrow\infty$. Consequently, we can conclude that a finite integer $k_{*}$ exists, which achieves the minimum of $\Psi(k,y^{\delta})$.
-3.2. Analysis for Rule 1
+as $k\rightarrow\infty$. Consequently, a finite integer $k_{*}$ exists that achieves the minimum of $\Psi(k,y^{\delta})$.
+Remark 2.3.
+It is worth emphasizing that Assumption 2.2, while abstract and difficult to verify analytically, currently offers the most practical framework for theoretical analysis, as also noted in [32].
+Its validity can, at present, only be supported through numerical experiments.
+A detailed numerical investigation of heuristic parameter choice rules, beyond the Hanke–Raus rule, for the nonlinear Landweber iteration is provided in [12].
+2.2. Analysis for Heuristic Rule 1
 This subsection aims to demonstrate that
-((3.4)) $\Psi(k_{*}(y^{\delta}),y^{\delta})\rightarrow 0\quad\text{as}\quad\delta\rightarrow 0.$
-Additionally, we demonstrate that IRSGD iterates $u_{k}^{\delta}$ remains in a ball around $u_{0}$ under certain conditions.
-In order to show (3.4), we propose an auxiliary integer $\hat{k}_{\delta}$ defined by the stopping rule
-((3.5)) $\|F(u^{\delta}_{\hat{k}_{\delta}})-y^{\delta}\|+\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y^{\delta}-y^{\dagger}\|<\|F(u^{\delta}_{k})-y^{\delta}\|+\frac{M}{k+a}\quad$
-for $0\leq k<\hat{k}_{\delta}$, where $M=\frac{\varrho c(\rho)^{2}\tau}{\Omega(1+\varrho)(1+\eta)},\varrho>0$ is a constant, $\Omega>0,\eta$ is same as in (3.3), $c(\rho)$ is defined in (3.6) and $\tau>1$ is a real number (not to be confused with the one in (1.7)).
-Adding an extra term in the classical discrepancy principle produces a minor variation of the discrepancy principle (1.7).
-An essential component in establishing (3.4) is provided by this additional term, which ensures that $\hat{k}_{\delta}\rightarrow\infty$ as $\delta\rightarrow 0$.
+((2.3)) $\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\rightarrow 0\quad\text{as}\quad\delta\rightarrow 0.$
+We show the same in Lemma 2.2. Additionally, we demonstrate that IRSGD iterates $u_{k}^{\delta}$ remains in a ball around $u_{0}$ under certain conditions.
+To establish ((2.3)), we introduce an auxiliary index $\hat{k}_{\delta}$, determined by the following stopping criterion.
+Rule 2 (Modified discrepancy).
+The stopping index $\hat{k}_{\delta}$ is defined as the smallest integer satisfying
+((2.4)) $\|F_{i}(u^{\delta}_{\hat{k}_{\delta}})-y_{i}^{\delta}\|+\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|<\|F_{i}(u^{\delta}_{k})-y_{i}^{\delta}\|+\frac{M}{k+a},\quad\forall i=0,1,\ldots,P-1,$
+for all $0\leq k<\hat{k}_{\delta}$, where
+$M=\frac{\varrho\,c(\rho)^{2}\,\tau}{\Omega\,(1+\varrho)\,(1+\eta)},$
+with $\varrho>0$, $\Omega>0$, $\eta$, $\tau>1$, and $c(\rho)$ defined in ((2.2)), ((1.7)), and ((2.5)), respectively.
+This modified discrepancy principle represents a slight augmentation of the classical version given in ((1.7)), achieved by incorporating an additional penalization term.
+This term plays a pivotal role in the convergence analysis, particularly in establishing the result in ((2.3)), as it guarantees that $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$.
+Remark 2.4.
+It is important to emphasize that both the stopping indices $k_{*}$ and $\hat{k}_{\delta}$ depend on the iterates $u_{k}^{\delta}$, which are themselves random variables.
+Consequently, $k_{*}$ and $\hat{k}_{\delta}$ are also random quantities, adding a substantial layer of complexity to the theoretical analysis compared to the deterministic settings considered in [18, 19, 32].
+[…]
+Algorithm 2 IRSGD with modified discrepancy stopping
+1:Given: operators $F_{i}$ and noisy data $y_{i}^{\delta}$ for $i=0,1,\dots,P-1$; constant $a\geq 1$; parameters $\omega,\Omega>0$, $\lambda_{\max}\in[0,\tfrac{1}{2})$; discrepancy parameters $\tau>1$ and $M>0$ (as in ((2.4))).
+2:Initialize: $u_{0}^{\delta}=u_{0}$ (initial guess), set $k\leftarrow 0$.
+3:repeat
+4:  Choose $\omega_{k}\in[\omega,\Omega]$ and $\lambda_{k}$ with $0\leq\lambda_{k}\leq\lambda_{\max}$.
+5:  Select index $i_{k}$ uniformly at random from $\{0,1,\dots,P-1\}$.
+6:  Compute update
+$u_{k+1}^{\delta}\;=\;u_{k}^{\delta}-\omega_{k}\,F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}\big{(}F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\big{)}-\lambda_{k}\,(u_{k}^{\delta}-u_{0}).$
+7:  Set $k\leftarrow k+1$.
+8:until the modified-discrepancy condition ((2.4)) is satisfied for the current iterate, i.e., the index $\hat{k}_{\delta}$ defined by
+$\|F_{i}(u^{\delta}_{\hat{k}_{\delta}})-y_{i}^{\delta}\|+\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|\quad\text{for all }i=0,1,\dots,P-1,$
+holds (see Rule 2).
+9:return  $u^{\delta}_{\hat{k}_{\delta}}$
 Lemma 2.1.
-Let Assumption 3.1 hold and let $\Omega,\omega$ be positive constants such that $\Omega\geq\omega$. Moreover, let $\kappa\in(0,1)$ and $\varrho>0$ be a sufficiently small constant and
+Let Assumption 2.1 hold and let $\Omega,\omega$ be positive constants such that $\Omega\geq\omega_{k}\geq\omega$.
+Moreover, let $\kappa\in(0,1)$ and $\varrho>0$ be a sufficiently small constant (see Remark 2.5) and
 ((2.5)) $c(\rho):=\rho\lambda_{\textbf{max}}\frac{(1-\lambda_{\textbf{max}})+\sqrt{(1-\lambda_{\textbf{max}})^{2}+(\frac{1}{1+\varrho}-(1-\lambda_{\text{max}})^{2})(1+\kappa^{-2}\Omega^{2}L^{2})}}{(\frac{1}{1+\varrho}-(1-\lambda_{\text{max}})^{2})}.$
@@
 ((2.7)) $\sum_{k=0}^{\infty}\lambda_{k}\leq\lambda<\infty,$
-there holds
-  • (a)
-    the stopping index $\hat{k}_{\delta}$ defined in (3.5) is finite for $\delta>0$ and $\hat{k}_{\delta}\rightarrow\infty$ as $\delta\rightarrow 0.$
-  • (b)
-    for a solution $\hat{u}$ to (1.2) and for any $0\leq k<\hat{k}_{\delta}$, if $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$, then $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})\subset B_{\rho+c(\rho)}(u_{0}).$
+[…]
 Proof.
-First we prove assertion (b).
-For $k=0$, result trivially follows from (iii) of Assumption 3.1.
-By assuming that $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<\hat{k}_{\delta}$, we demonstrate that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u}).$ From (3.1), we have
-$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $=\|u_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0})-\hat{u}\|^{2}$
-$=(1-\lambda_{k})^{2}\|u^{\delta}_{k}-\hat{u}\|^{2}+\lambda_{k}^{2}\|u_{0}-\hat{u}\|^{2}+2\lambda_{k}(1-\lambda_{k})\langle u^{\delta}_{k}-\hat{u},u_{0}-\hat{u}\rangle$
-((3.9)) $\hskip 14.22636pt+\omega_{k}^{2}\|F^{\prime}_{i_{k}}(u^{\delta}_{k})^{*}(y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}))\|^{2}+2\omega_{k}(1-\lambda_{k})\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F^{\prime}_{i_{k}}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle$
+[…]
+$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $=(1-\lambda_{k})^{2}\|u^{\delta}_{k}-\hat{u}\|^{2}+\lambda_{k}^{2}\|u_{0}-\hat{u}\|^{2}+2\lambda_{k}(1-\lambda_{k})\langle u^{\delta}_{k}-\hat{u},u_{0}-\hat{u}\rangle$
+[…]
 $\hskip 14.22636pt+2\omega_{k}\lambda_{k}\langle u_{0}-\hat{u},F^{\prime}_{i_{k}}(u^{\delta}_{k})^{*}(y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}))\rangle.$
-To estimate 3.2, we can write the last two terms of the sum as
-$\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F_{i_{k}}^{\prime}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle$ $=\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),y_{i_{k}}^{\delta}-y_{i_{k}}^{\dagger}\rangle-\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
-((3.10)) $\hskip 5.69054pt-\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F_{i_{k}}(u^{\delta}_{k})-F_{i_{k}}(\hat{u})-F_{i_{k}}^{\prime}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle,$
+To estimate 2.2, we can write the last two terms of the sum as
+[…]
+[…]
 $2\omega_{k}\lambda_{k}\langle u_{0}-\hat{u},F^{\prime}_{i_{k}}(u^{\delta}_{k})^{*}(y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}))\rangle$ $\leq 2\lambda_{k}\omega_{k}L\|u_{0}-\hat{u}\|\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|$
 ((2.10)) $\leq\kappa^{-2}\lambda_{k}^{2}\omega_{k}^{2}L^{2}\|u_{0}-\hat{u}\|^{2}+\kappa^{2}\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2},$
-where we have used (3.2) and the inequality $2ab\leq a^{2}+b^{2}$ for $a,b\in\mathbb{R}$ in obtaining 3.2.
-By incorporating 3.2, 3.2, (3.3) and (3.2) in 3.2, we attain that
+where we have used (2.1) and the inequality $2ab\leq a^{2}+b^{2}$ for $a,b\in\mathbb{R}$ in obtaining 2.2.
+By incorporating (2.1), (2.2), 2.2, and 2.2 in 2.2, we attain that
 $\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u^{\delta}_{k}-\hat{u}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\|u_{0}-\hat{u}\|^{2}$
-$+2\lambda_{k}(1-\lambda_{k})\|u^{\delta}_{k}-\hat{u}\|\|u_{0}-\hat{u}\|$
-((3.12)) $+\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|\big{(}2\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
-$-\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\big{)}.$
+$\quad+2\lambda_{k}(1-\lambda_{k})\|u^{\delta}_{k}-\hat{u}\|\|u_{0}-\hat{u}\|$
+((2.11)) $\quad+\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|\big{(}2\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
+$\quad-\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\big{)}.$
 Next, according to the definition of $\hat{k}_{\delta}$ in (2.4), for $k<\hat{k}_{\delta}$ we further derive that
 ((2.12)) $\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|\leq\frac{1}{\tau}\left(\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}+\frac{M}{k+a}\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|\right).$
-Using the fact that $\omega_{k}\in[\omega,\Omega],u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ and $\hat{u}\in B_{\rho}(u_{0})$ along with inequality (3.13) in 3.2, we have
+[…]
 $\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq(1-\lambda_{k})^{2}c(\rho)^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\rho^{2}+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)$
-((3.14)) $-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-2\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
-$+\frac{2M\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|}{\tau(k+a)}.$
+((2.13)) $\quad-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-2\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
+$\quad+\frac{2M\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|}{\tau(k+a)}.$
 We note that
@@
 $c\bigg{(}\frac{1}{1+\varrho}-(1-\lambda_{k})^{2}\bigg{)}\geq\rho\lambda_{k}(1-\lambda_{k})+\sqrt{(1-\lambda_{k})^{2}+\bigg{(}\frac{1}{1+\varrho}-(1-\lambda_{k})^{2}\bigg{)}(1+\kappa^{-2}\Omega^{2}L^{2})},$
-which is true for $c(\rho)$. From (3.6) along with the inequality $2ab\leq a^{2}+b^{2},\forall\hskip 5.69054pta,b\in\mathbb{R}$ and 3.2, we get
-$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-2\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
-$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau}\left(\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}+\frac{1}{(k+a)^{2}}\right).$
-$\leq-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-(2+M)\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
-$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
-As $\omega_{k}\in[\omega,\Omega],0\leq\lambda_{k}\leq\lambda_{\text{max}}<1$ and $1-\lambda_{k}\leq 1$, we further attain
-$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
-$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\Omega(1+\eta)}{\tau(k+a)^{2}}.$
-Moreover, given the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$, we can write the last inequality as
-$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}|\mathcal{F}_{k}]$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\|y^{\delta}-F(u^{\delta}_{k})\|^{2}$
+which is true for $c(\rho)$.
+From (2.5) along with the inequality $ab\leq a^{2}+b^{2},\forall\hskip 5.69054pta,b\in\mathbb{R}, 2.2$, $\omega_{k}\in[\omega,\Omega],0\leq\lambda_{k}\leq\lambda_{\text{max}}<1$ and $1-\lambda_{k}\leq 1$, we get
+$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-(2+M)\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
+$\quad+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}$
+$\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
 $\hskip 28.45274pt+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\Omega(1+\eta)}{\tau(k+a)^{2}}.$
-Finally, by taking into consideration the entire conditional, we arrive at
-$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\mathbb{E}[\|y^{\delta}-F(u^{\delta}_{k})\|^{2}]$
-$\hskip 28.45274pt+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\Omega(1+\eta)}{\tau(k+a)^{2}}.$
-This with (3.7), (3.5) for $M=\frac{\varrho c(\rho)^{2}\tau}{\Omega(1+\varrho)(1+\eta)}$ and the inequality $k+a\geq 1$ implies that
-$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]\leq\frac{c(\rho)^{2}}{1+\varrho}+\frac{\varrho c(\rho)^{2}}{1+\varrho}=c(\rho)^{2}.$
-By incorporating Jensen’s inequality (2.2) with $f(x)=x^{2}$, we conclude that
-$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|]\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]^{1/2}\leq c(\rho).$
-This means that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})$. Further, by engaging the triangular inequality (2.1) and the fact that $\|u_{0}-\hat{u}\|\leq\rho$, we conclude that
-$\mathbb{E}[\|u^{\delta}_{k+1}-u_{0}\|]$ $\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|]+\mathbb{E}[\|\hat{u}-u_{0}\|]\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]^{1/2}+\|\hat{u}-u_{0}\|\leq\rho+c(\rho),$
-which implies that $u_{k+1}^{\delta}\in B_{\rho+c(\rho)}(u_{0})$.
-Next, we will prove that $\hat{k}_{\delta}$ is finite. To establish this, we can utilize 3.2 with (3.13) and (3.7) along with the fact that $1-\lambda_{k}\leq 1$ for $\lambda\in[0,\frac{1}{2})$ to get
-$\|u_{k+1}^{\delta}-\hat{u}\|^{2}+D\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\rho^{2}$
-$+2\lambda_{k}(1-\lambda_{k})\|u_{k}^{\delta}-\hat{u}\|\rho+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
-((3.15)) $\leq\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\rho^{2}+2\lambda_{k}\rho c(\rho)+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.\hskip 56.9055pt$
-Using the inequality that $\lambda_{k}^{2}\leq\lambda_{k}$ and the definition of the measurability of $u_{k}^{\delta}$ along with the additivity of expected values with respect to $\mathcal{F}_{k}$ in 3.2 to get that
-((3.16)) $\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}|\mathcal{F}_{k}]+D\|y^{\delta}-F(u_{k}^{\delta})\|^{2}$ $\leq\mathbb{E}[\|u_{k}^{\delta}-\hat{u}\|^{2}|\mathcal{F}_{k}]+\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}$
-$\hskip 56.9055pt+2\rho c(\rho)\lambda_{k}+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
-Taking the sum of (3.16) from $k=0$ to $k=\hat{k}_{\delta}-1$ to deduce that
-((3.17)) $\sum_{k=0}^{\hat{k}_{\delta}-1}\|y^{\delta}-F(u^{\delta}_{k})\|^{2}\leq D^{-1}\left(\rho^{2}+\left(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)\right)\sum_{k=0}^{\hat{k}_{\delta}-1}\lambda_{k}+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1-\lambda_{k}}{(k+a)^{2}}\right).$
-From (3.5) and inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})\hskip 5.69054pt\forall\hskip 5.69054pta,b\in\mathbb{R}$, we get
-((3.18)) $\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\leq\left(\|F(u_{k}^{\delta})-y^{\delta}\|+\frac{M}{k+a}\right)^{2}\leq 2\left(\|F(u_{k}^{\delta})-y^{\delta}\|^{2}+\frac{M^{2}}{(k+a)^{2}}\right).$
-By taking the sum of (3.18) from $k=0$ to $\hat{k}_{\delta}-1$, we get
-((3.19)) $\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}\leq 2\left(\sum_{k=0}^{\hat{k}_{\delta}-1}\|F(u_{k}^{\delta})-y^{\delta}\|^{2}+M^{2}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right).$
-Using (3.17) in (3.19) to get
-$\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}$ $\leq 2\left[D^{-1}\left\{\rho^{2}+\left(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)\right)\sum_{k=0}^{\hat{k}_{\delta}-1}\lambda_{k}\right.\right.$
-$\left.\left.\hskip 85.35826pt+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right\}+M^{2}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right].$
-By applying (3.8) and the result $\sum_{k=0}^{\infty}\frac{1}{(k+a)^{2}}\leq\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$ for $a\geq 1$, we further get
-$\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}<\infty.$
-Hence $\hat{k}_{\delta}$ is finite. Finally, from the definition of $\hat{k}_{\delta}$ in (3.5), we get
-$\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y^{\delta}-y^{\dagger}\|\to 0\text{\quad as\quad}\delta\to 0.$
-Therefore, we must have $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$. ∎
+This with (2.6), $M=\frac{\varrho c(\rho)^{2}\tau}{\Omega(1+\varrho)(1+\eta)}$ and the inequality $k+a\geq 1$ implies that
+$\|u^{\delta}_{k+1}-\hat{u}\|^{2}\leq\frac{c(\rho)^{2}}{1+\varrho}+\frac{\varrho c(\rho)^{2}}{1+\varrho}=c(\rho)^{2}.$
+[…]
+$\|u^{\delta}_{k+1}-u_{0}\|$ $\leq\|u^{\delta}_{k+1}-\hat{u}\|+\|\hat{u}-u_{0}\|\leq\rho+c(\rho),$
+which implies that $u_{k+1}^{\delta}\in B_{\rho+c(\rho)}(u_{0})$. The desired result is thus established. ∎
 Remark 2.5.
@@
 Therefore, a sufficiently small positive constant $\varrho$ is required in (2.5), due to the restriction that $\lambda_{k}\in[0,\frac{1}{2})$.
+We first show that Algorithm 2 is well-defined. Specifically, for
+$\|y_{i}^{\delta}-y_{i}^{\dagger}\|_{i\in\{0,1,\ldots,P-1\}}>0,$
+the modified discrepancy principle ((2.4)) yields a finite stopping index, ensuring that the algorithm terminates in finitely many steps.
+Moreover, we prove that $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$.
+Here, the expectation $\mathbb{E}$ and probability $\mathbb{P}$ are taken with respect to the filtration $\{\mathcal{F}_{k}\}_{k\geq 0}$, where $\mathcal{F}_{k}$ is the $\sigma$-algebra generated by the random indices $i_{n}$ for $0\leq n<k$.
+The formal result is stated below.
+Theorem 2.1.
+Under the assumptions of Lemma 2.1, the stopping index $\hat{k}_{\delta}$ as defined in (2.4) almost surely remains finite for $\|y_{i}^{\delta}-y_{i}^{\dagger}\|_{i\in\{0,1,\ldots,P-1\}}>0$ and $\hat{k}_{\delta}\rightarrow\infty$ as $\delta\rightarrow 0.$
+Proof.
+Our objective is to establish that
+$\mathbb{P}(\hat{k}_{\delta}=\infty)=0\quad\text{for }\|y_{i}^{\delta}-y_{i}^{\dagger}\|>0,\quad i=0,1,\ldots,P-1.$
+By the definition of $\hat{k}_{\delta}$, the event $\Psi:=\{\hat{k}_{\delta}=\infty\}$ can be equivalently expressed as
+$\Psi=\left\{\|F_{i}(u_{k}^{\delta})-y_{i}^{\delta}\|+\frac{M}{(k+a)}>\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|;i=0,1,\ldots P-1;\quad\text{for all integer }k\geq 0\right\}.$
+We now establish the desired result. Recall from (2.2), ((2.12)), and ((2.6)) that for $\lambda_{k}\in[0,\tfrac{1}{2})$ and $\omega_{k}\in[\omega,\Omega]$, we have
+$\|u_{k+1}^{\delta}-\hat{u}\|^{2}+D\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}\big{(}1+\kappa^{-2}\omega_{k}^{2}L^{2}\big{)}\rho^{2}$
+$\quad+2\lambda_{k}(1-\lambda_{k})\|u_{k}^{\delta}-\hat{u}\|\,\rho+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
+Since $(1-\lambda_{k})^{2}\leq 1$ and $\|u_{k}^{\delta}-\hat{u}\|\leq c(\rho)$ by Lemma 2.1, the above yields
+$\|u_{k+1}^{\delta}-\hat{u}\|^{2}+D\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}$ $\leq\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}\big{(}1+\kappa^{-2}\omega_{k}^{2}L^{2}\big{)}\rho^{2}$
+((2.14)) $\quad+2\lambda_{k}\rho\,c(\rho)+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
+Next, using $\lambda_{k}^{2}\leq\lambda_{k}$ and $\omega_{k}\leq\Omega$, and noting that $u_{k}^{\delta}$ is $\mathcal{F}_{k}$–measurable while $i_{k}$ is chosen uniformly from $\{0,1,\dots,P-1\}$ independent of $\mathcal{F}_{k}$, we take conditional expectation with respect to $\mathcal{F}_{k}$.
+By the independence of $i_{k}$, the expected residual term becomes the average over all $P$ components.
+Thus (2.2) implies
+$\mathbb{E}\big{[}\|u_{k+1}^{\delta}-\hat{u}\|^{2}\,\big{|}\,\mathcal{F}_{k}\big{]}+\frac{D}{P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u_{k}^{\delta})\|^{2}$
+((2.15)) $\quad\leq\mathbb{E}\big{[}\|u_{k}^{\delta}-\hat{u}\|^{2}\,\big{|}\,\mathcal{F}_{k}\big{]}+\rho^{2}\big{(}1+\kappa^{-2}\Omega^{2}L^{2}\big{)}\lambda_{k}+2\rho\,c(\rho)\,\lambda_{k}+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
+We now take full expectation in (2.2) and sum from $k=0$ to t1t-1.
+The left–hand side telescopes in the first term, while the residual terms remain additive.
+Using $\mathbb{E}[\|u_{t}^{\delta}-\hat{u}\|^{2}]\geq 0$, we arrive at
+((2.16)) $\begin{split}\sum_{k=0}^{t-1}\mathbb{E}\left[\frac{1}{P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u^{\delta}_{k})\|^{2}\right]&\leq\frac{1}{D}\bigg{(}\mathbb{E}[\|u_{0}^{\delta}-\hat{u}\|^{2}]\\ &\quad+\big{(}\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho\,c(\rho)\big{)}\sum_{k=0}^{t-1}\lambda_{k}\\ &\quad+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{t-1}\frac{1-\lambda_{k}}{(k+a)^{2}}\bigg{)}.\end{split}$
+Finally, recalling that the initial guess $u_{0}^{\delta}$ is deterministic (or $\mathcal{F}_{0}$–measurable) and satisfies $\|u_{0}^{\delta}-\hat{u}\|\leq\rho$.
+Hence $\mathbb{E}[\|u_{0}^{\delta}-\hat{u}\|^{2}]=\|u_{0}^{\delta}-\hat{u}\|^{2}\leq\rho^{2}$.
+So the bound ((2.16)) simplifies to
+((2.17)) $\begin{split}\sum_{k=0}^{t-1}\mathbb{E}\left[\frac{1}{P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u^{\delta}_{k})\|^{2}\right]&\leq\frac{1}{D}\bigg{(}\rho^{2}+\big{(}\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho\,c(\rho)\big{)}\sum_{k=0}^{t-1}\lambda_{k}\\ &\quad+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{t-1}\frac{1-\lambda_{k}}{(k+a)^{2}}\bigg{)}.\end{split}$
+By applying (2.7) and the result $\sum_{k=0}^{\infty}\frac{1}{(k+a)^{2}}\leq\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$ for $a\geq 1$ in ((2.17)), we further get
+((2.18)) $\sum_{k=0}^{t-1}\mathbb{E}\left[\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}\right]=\sum_{k=0}^{t-1}\mathbb{E}\left[\frac{1}{P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u^{\delta}_{k})\|^{2}\right]<\infty.$
+Now, define $\chi_{\Psi}$ as the characteristic function associated with the set $\Psi$, such that
+$\chi_{\Psi}(w)=\begin{cases}1,&\text{if }w\in\Psi,\\ 0,&\text{if }w\notin\Psi.\end{cases}$
+By definition of $\Psi$, for every integer $k\geq 0$ we have, for all $i=0,1,\dots,P-1$,
+$\|y_{i}^{\delta}-F_{i}(u_{k}^{\delta})\|>\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|-\frac{M}{k+a}.$
+Using the elementary inequality $(l-m)^{2}\geq\tfrac{1}{2}l^{2}-m^{2}$ with $l=\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|$ and $m=\dfrac{M}{k+a}$, we obtain for each $i$ on $\Psi$
+$\|y_{i}^{\delta}-F_{i}(u_{k}^{\delta})\|^{2}\geq\tfrac{1}{2}\tau^{2}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}-\frac{M^{2}}{(k+a)^{2}}.$
+Multiplying by the indicator $\chi_{\Psi}$ and taking expectation, then averaging over the uniformly sampled index $i_{k}$, gives
+$\begin{split}\mathbb{E}\big{[}\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}\big{]}&\geq\mathbb{E}\big{[}\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}\chi_{\Psi}\big{]}\\ &=\mathbb{E}\Big{[}\frac{1}{P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u_{k}^{\delta})\|^{2}\chi_{\Psi}\Big{]}\\ &\geq\frac{\tau^{2}}{2P}\,\mathbb{E}\Big{[}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}\chi_{\Psi}\Big{]}-\frac{M^{2}}{(k+a)^{2}}\,\mathbb{E}[\chi_{\Psi}]\\ &\geq\frac{\tau^{2}}{2P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}\,\mathbb{P}(\Psi)-\frac{M^{2}}{(k+a)^{2}}\,\mathbb{P}(\Psi).\end{split}$
+Sum the last inequality over $k=0,\dots,t-1$ and use linearity of expectation, we arrive at
+((2.19)) $\sum_{k=0}^{t-1}\mathbb{E}\big{[}\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}\big{]}\geq\frac{\tau^{2}}{2P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}t\,\mathbb{P}(\Psi)-M^{2}\,\mathbb{P}(\Psi)\sum_{k=0}^{t-1}\frac{1}{(k+a)^{2}}.$
+By ((2.18)) there exists a constant $\mathfrak{C}>0$ independent of $t$ such that
+$\sum_{k=0}^{t-1}\mathbb{E}\big{[}\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}\big{]}\leq\mathfrak{C}.$
+Combining this with ((2.19)) gives
+$\mathfrak{C}\geq\frac{\tau^{2}}{2P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}\,t\,\mathbb{P}(\Psi)-M^{2}\,\mathbb{P}(\Psi)\sum_{k=0}^{t-1}\frac{1}{(k+a)^{2}}.$
+Rearrange to obtain
+$\mathbb{P}(\Psi)\left(\frac{\tau^{2}}{2P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}\,t-M^{2}\sum_{k=0}^{t-1}\frac{1}{(k+a)^{2}}\right)\leq\mathfrak{C}.$
+The series $\sum_{k=0}^{\infty}\frac{1}{(k+a)^{2}}\leq\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$, hence
+$\mathbb{P}(\Psi)\left(\frac{\tau^{2}}{2P}\sum_{i=0}^{P-1}\|y_{i}^{\delta}-y_{i}^{\dagger}\|^{2}\,t-\frac{M^{2}\pi^{2}}{6}\right)\leq\mathfrak{C}.$
+Letting $t\to\infty$, we observe that the term in parentheses on the left–hand side grows without bound whenever $\mathbb{P}(\Psi)>0$.
+This would contradict the boundedness of the left–hand side unless $\mathbb{P}(\Psi)=0$.
+Therefore, we conclude that
+$\mathbb{P}(\Psi)=0,\quad\text{i.e.,}\quad\mathbb{P}(\hat{k}_{\delta}=\infty)=0.$
+This implies that the complement of the event $\Psi$ occurs with probability one, i.e., along almost every sample path, ((2.4)) holds for some finite $\hat{k}_{\delta}$.
+Finally, from the definition of $\hat{k}_{\delta}$ in (2.4), we get
+$\frac{M}{(\hat{k}_{\delta}+a)}\leq\tau\|y_{i}^{\delta}-y_{i}^{\dagger}\|\leq\tau\|y^{\delta}-y^{\dagger}\|\to 0\text{\quad as \quad}\delta\to 0.$
+Consequently, we must have $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$. This completes the proof. ∎
+The proof of the following lemma draws inspiration from the methodologies outlined in [36] and [32].
+In this result, we examine the asymptotic behavior of the stopping index prescribed by Rule 1 when applied to noisy data satisfying Assumption 2.2.
+It is noteworthy that the proof of this lemma relies essentially on Theorem 2.1.
 Lemma 2.2.
-Let Assumption 3.1 hold and let the sequence $\{\lambda_{k}\}$ be chosen such that (3.7) and (3.8) hold.
-Additionally, let $\{y^{\delta}\}$ represent a family of noisy data that fulfill Assumption 3.2.
-Furthermore, assume that $k_{*}:=k_{*}(y^{\delta})$ is the integer determined by Rule 1.
+Let the sequence $\{\lambda_{k}\}$, $\omega$ and $\Omega$ be chosen such that (2.6) and (2.7) hold and let Assumption 2.1 hold.
+Additionally, let $\{y^{\delta}\}$ represent a set of noisy data that fulfill Assumption 2.2.
+Furthermore, assume that $k_{*}:=k_{*}(y^{\delta})$ is the stopping index as defined in Rule 1.
 Then
-$\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0\quad\text{as}\quad\delta\to 0.$
-Consequently, we have
-$\|F(u^{\delta}_{k_{*}(y^{\delta})})-y^{\delta}\|\to 0\quad\text{and}\quad k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\to 0\quad\text{as}\quad\delta\to 0.$
+$\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\to 0\quad\text{as}\quad\delta\to 0.$
+Thus, we have
+$\mathbb{E}\left[\|F(u^{\delta}_{k_{*}(y^{\delta})})-y^{\delta}\|\right]\to 0\quad\text{and}\quad\mathbb{E}\left[k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\right]\to 0\quad\text{as}\quad\delta\to 0.$
 Proof.
 By employing the condition (2.7) and the fact that $\sum_{k=0}^{\infty}\frac{1}{(k+a)^{2}}\leq\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$ for $a\geq 1$ in (2.17), we arrive at
-((3.20)) $\sum_{k=0}^{\hat{k}_{\delta}-1}\|y^{\delta}-F(u^{\delta}_{k})\|^{2}\leq D^{-1}C_{1},$
+((2.20)) $\sum_{k=0}^{\hat{k}_{\delta}-1}\mathbb{E}\left[\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u^{\delta}_{k})\|^{2}\right]\leq PD^{-1}C_{1},$
 where $C_{1}=\left(\rho^{2}+\left(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)\right)\lambda+\frac{M\Omega\pi^{2}(1+\eta)}{6\tau}\right)$.
 By minimality of $\Psi(k_{*},y^{\delta})$, it follows from Rule 1 that
-$\|F(u_{k}^{\delta})-y^{\delta}\|^{2}=\frac{\Psi(k,y^{\delta})}{k+a}\geq\frac{\Psi(k_{*}(y^{\delta}),y^{\delta})}{k+a}.$
+$\sum_{i=0}^{P-1}\|F_{i}(u_{k}^{\delta})-y_{i}^{\delta}\|^{2}=\frac{\Psi(k,y^{\delta})}{k+a}\geq\frac{\Psi(k_{*}(y^{\delta}),y^{\delta})}{k+a}.$
 This and (2.20) further yields
-((3.21)) $\Psi(k_{*}(y^{\delta}),y^{\delta})\sum^{\hat{k}_{\delta}-1}_{k=0}\frac{1}{k+a}\leq\sum_{k=0}^{\hat{k}_{\delta}-1}\|F(u^{\delta}_{k})-y^{\delta}\|^{2}\leq D^{-1}C_{1}.$
+((2.21)) $\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\sum^{\hat{k}_{\delta}-1}_{k=0}\frac{1}{k+a}\leq\sum_{k=0}^{\hat{k}_{\delta}-1}\mathbb{E}\left[\sum_{i=0}^{P-1}\|y_{i}^{\delta}-F_{i}(u^{\delta}_{k})\|^{2}\right]\leq PD^{-1}C_{1}.$
 Note that
-[…]
+$\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{k+a}\geq\sum_{k=0}^{\hat{k}_{\delta}-1}\int_{k}^{k+1}\frac{1}{t+a}dt=\int_{0}^{\hat{k}_{\delta}}\frac{1}{t+a}dt=\log\frac{\hat{k}_{\delta}+a}{a}.$
 Plugging the last estimate in (2.21) yields
-$\bigg{(}\log\frac{\hat{k}_{\delta}+a}{a}\bigg{)}\Psi(k_{*}(y^{\delta}),y^{\delta})\leq D^{-1}C_{1}.$
-Since $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$, we must have $\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0$ as $\delta\to 0$. Further, we note that
-$a\|F(u^{\delta}_{k^{*}(y^{\delta})})-y^{\delta}\|^{2}\leq\Psi(k_{*}(y^{\delta}),y^{\delta}).$
+$\bigg{(}\log\frac{\hat{k}_{\delta}+a}{a}\bigg{)}\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\leq PD^{-1}C_{1}.$
+Since $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$, we must have $\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\to 0$ as $\delta\to 0$. Further, we note that
+$a\sum_{i=0}^{P-1}\|F_{i}(u^{\delta}_{k_{*}(y^{\delta})})-y_{i}^{\delta}\|^{2}=a\|F(u^{\delta}_{k_{*}(y^{\delta})})-y^{\delta}\|^{2}\leq\Psi(k_{*}(y^{\delta}),y^{\delta}).$
 Also, Assumption 2.2 implies that
 $(k_{*}(y^{\delta})+a)\varkappa^{2}\|y^{\delta}-y^{\dagger}\|^{2}\leq\Psi(k_{*}(y^{\delta}),y^{\delta}).$
-Hence, we conclude that $\|F(u_{k_{*}(y^{\delta})}^{\delta})-y^{\delta}\|\to 0$ and $k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\to 0$ as $\delta\to 0,$ since $\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0$ as $\delta\to 0$.
+Hence, we conclude that $\mathbb{E}\left[\|F(u_{k_{*}(y^{\delta})}^{\delta})-y^{\delta}\|\right]\to 0$ and $\mathbb{E}\left[k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\right]\to 0$ as $\delta\to 0,$ since $\mathbb{E}\left[\Psi(k_{*}(y^{\delta}),y^{\delta})\right]\to 0$ as $\delta\to 0$.
 Thus, result.
 ∎
-The next result demonstrates that IRSGD iterates $u_{k}^{\delta}$ remains in a ball around $u_{0}$ for any $0\leq k<k_{*}$, where $k_{*}$ is defined via Rule 1 (note that (b) of Lemma 3.1 shows the same for $0\leq k<\hat{k_{\delta}}$).
+The next result demonstrates that the iterates $u_{k}^{\delta}$ generated by Algorithm 1 remains in a ball around $u_{0}$ for any $0\leq k<k_{*}$, where $k_{*}$ is defined via Rule 1 (note that Lemma 2.1 shows the same for $0\leq k<\hat{k_{\delta}}$).
 Lemma 2.3.
-Let all the assumptions of Lemma 3.2 hold. Additionally, assume that
-((3.22)) $D_{1}:=\kappa^{2}-\Omega(1-\lambda_{k})(1-3\eta)+\omega^{2}L^{2}\geq 0.$
-Then, if $u^{\delta}_{k}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<k_{*}$, it follows that $u^{\delta}_{k+1}\in B_{c(\rho)}(\hat{u})\subset B_{\rho+c(\rho)}(u_{0})$, provided $\delta$ is sufficiently small.
+Let all the conditions of Lemma 2.2 be satisfied. Additionally, assume that
+((2.22)) $D_{1}:=\kappa^{2}-\Omega(1-\lambda_{\text{max}})(1-3\eta)+\omega^{2}L^{2}>0.$
+[…]
 Proof.
-For $k=0$, result is trivial.
-By assuming $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<k^{*},$ we show that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})$.
-From 3.2 and the inequalities $\|u_{k}^{\delta}-\hat{u}\|\leq c(\rho)$ and $\|u_{0}-\hat{u}\|\leq\rho$, it follows that
+[…]
 $\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}$ $-\|u_{k}^{\delta}\ -\hat{u}\|^{2}$
 ((2.23)) $\leq\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}^{2}+2\omega_{k}(1-\lambda_{k})(1+\eta)\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
-$+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}$
-$\leq\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}^{2}+\omega_{k}(1-\lambda_{k})(1+\eta)(\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}+\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|^{2})$
-$+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}$
+$\quad+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}$
 $\leq\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}^{2}+\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|^{2}+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)$
-$+(\omega_{k}(1-\lambda_{k})(1+\eta)-\left(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\right)\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}.$
-Using the fact that $1-\lambda_{k}\leq 1$ and $\lambda_{k}^{2}\leq\lambda_{k}$ for $\lambda\in[0,\frac{1}{2})$, we get
+$\quad+(\omega_{k}(1-\lambda_{k})(1+\eta)-\left(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\right)\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}.$
+Using the fact that $1-\lambda_{k}\leq 1$ and $\lambda_{k}^{2}\leq\lambda_{k}$ for $\lambda_{k}\in[0,\frac{1}{2})$, we get
 ((2.24)) $\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}$ $-\|u_{k}^{\delta}\ -\hat{u}\|^{2}\leq(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda_{k}+\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|^{2}$
 $+(\omega_{k}(1-\lambda_{k})(1+\eta)-\left(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\right)\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}.$
-Further, the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$ implies that
-$\mathbb{E}[\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}-\|u_{k}^{\delta}\ -\hat{u}\|^{2}|\mathcal{F}_{k}]$ $\leq(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda_{k}+\Omega(1-\lambda_{k})(1+\eta)\|y^{\dagger}-y^{\delta}\|^{2}$
-$\hskip 28.45274pt+\left(\Omega(1-\lambda_{k})(1-3\eta)-\omega^{2}L^{2}-\kappa^{2})\right)\|F(u_{k}^{\delta})-y^{\delta}\|^{2}.$
-By considering the conditional in its whole and utilizing $1-\lambda_{k}\leq 1$, we further reach at
+By taking the total conditional and utilizing $1-\lambda_{k}\leq 1$, we further reach at
 $\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}]-\mathbb{E}[\|u_{k}^{\delta}-\hat{u}\|^{2}]$ $\leq(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda_{k}+\Omega(1+\eta)\mathbb{E}[\|y^{\dagger}-y^{\delta}\|^{2}]$
@@
 and hence by using (2.7) and the fact that $k<k_{*}$, where $k_{*}$ is a finite integer, we get
-$\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\Omega(1+\eta)k_{*}\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}].$
+$\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\Omega(1+\eta)\mathbb{E}[k_{*}\|y^{\delta}-y^{\dagger}\|^{2}].$
 From Lemma 2.2, one can note that for sufficiently small $\delta$, there exist a constant $\xi$ such that
-((3.26)) $\Omega(1+\eta)k_{*}(y^{\delta})\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}]\leq\xi,\ \text{with}\ \rho^{2}+\ (\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\xi\leq c(\rho)^{2}.$
-Hence as a result, we get
-$\mathbb{E}[\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}]\leq c(\rho)^{2},$
-which completes the proof. ∎
+((2.26)) $\Omega(1+\eta)\mathbb{E}[k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}]\leq\xi,\ \text{with}\ \rho^{2}+\ (\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\xi\leq c(\rho)^{2}.$
+Hence, as a result, we get $\mathbb{E}[\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}]\leq c(\rho)^{2}.$ By incorporating Jensen’s inequality with $f(x)=x^{2}$, we conclude that
+$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|]\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]^{1/2}\leq c(\rho).$
+This concludes the proof. ∎
+3. Convergence Analysis
+In this section, we prove that $u^{\delta}_{k_{*}}$ converges almost surely to a solution of (1.1) as $\delta\to 0$.
+We first analyze the exact-data analogue of Algorithm 1 and establish its almost-sure convergence.
+Next, we show a stability property along each sample path that connects the noisy- and exact-data cases.
+These results together yield the main convergence theorem.
+The exact-data version of Algorithm 1 is given below.
+Algorithm 3 Exact-data IRSGD
+1:Given: $F_{i},y_{i}^{\dagger}$ for all $i\in\{0,1,\ldots,P-1\}$
+2:Initialize: $u_{0}=u^{(0)}$
+3:for $k\geq 0$ do
+4:  Choose $\omega_{k}\in[\omega,\Omega]$ and the weight parameter $\lambda_{k}$ such that
+$0\leq\lambda_{k}\leq\lambda_{\text{max}}<\frac{1}{2}.$
+5:  Randomly pick an index $i_{k}$ from $\{0,1,\ldots,P-1\}$ uniformly.
+6:  Update
+$u_{k+1}=u_{k}-\omega_{k}F^{\prime}_{i_{k}}(u_{k})^{*}(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger})-\lambda_{k}(u_{k}-u^{(0)}),$
+where $F^{\prime}_{i_{k}}$ is the Fréchet derivative of $F_{i_{k}}$, and $u^{(0)}$ is the initial guess.
+7:end for
+[…]
+((3.1)) $\daleth=\left\{\sum_{l=0}^{\infty}\sum_{i=0}^{P-1}\|F_{i}(u_{l})-y_{i}^{\dagger}\|^{2}<\infty\right\},$
+and show that this event occurs almost surely, i.e., $\mathbb{P}(\daleth)=1$.
 Corollary 3.1.
-Let the assumptions of Lemma 3.1 be fulfilled. Moreover assume that IRSGD method (3.1) is stopped according to Rule 1, then for sufficiently small $\delta$, we have
-$\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq D_{1}^{-1}\left((\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{k}+\rho^{2}+\xi\right)<\infty.$
-In particular, for the case $y^{\delta}=y^{\dagger}$ ((((i.e., if $\delta=0)$, we have
-$\sum_{l=0}^{\infty}\mathbb{E}[\|F(u_{l})-y^{\dagger}\|^{2}]\leq D_{1}^{-1}\left(\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{\infty}\lambda_{k}\right)<\infty,$
-where $D_{1}$ is defined in (3.22).
+Let the assumptions of Lemma 2.1 be satisfied. Additionally, assume that IRSGD method (1.11) is stopped according to Rule 1, then for sufficiently small $\delta$, we have
+$\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq D_{1}^{-1}\left((\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{k}+\rho^{2}+\xi\right)<\infty,$
+where $D_{1}$ is defined in (2.22). In particular, for the case $y^{\delta}=y^{\dagger}$, we have
+$\sum_{l=0}^{\infty}\mathbb{E}[\|F(u_{l})-y^{\dagger}\|^{2}]\leq D_{1}^{-1}\left(\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{\infty}\lambda_{k}\right)<\infty.$
+[…]
 Proof.
 From (2.25), we get
-((3.27)) $D_{1}\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{l}+\Omega(1+\eta)k_{*}\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}].$
+$D_{1}\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{l}+\Omega(1+\eta)\mathbb{E}[k_{*}\|y^{\delta}-y^{\dagger}\|^{2}].$
 Using (2.26), we note that
-((3.28)) $\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq D_{1}^{-1}\left((\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{k}+\rho^{2}+\xi\right)<\infty.$
-Next, for the case $\delta=0$, it follows directly from (3.27) that
-((3.29)) $\sum_{l=0}^{\infty}\mathbb{E}[\|F(u_{l})-y^{\dagger}\|^{2}]\leq D_{1}^{-1}\left(\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{\infty}\lambda_{k}\right)<\infty.$
-Here we have used the fact that $\sum_{l=0}^{\infty}\lambda_{k}<\infty.$ Thus, (3.28) and (3.29) complete the proof. ∎
+$\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq D_{1}^{-1}\left((\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{k}+\rho^{2}+\xi\right)<\infty.$
+Thus, result. The case $y^{\delta}=y^{\dagger}$ can be handled similarly. ∎
 Remark 3.1.
 If $\lambda_{\text{max}}=0$, the parameter $\kappa$ in conditions (2.6) and (2.22) can be set to zero.
-This adjustment mirrors the conditions required to establish the same results for the stochastic gradient descent (SGD) method (1.9).
-3.3. Convergence for non noisy case
-The following result illustrate that the sequence of IRSGD iterates $\{u_{k}\}_{k\geq 1}$ in (3.1) with exact data $y$ forms a Cauchy sequence.
+This adjustment mirrors the conditions required to establish the same results for SGD method (1.9).
+3.1. Convergence for exact-data case
+In this subsection, we establish the convergence of Algorithm 3. Let us assume that
+((3.2)) $D_{2}:=(2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})>0.$
 Lemma 3.1.
-Let all the assumptions of Lemma 3.1 hold. Additionally assume that
-((3.30)) $D_{2}:=(2\omega(1-\lambda_{k})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})\geq 0.$
-Then the IRSGD method (3.1) for $\delta=0$ produces a sequence of iterates $\{u_{k}\}_{k\geq 1}$ which is a Cauchy sequence.
+Suppose all the conditions of Lemma 2.1 are satisfied.
+Additionally, assume that ((3.2)) holds.
+Then, the sequence of iterates $\{u_{k}\}_{k\geq 1}$ produced by Algorithm 3 is a Cauchy sequence almost surely.
 Proof.
-For a solution $u^{\dagger}$ of (1.2), we define $e_{k}:=u_{k}-u^{\dagger}$.
-As indicated in part (b) of Lemma 3.1, $\mathbb{E}[\|e_{k}\|^{2}]$ is a bounded sequence and hence it has a convergent subsequence $\mathbb{E}[\|e_{k_{n}}\|^{2}]$.
+Firstly, we define $e_{k}:=u_{k}-u^{\dagger}$, for a solution $u^{\dagger}$ of (1.1).
+As shown in Lemma 2.1, $\mathbb{E}[\|e_{k}\|^{2}]$ is a bounded sequence and hence it has a convergent subsequence $\mathbb{E}[\|e_{k_{n}}\|^{2}]$.
 Let it converge to some $\epsilon\geq 0$.
-It follows directly from 3.2 that, for $\delta=0$ we have
+It follows directly from 2.2 that, for $\delta=0$ we have
 $\|u_{k+1}-u^{\dagger}\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u_{k}-u^{\dagger}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
@@
 $-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|y_{i_{k}}-F_{i_{k}}(u_{k})\|^{2}.$
-From the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$, we deduce that
+From the definition of $F$ and the measurability of $u_{k}^{\delta}$ concerning $\mathcal{F}_{k}$, we deduce that
 $\mathbb{E}[\|u_{k+1}-u^{\dagger}\|^{2}|\mathcal{F}_{k}]$ $\leq\mathbb{E}[(1-\lambda_{k})^{2}\|u_{k}-u^{\dagger}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
-$+2\lambda_{k}(1-\lambda_{k})\|u_{k}-u^{\dagger}\|\|u_{0}-u^{\dagger}\||\mathcal{F}_{k}]$
-$-\|y-F(u_{k})\|^{2}(2\omega(1-\lambda_{k})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2}).$
-By using (3.30) and considering the full conditional, we reach at
-$\mathbb{E}[\|u_{k+1}-u^{\dagger}\|^{2}]+D_{2}\mathbb{E}[\|y-F(u_{k})\|^{2}]$ $\leq(1-\lambda_{k})^{2}\mathbb{E}[\|u_{k}-u^{\dagger}\|^{2}]+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
+$\quad+2\lambda_{k}(1-\lambda_{k})\|u_{k}-u^{\dagger}\|\|u_{0}-u^{\dagger}\||\mathcal{F}_{k}]$
+$\quad-\frac{1}{P}\|y-F(u_{k})\|^{2}(2\omega(1-\lambda_{k})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2}).$
+Taking the total conditional and using (3.2), we get
+$\mathbb{E}[\|u_{k+1}-u^{\dagger}\|^{2}]+\frac{D_{2}}{P}\mathbb{E}[\|y-F(u_{k})\|^{2}]$ $\leq(1-\lambda_{k})^{2}\mathbb{E}[\|u_{k}-u^{\dagger}\|^{2}]+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
 $+2\lambda_{k}(1-\lambda_{k})\mathbb{E}[\|u_{k}-u^{\dagger}\|]\|u_{0}-u^{\dagger}\|.$
@@
 ((3.3)) $\mathbb{E}[\|e_{k}\|^{2}]^{1/2}\leq\mathbb{E}[\|e_{m}\|^{2}]^{1/2}\prod_{l=m}^{k-1}(1-\lambda_{l})+\rho\sqrt{1+\kappa^{-2}\Omega^{2}L^{2}}(1-\prod_{l=m}^{k-1}(1-\lambda_{l}))$
-for $m<k,$ where we have used the formula
+for $m<k,$ where we have used the following result
 $1-\prod_{l=m}^{n}(1-\lambda_{l})=\sum_{j=m}^{n}\lambda_{j}\prod_{l=j+1}^{n}(1-\lambda_{l})$
@@
 $\mathbb{E}[\|e_{k}\|^{2}]^{1/2}$ $\geq\mathbb{E}[\|e_{k_{n}}\|^{2}]^{1/2}-(\rho\sqrt{1+\kappa^{-2}\Omega^{2}L^{2}}-\mathbb{E}[\|e_{k_{n}}\|^{2}]^{1/2})\left(1-\prod_{l=k}^{k_{n-1}}(1-\lambda_{l})\right).$
-These estimates, along with the convergence of $\mathbb{E}[\|e_{k_{n}}\|^{2}]$, immediately imply
+These bounds, along with the convergence of $\mathbb{E}[\|e_{k_{n}}\|^{2}]$, directly imply
 $\lim_{k\to\infty}\mathbb{E}[\|e_{k}\|^{2}]^{1/2}=\epsilon^{1/2},$
@@
 ((3.4)) $\prod_{l=m}^{k-1}(1-\lambda_{l})\to 1\text{ as }m\to\infty,\;k>m.$
-Further, to show the convergence of the iterates $u_{k}$, it is crucial to note that $\mathbb{E}[\langle\cdot,\cdot\rangle]$ serves as an inner product.
-For any $n\geq k$, we select an index $m$ satisfying $n\geq m\geq k$ such that
-((3.33)) $\mathbb{E}[\|y-F(u_{m})\|^{2}]\leq\mathbb{E}[\|y-F(u_{i})\|^{2}]\quad\forall k\leq i\leq n.$
-Using inequality $\mathbb{E}[\|e_{n}-e_{k}\|^{2}]^{1/2}\leq\mathbb{E}[\|e_{n}-e_{m}\|^{2}]^{1/2}+\mathbb{E}[\|e_{m}-e_{k}\|^{2}]|^{1/2}$ and the identities
-$\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{j},e_{m}\rangle]+\mathbb{E}[\|e_{n}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
-$\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]+\mathbb{E}[\|e_{k}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
-it is sufficient to establish that $\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ and $\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ tend to zero as $k\rightarrow\infty$.
-As $k\rightarrow\infty$, the last two terms on the right-hand sides of the preceding two identities tend to $\epsilon-\epsilon=0$.
-This is due to the convergence of $\mathbb{E}[\|e_{k}\|^{2}]$ to $\epsilon$, as discussed above.
-Next, we proceed to demonstrate that the term $\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]$ also approaches to zero as $k\rightarrow\infty$.
-In fact, by virtue of the definition of $u_{k}$, we have the following representation
-$e_{k}=e_{m}\prod_{l=m}^{k-1}(1-\lambda_{l})+\sum_{j=m}^{k-1}\omega_{j}F_{i_{j}}^{\prime}(u_{j})^{\ast}(y_{i_{j}}-F_{i_{j}}(u_{j}))\prod_{l=j+1}^{k-1}(1-\lambda_{l})+(u_{0}-u^{\dagger})\prod_{l=m}^{k-1}(1-\lambda_{l}).$
-for $m<k$. Using this formula, we obtain the estimate
+Note that the operator $\mathbb{E}[\langle\cdot,\cdot\rangle]$ defines an inner product. Now, for any $n\geq k$, we pick an index $m$ such that $k\leq m\leq n$ and
+((3.5)) $\mathbb{E}[\|y-F(u_{m})\|^{2}]\leq\mathbb{E}[\|y-F(u_{i})\|^{2}]\quad\text{for all }k\leq i\leq n.$
+Using the triangular inequality
+((3.6)) $\mathbb{E}[\|e_{n}-e_{k}\|^{2}]^{1/2}\leq\mathbb{E}[\|e_{n}-e_{m}\|^{2}]^{1/2}+\mathbb{E}[\|e_{m}-e_{k}\|^{2}]^{1/2},$
+along with the identities
+((3.7)) $\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]+\mathbb{E}[\|e_{n}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
+((3.8)) $\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]+\mathbb{E}[\|e_{k}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
+we aim to prove the desired result. To that end, it suffices to show that both $\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ and $\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ converge to zero as $k\to\infty$.
+Since $\mathbb{E}[\|e_{k}\|^{2}]\to\epsilon$ as $k\to\infty$, the final two terms on the right-hand sides of the above identities vanish in the limit, i.e., $\epsilon-\epsilon=0$.
+Thus, it remains to show that $\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]$ also tends to zero as $k\to\infty$.
+The same argument applies to $\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]$.
+To show this, we recall the recursive structure of $u_{k}$, from which we obtain
+$e_{k}=e_{m}\prod_{l=m}^{k-1}(1-\lambda_{l})+\sum_{j=m}^{k-1}\omega_{j}F_{i_{j}}^{\prime}(u_{j})^{\ast}(y_{i_{j}}-F_{i_{j}}(u_{j}))\prod_{l=j+1}^{k-1}(1-\lambda_{l})+(u_{0}-u^{\dagger})\prod_{l=m}^{k-1}(1-\lambda_{l}),$
+for $m<k$. This formula allows us to determine that
 $|\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]|$ $\leq\left(1-\prod_{l=m}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$
-$+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y_{i_{j}}-F_{i_{j}}(u_{j}),F_{i_{j}}^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle],$
+$\quad+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y_{i_{j}}-F_{i_{j}}(u_{j}),F_{i_{j}}^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle],$
 $|\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]|$ $\leq\left(1-\prod_{l=m}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$
-$+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y-F(u_{j}),F^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle]|.$
-Since $|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$ is bounded, the first term on the right-hand side tends to zero due to (3.32).
-Since $1-\lambda_{l}\leq 1$, the second term on the right-hand side can be estimated by
+$\quad+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y-F(u_{j}),F^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle]|.$
+Due to (3.4), the first term on the right-hand side approaches zero because $|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$ is bounded.
+In addition to that, by using the inequality $1-\lambda_{k}\leq 1$ along with (2.7), one can estimate the second term on the right-hand side as
 ((3.9)) $\Omega(1+3\eta)\sum_{j=m}^{n-1}\mathbb{E}[\|y-F(u_{j})\|^{2}].$
-Due to (3.8) and Corollary 3.1, it follows that (3.34) tends to zero as $m\to\infty$.
+Now from Corollary 3.1, it follows that (3.9) approaches zero as $k\to\infty$.
 This means $\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]\to 0$.
 Analogously, we have $\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]\to 0$.
-Thus, result holds.
-∎
-In the following theorem, we establish the convergence of non-noisy iterates.
+Hence from ((3.7)) and ((3.8)) we have
+$\mathbb{E}[\|e_{n}-e_{m}\|^{2}]\to 0\quad\text{and}\quad\mathbb{E}[\|e_{m}-e_{k}\|^{2}]\to 0$
+as $k\to\infty.$ Hence, the desired result follows directly from inequality ((3.6)). ∎
+[…]
 Theorem 3.1.
-Suppose that all the assumptions of Lemma 3.1 and (3.30) hold.
-Then, the IRSGD method (3.1) corresponding to $\delta=0$ generates a sequence of iterates $\{u_{k}\}_{k\geq 1}$ that converges to an element of $\mathcal{D}_{\rho+c(\rho)}(u^{\dagger})$ in $U$, i.e.,
-$\lim_{k\to 0}\mathbb{E}[\|u_{k}-\hat{u}\|^{2}]=0,\quad\hat{u}\in\mathcal{D}_{\rho+c(\rho)}(u^{\dagger}).$
+[…]
+[…]
 Proof.
-The proof is similar to [2, Theorem 1]. ∎
-3.4. Convergence for noisy case
-The next result demonstrates the pathwise stability of the IRSGD iterates $u_{k}^{\delta}$ as $\delta\to 0$.
-Specifically, for any fixed integer $k\geq 0$ and any sample path $(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}$, it holds that $\|u_{k}^{\delta}-u_{k}\|^{2}\to 0$ as $\delta\to 0$.
-As a result, the expectation $\mathbb{E}[\|u_{k}^{\delta}-u_{k}\|^{2}]\to 0$ as $\delta\to 0$, confirming the pathwise stability of the iterates for each realization within the filtration $\mathcal{F}_{k}$.
+Lemma 3.1 establishes that the sequence of iterates $\{u_{k}\}_{k\geq 1}$ produced by Algorithm 3 forms a Cauchy sequence in $U$.
+Consequently, the sequence admits a limit, denoted by $u^{\dagger}\in U$.
+Since the operator $F$ is continuous, it follows that $F(u_{k})\to F(u^{\dagger})$ as $k\to\infty$.
+Moreover, Corollary 3.1 guarantees that
+$\lim_{k\to\infty}\mathbb{E}\left[\|F(u_{k})-y^{\dagger}\|^{2}\right]=0,$
+[…]
+3.2. Regularization Quality
+The primary goal of this subsection is to demonstrate that the IRSGD method possesses the regularizing property when employed with the proposed heuristic stopping rule.
+Our approach begins by establishing the pathwise stability of the IRSGD iterates $u_{k}^{\delta}$ as $\delta\to 0$, thereby linking Algorithm 1 for noisy data to its exact-data counterpart in Algorithm 3.
+The proof structure is adapted from the methodology outlined in [16, Lemma 3.6].
 Lemma 3.2.
-Let all the assumptions of Lemma 3.1 be satisfied.
-For any fixed $k\in\mathbb{N}$ and any path $(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}$, let $u_{k}$ and $u_{k}^{\delta}$ be the IRSGD iterates along the path for exact data $y$ and noisy data $y^{\delta}$, respectively.
+Assume that every condition in Lemma 2.1 is met.
+For any given $k\in\mathbb{N}$ and any realization of the sequence $(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}$, let $u_{k}$ and $u_{k}^{\delta}$ be the iterates produced by the IRSGD method (1.11) for the exact and noisy data, respectively.
 Then
@@
 For $k=0$ it is trivial.
-Suppose that it holds for all indices up to $k$ and any path in $\mathcal{F}_{k}$.
-From (3.1), for any fixed path $(i_{1},\ldots,i_{k})$, we have
+Assume that it is true for any path in $\mathcal{F}_{k}$ and for all indices up to $k$.
+According to (1.11), for any fixed path $(i_{1},\ldots,i_{k})$, we get
 $u_{k+1}^{\delta}-u_{k+1}$ $=(1-\lambda_{k})(u_{k}^{\delta}-u_{k})-\omega_{k}((F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}-F_{i_{k}}^{\prime}(u_{k})^{*})(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})$
-$+F_{i_{k}}^{\prime}(u_{k})^{*}((F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}))).$
+$\quad+F_{i_{k}}^{\prime}(u_{k})^{*}((F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}))).$
 Thus, by the triangle inequality, we get
@@
 ((3.10)) $+\omega_{k}\|F_{i_{k}}^{\prime}(u_{k})^{*}\|\|(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger})\|.$
-Next, we claim that for any specific value of $k$, (i) $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}\|$ is bounded, and (ii) $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}^{\delta}-u_{k}\|$ is uniformly bounded.
-Indeed, by (3.2) and (3.3), we derive that
+We proceed to establish the following statements for any fixed $k$
+  1. (i)
+    The supremum $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}\|$ is bounded.
+  2. (ii)
+    The supremum of the difference $\|u_{k}^{\delta}-u_{k}\|$ over all realizations in $\mathcal{F}_{k}$, i.e., $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}^{\delta}-u_{k}\|$, is uniformly bounded.
+To prove claim (i), we invoke ((2.1)) and ((2.2)), which yields that
 $\|u_{k+1}-\hat{u}\|$ $\leq\|u_{k}-\hat{u}\|+\omega_{k}\|F_{i_{k}}^{\prime}(u_{k})^{*}\|\|F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}\|+\lambda_{k}\|u_{0}-u_{k}\|$
 $\leq\left(1+\lambda_{k}+\omega_{k}\frac{L^{2}}{1-\eta}\right)\|u_{k}-\hat{u}\|+\lambda_{k}\|u_{0}-\hat{u}\|$.
-This, together with an induction argument, establishes the claim (i). Similarly, we note that
+This recursive inequality implies that the sequence $\{\|u_{k}-\hat{u}\|\}$ grows at most linearly, and by applying an induction argument, we conclude that claim (i) is established.
+Similarly, to establish claim (ii), we note that
 $\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|$ $\leq\|F_{i_{k}}(u_{k}^{\delta})-F_{i_{k}}(u_{k})\|+\|F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}\|+\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
 ((3.11)) $\leq\frac{L}{1-\eta}\left(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|\right)+\|y^{\delta}-y^{\dagger}\|,$
-and consequently 3.4 and 3.4 derive that
+and consequently 3.2 and 3.2 derive that
 $\|u_{k+1}^{\delta}-u_{k+1}\|$ $\leq\omega_{k}\left(\frac{L}{1-\eta}\left(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|\right)+\|y^{\delta}-y^{\dagger}\|\right)\|F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}-F_{i_{k}}^{\prime}(u_{k})^{*}\|$
 $+\omega_{k}L\|(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger})\|+(1-\lambda_{k})\|u_{k}^{\delta}-u_{k}\|.$
-This, along with mathematical induction establishes claim (ii).
-Let $c=\frac{L}{1-\eta}\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}^{k}}(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|)+\|y^{\delta}-y^{\dagger}\|$.
-Then, it follows from 3.4, 3.4 and (3.2) that
+This, together with a standard induction argument, confirms claim (ii). Consequently, leveraging both claims, we may now proceed to let
+$c=\frac{L}{1-\eta}\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}^{k}}(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|)+\|y^{\delta}-y^{\dagger}\|.$
+Then, it follows from 3.2, 3.2 and (2.1) that
 $\lim_{\delta\to 0^{+}}\|u_{k+1}^{\delta}-u_{k+1}\|$ $\leq(1-\lambda_{k})\lim_{\delta\to 0^{+}}\|u_{k}^{\delta}-u_{k}\|+c\omega_{k}\lim_{\delta\to 0^{+}}\|F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}-F_{i_{k}}^{\prime}(u_{k})^{*}\|$
 $+\omega_{k}L\lim_{\delta\to 0^{+}}\|(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger})\|.$
-Finally, the result follows from the continuity of the operators $F_{i}$ and $F_{i}^{\prime}$, i.e.,
-$\lim_{\delta\to 0^{+}}\|u_{k}^{\delta}-u_{k}\|=0\implies\lim_{\delta\to 0^{+}}\mathbb{E}[\|u_{k}^{\delta}-u_{k}\|^{2}]=0.$
-This completes the proof. ∎
-The final result of this section establishes the regularizing nature of IRSGD method.
+Finally, by the continuity of the operators $F_{i}$ and $F_{i}^{\prime}$, we obtain $\lim_{\delta\to 0^{+}}\|u_{k}^{\delta}-u_{k}\|=0,$ which, in turn, implies $\lim_{\delta\to 0^{+}}\mathbb{E}\big{[}\|u_{k}^{\delta}-u_{k}\|^{2}\big{]}=0.$ This completes the proof.
+∎
+We are now prepared to formally demonstrate the central result of this subsection, which concerns the regularization property of the IRSGD method.
 Theorem 3.2.
-Let all the assumptions of Lemma 3.1 hold.
-Let $\{y^{\delta}\}$ be the sequence of noisy data such that Assumption 3.2 hold and let IRSGD iteration corresponding to noisy data is stopped at $k_{*}(y^{\delta})$ defined via Rule 1.
+Assume that every condition in Lemma 2.1 is met.
+Suppose that $\{y^{\delta}\}$ is the noisy data sequence such that Assumption 2.2 hold and let IRSGD iteration corresponding to noisy data is stopped at $k_{*}(y^{\delta})$ defined via Rule 1.
 Then there exist a solution $\hat{u}$ of (1.2) such that
@@
 Proof.
-Let $\{y^{\delta_{m}}\}$ be a subsequene of $\{y^{\delta}\}$, and let $K:=\lim\inf_{m\to\infty}k_{*}(y^{\delta_{m}}).$ By taking a subsequence of $\{y^{\delta_{m}}\}$ if necessary, we may assume that $K=\lim_{m\to\infty}k_{*}(y^{\delta_{m}})$ and, according to Lemma 3.5, we have
-((3.37)) $\mathbb{E}[\|u_{k}^{\delta_{m}}-u_{k}\|^{2}]\to 0\quad\text{as}\quad m\to\infty.$
-for all $0\leq k\leq K.$ Let $\hat{u}$ be the limit of $u_{k}$, which is a solution of (1.2) in $B_{\rho+c(\rho)}(u_{0}).$ First, we assume that $K<\infty,$ then we have $k_{*}(y^{\delta_{m}})=K$ for large $m.$ Thus, from Lemma 3.2 it follows that
-$\|F(u_{K}^{\delta_{m}})-y^{\delta_{m}}\|\rightarrow 0\quad\text{as}\quad m\rightarrow\infty.$
-By engaging (3.37) and the continuity of $F$, we have $F(u_{K})=y^{\dagger}.$ In other words, $K$th iteration of IRSGD method with exact data is a solution of $F(u)=y^{\dagger}.$ Hence, the iteration terminates with $\hat{u}=u_{K}$ and $u_{K}^{\delta_{m}}\to\hat{u}$ as $m\to\infty.$
-If $K=\infty,$ then Theorem 3.1 and Lemma 3.5 guarantees that, for each $\epsilon>0,$ there exist $\hat{k}\in\mathbb{N}$ and $\overline{m}\in\mathbb{N}$ such that
+[…]
+$\lim_{\delta\to 0}\|u_{k_{*}(y^{\delta})}^{\delta}-\hat{u}\|^{2}]=0\quad\text{on the event}\daleth.$
+[…]
+((3.12)) $\mathbb{E}[\|u_{k}^{\delta_{m}}-u_{k}\|^{2}]\to 0\quad\text{as}\quad m\to\infty\ \ \text{for all}\ 0\leq k\leq K.$
+There are two possible cases to consider:
+Case I: Suppose $K<\infty$. Then, for sufficiently large $m$, we have $k_{*}(y^{\delta_{m}})=K$. Thus, from Lemma 2.2 it follows that
+$\mathbb{E}\left[\|F(u_{K}^{\delta_{m}})-y^{\delta_{m}}\|\right]\rightarrow 0\quad\text{as}\quad m\rightarrow\infty.$
+By engaging (3.12) and the continuity of $F$, we have $F(u_{K})=y^{\dagger}.$ In other words, $K$th iteration of IRSGD method with exact data is a solution of $F(u)=y^{\dagger}.$ Hence, the method terminates with $\hat{u}=u_{K}$ and $u_{K}^{\delta_{m}}\to\hat{u}$ as $m\to\infty.$
+Case II: Suppose $K=\infty$. Then Theorem 3.1 and Lemma 3.2 guarantees that, for each $\epsilon>0,$ there exist $\hat{k}\in\mathbb{N}$ and $\overline{m}\in\mathbb{N}$ such that
 ((3.13)) $\mathbb{E}[\|u_{\hat{k}}-\hat{u}\|^{2}]<\frac{\epsilon}{8},\quad\text{and}\quad\mathbb{E}[\|u_{\hat{k}}^{\delta_{m}}-u_{\hat{k}}\|^{2}]<\frac{\epsilon}{8}\quad\text{for all}\quad m\geq\overline{m}.$
-At this stage, we apply the full conditional on (3.16) to arrive at
+At this stage, we apply the full conditional on 2.2 to arrive at
 ((3.14)) $\mathbb{E}[\|u_{K}^{\delta_{m}}-\hat{u}\|^{2}]\leq 2\mathbb{E}[\|u_{\hat{k}}^{\delta_{m}}-u_{\hat{k}}\|^{2}+2\mathbb{E}[\|u_{\hat{k}}-\hat{u}\|^{2}+C_{2}\sum_{l=\hat{k}}^{K-1}\lambda_{l}+C_{3}\sum_{l=\hat{k}}^{K-1}\frac{1}{(k+a)^{2}},$
-where $C_{2}=2\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})$ and $C_{3}=\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau}.$ By choosing $\hat{k}$ such that
+where $C_{2}=\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)$ and $C_{3}=\frac{M\Omega(1+\eta)}{\tau}.$ By choosing $\hat{k}$ such that
 $\sum_{l=\hat{k}}^{K-1}\lambda_{l}\leq\frac{\epsilon}{4C_{2}}\quad\text{and}\quad\sum_{l=\hat{k}}^{K-1}\frac{1}{(k+a)^{2}}\leq\frac{\epsilon}{4C_{3}},$
@@
 $\mathbb{E}[\|u_{K}^{\delta_{m}}-\hat{u}\|^{2}]\leq\epsilon\quad\text{for all}\quad m\geq\overline{m}.$
-This completes the proof. ∎
-4. Convergence rates
-The main results of this section are presented in Theorem 4.3 and Theorem 4.4, which respectively address the rate of convergence under the discrepancy principle (1.7) and the rate of convergence under the heuristic rule 1.
-As a by-product, convergence rates for SGD (1.9) under these two stopping criteria can also be derived by selecting $\lambda_{k}=0$ in the iterative scheme (3.1).
-These outcomes signify the second major contribution of the work.
-Organisation: This section is organised as follows.
-At the outset, we introduce additional assumptions necessary for the analysis of convergence rates.
-Subsections 4.1 and 4.2 are dedicated to Theorems 4.1, 4.2, respectively and these theorems play an important role in deriving the convergence rates.
-In subsection 4.3, we establish convergence rates under a-posteriori stopping criteria, while subsection 4.4 focuses on deriving convergence rates using heuristic stopping criteria.
-We employ the following shorthand notation. Let
-((4.1)) $K_{i}=F_{i}^{\prime}(u^{\dagger}),\quad K=\begin{pmatrix}K_{0}\\ \vdots\\ K_{P-1}\end{pmatrix}\ \ \text{and}\quad B=K^{*}K=\sum_{i=0}^{P-1}K_{i}^{*}K_{i}.$
-Additionally, we commonly use the abbreviated notation
-((4.2)) $\prod_{j}^{k-1}(B_{\lambda})=\prod_{i=j}^{k-1}(I-\lambda_{i}-\omega_{i}B),$
-with the convention $\prod_{j}^{k}(B_{\lambda})=I$ for $j>k$. Also, for $h\geq 0,$ and $j\in\mathbb{N},$ we define
-((4.3)) $\tilde{h}=h+\frac{1}{2}\quad\text{and}\quad\Theta_{j}^{h}=\|B^{h}\prod_{j+1}^{k-1}(B_{\lambda})\|.$
-Note that our primary focus is on the cases where $h=0$ and $h=\frac{1}{2}$.
-Following the approach in [14], we assume the following conditions.
-Assumption 4.1 (Range invariance condition).
-There is a family of bounded operators $\{R_{u}^{i}:u\in B_{\rho+c(\rho)}(u_{0})\}$ for all $i\in\{0,1,\ldots,P-1\}$ such that
-$F_{i}^{\prime}(u)=R_{u}^{i}F_{i}^{\prime}(u^{\dagger})\quad\text{and}\quad R_{u}=\text{diag}(R_{u}^{i}):Y^{P}\to Y^{P},$
-with
-$\|R_{u}-I\|\leq c_{R}\|u-u^{\dagger}\|,\quad u\in B_{\rho+c(\rho)}(u_{0})$
-for some positive constant $c_{R}>0$ and $R_{u}=\text{diag}(R_{u}^{i})$ (diagonal matrix).
-This assumption has been validated for a range of nonlinear inverse problems, such as parameter identification and nonlinear integral equations (see [8]).
-Assumption 4.2 (Source condition [16]).
-There exist $\gamma\in(0,1/2)$ and $z\in U$ such that
-((4.4)) $u^{\dagger}-u_{0}=(F^{\prime}(u^{\dagger})^{*}F^{\prime}(u^{\dagger}))^{\gamma}z.$
-This assumption reflects a particular smoothness requirement for the exact solution $u^{\dagger}$. The restriction on $\gamma$ arises due to technical considerations (see, for example, (7.9)).
-Assumption 4.3 (Stochastic range invariance condition).
-Under the notations defined in Assumption 4.1, there exist $\theta\in(0,1]$ and $c_{R}>0$ such that for the function $Q(u)=F(u)-y^{\delta}$ or $Q(u)=F^{\prime}(u^{\dagger})(u-u^{\dagger}),u_{k}^{\delta}\in B_{\rho+c(\rho)}(u_{0})$ and $w_{s}=su_{k}^{\delta}+(1-s)u^{\dagger},s\in[0,1],$ there holds
-$\mathbb{E}\left[\|(I-R_{w_{s}})Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}}$ $\leq c_{R}\mathbb{E}\left[\|u^{\delta}_{k}-u^{\dagger}\|^{2}\right]^{\frac{\theta}{2}}\mathbb{E}\left[\|Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}},$
-$\mathbb{E}\left[\|(I-R^{*}_{w_{s}})Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}}$ $\leq c_{R}\mathbb{E}\left[\|u^{\delta}_{k}-u^{\dagger}\|^{2}\right]^{\frac{\theta}{2}}\mathbb{E}\left[\|Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}}.$
-This assumption plays a crucial role in the convergence rates analysis (see [14, Lemma 4.6]) and allows for bounding the terms that involves conditional dependence.
-4.1. Recursion on the bias
-In this subsection, we derive a recursive formula for the bias of the IRSGD iterate $u_{k}^{\delta}.$ The following lemma gives a useful representation of the mean $\mathbb{E}[e_{k}^{\delta}]$.
-Lemma 4.1.
-Under the conditions of Assumption 4.1, the error $e_{k}^{\delta}:=u_{k}^{\delta}-u^{\dagger}$ exhibits
-((4.5)) $\mathbb{E}[e_{k}^{\delta}]=\Pi^{k-1}_{j=0}(B_{\lambda})e_{0}+\sum_{j=0}^{k-1}\omega_{j}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)-\sum_{j=0}^{k-1}\lambda_{j}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0}),$
-with the vector $v_{k}\in Y^{P}$ is given by
-((4.6)) $v_{k}=\underbrace{-\left(F(u^{\delta}_{k})-F(u^{\dagger})-K(u^{\delta}_{k}-u^{\dagger})\right)}_{\text{Linearization error}}+\underbrace{(I-R_{u^{\delta}_{k}}^{*})(F(u^{\delta}_{k})-y^{\delta})}_{\text{Range invariance}}.$
-Proof.
-From the definition of IRSGD iterates $u_{k}^{\delta}$ in (3.1) we have
-$e_{k+1}^{\delta}$ $=(1-\lambda_{k})e_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u^{\dagger}-u_{0}).$
-The relation $F^{\prime}_{i_{k}}(u^{\delta}_{k})^{\ast}=(R^{i_{k}}_{u^{\delta}_{k}}F^{\prime}_{i_{k}}(u^{\dagger}))^{\ast}=K^{\ast}_{i_{k}}({R^{i_{k}}_{u^{\delta}_{k}}})^{*}$ from Assumption 4.1 and (4.1) leads to
-$e^{\delta}_{k+1}=(1-\lambda_{k})e^{\delta}_{k}-\omega_{k}\{K^{\ast}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})+K^{\ast}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})-K^{\ast}_{i_{k}}v_{k,i_{k}}\}-\lambda_{k}(u^{\dagger}-u_{0}),$
-with the random variable $v_{k,i_{k}}$ defined by
-((4.7)) $v_{k,i_{k}}=-\left(F_{i_{k}}(u^{\delta}_{k})-F_{i_{k}}(u^{\dagger})-K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})\right)+(I-({R^{i_{k}}_{u^{\delta}_{k}}})^{*})(F_{i_{k}}(u^{\delta}_{k})-y^{\delta}_{i_{k}}).$
-Consequently, by the measurability of $u^{\delta}_{k}$ (and hence $e^{\delta}_{k}$) with regard to $\mathcal{F}_{k}$, we write
-$\mathbb{E}[e^{\delta}_{k+1}\mid\mathcal{F}_{k}]$ $=(I-\lambda_{k}-\omega_{k}B)e^{\delta}_{k}-\omega_{k}K^{\ast}(y^{\dagger}-y^{\delta})+\omega_{k}K^{\ast}v_{k}-\lambda_{k}(u^{\dagger}-u_{0}).$
-Then taking the full conditional, using (4.2) and applying the recursion repeatedly yields
-$\mathbb{E}[e_{k}^{\delta}]=\Pi^{k-1}_{j=0}(B_{\lambda})e_{0}+\sum_{j=0}^{k-1}\omega_{j}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)-\sum_{j=0}^{k-1}\lambda_{j}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0}).$
-This brings the proof to a conclusion. ∎
-Remark 4.1.
-The term $v_{k}$ in equation (4.6) accounts for both the linearization error of the nonlinear operator $F$, expressed as $F(u^{\delta}_{k})-F(u^{\dagger})-K(u^{\delta}_{k}-u^{\dagger})$, and the range invariance of the derivative $F^{\prime}(u)$ as specified in (ii) of Assumption 3.1 and Assumption 4.1.
-The following conclusion provides a reasonable bound on the mean $\mathbb{E}[v_{j}]$.
-Lemma 4.2.
-Under Assumption 3.1, for $v_{j}$ defined in (4.6), the following inequality holds
-((4.8)) $\|\mathbb{E}[v_{j}]\|\leq\frac{(3-\eta)c_{R}}{2(1-\eta)}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{\frac{1}{2}}\mathbb{E}[\|B^{\frac{1}{2}}e^{\delta}_{j}\|^{2}]^{\frac{1}{2}}+c_{R}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{\frac{1}{2}}\delta.$
-Proof.
-See [14, Lemma 4.3] for the proof, where $c_{R}>0$ denotes the same constant as specified in Assumption 4.1. ∎
-Finally, in the next theorem, we establish a bound for $e_{k}^{\delta}$ in terms of a weighted norm.
-Theorem 4.1.
-Under Assumptions 3.1, 4.1 and 4.2, for any $h\geq 0,$ there holds
-((4.9)) $\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|\leq\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}(\delta+\|\mathbb{E}[v_{j}]\|),$
-where $\|\mathbb{E}[v_{j}]\|$ is same as given in Lemma 4.2.
-Proof.
-As $B^{h}$ is a bounded linear operator for $h\geq 0$, by Lemma 4.1, we have
-$B^{h}\mathbb{E}[e_{k}^{\delta}]$ $=B^{h}\Pi^{k-1}_{j=0}(B_{\lambda})e_{0}+\sum_{j=0}^{k-1}\omega_{j}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)$
-((4.10)) $\hskip 28.45274pt-\sum_{j=0}^{k-1}\lambda_{j}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0}):=I+II+III.$
-To establish the validity of the theorem, it is sufficient to bound all three terms $I,II$ and $III$. We will address each of these terms individually. From (4.4), we have
-((4.11)) $\|I\|=\|B^{h}\Pi^{k-1}_{j=0}(B_{\lambda})B^{\gamma}z\|\leq\|\Pi^{k-1}_{j=0}(B_{\lambda})B^{\gamma+h}\|\|z\|.$
-Next, we estimate the term $II$ as
-((4.12)) $\Omega\left\|\sum_{j=0}^{k-1}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)\right\|\leq\Omega\sum_{j=0}^{k-1}\|B^{\tilde{h}}\Pi^{k-1}_{j+1}(B_{\lambda})\|\left(\delta+\|\mathbb{E}[v_{j}]\|\right).$
-Likewise, we can also estimate the term $III$ using (4.4) as
-$\sum_{j=0}^{k-1}\|\lambda_{j}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0})\|\leq\sum_{j=0}^{k-1}\lambda_{j}\|B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})B^{\gamma}z\|\leq\sum_{j=0}^{k-1}\lambda_{j}\|\Pi^{k-1}_{j+1}(B_{\lambda})B^{\gamma+h}\|\|z\|.$
-Plugging this along with (4.11) and (4.12) in 4.1 together with the notation of $\Theta_{j}^{h}$ concludes the proof. ∎
-Remark 4.2.
-The weighted norm $\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|$ is valuable because the upper bound presented in Theorem 4.1 incorporates $\mathbb{E}[\|B^{\frac{1}{2}}e_{k}^{\delta}\|^{2}]$ (due to $\|\mathbb{E}[v_{j}]\|$) with $h=\frac{1}{2}$.
-In the context of linear inverse problems where $R_{u}=I$, $c_{R}=0$ and $\|\mathbb{E}[v_{j}]\|=0$ (due to Lemma 4.2), the recursion simplifies to
-$\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|\leq\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\delta,$
-where these terms correspond to the approximation error and the data error, respectively.
-4.2. Recursion on variance
-In this subsection, we address the computational variance $\mathbb{E}[\|B^{h}(u^{\delta}_{k}-\mathbb{E}[u^{\delta}_{k}])\|^{2}]$, which arises from the random index $i_{k}$.
-The variance was first constrained in terms of the iteration noises $H_{j,1}$ and $H_{j,2}$, which are specified below in (4.14).
-Lemma 4.3.
-Under Assumption 4.1, the IRSGD iterate $u^{\delta}_{k}$ as outlined in (3.1), satisfies
-$\mathbb{E}\left[\|B^{h}(u^{\delta}_{k}-\mathbb{E}[u^{\delta}_{k}])\|^{2}\right]$ $\leq\Omega^{2}\left(\sum_{j=0}^{k-1}(\Theta^{\tilde{h}}_{j})^{2}\mathbb{E}[\|H_{j,1}\|^{2}]+2\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta^{\tilde{h}}_{i}\Theta^{\tilde{h}}_{j}\mathbb{E}[\|H_{i,1}\|\|H_{j,2}\|]\right.$
-((4.13)) $\left.\hskip 142.26378pt+\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\Theta^{\tilde{h}}_{i}\Theta^{\tilde{h}}_{j}\mathbb{E}[\|H_{i,2}\|\|H_{j,2}\|]\right),$
-where the random variables $H_{j,1}$ and $H_{j,2}$, respectively, defined as
-((4.14)) $H_{j,1}$ $=(K(u^{\delta}_{j}-u^{\dagger})-K_{ij}(u^{\delta}_{j}-u^{\dagger})\varphi_{ij})+((y^{\dagger}-y^{\delta})-(y^{\dagger}_{i}-y^{\delta}_{i})\varphi_{i_{j}}),$
-$H_{j,2}$ $=-\mathbb{E}[v_{j}]+v_{j,i_{j}}\varphi_{i_{j}}.$
-Here $v_{k}$ and $v_{k,i}$ are defined in (4.6) and (4.7),respectively, and $\varphi_{i}=(0,\ldots,0,1,0,\ldots,0)$ represents the canonical $i$th basis vector in $\mathbb{R}^{P}$.
-Proof.
-From the IRSGD iterates as defined in (3.1) along with the Assumption 4.1 that $F^{\prime}_{i_{k}}(u^{\delta}_{k})^{\ast}=(R^{i_{k}}_{u^{\delta}_{k}}F^{\prime}_{i_{k}}(u^{\dagger}))^{\ast}=K^{\ast}_{i_{k}}({R^{i_{k}}_{u^{\delta}_{k}}})^{*}$, we have
-((4.15)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}K^{\ast}_{i_{k}}({R^{i_{k}}_{u^{\delta}_{k}}})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
-Rearranging the terms of (4.15) to write it as
-((4.16)) $u_{k+1}^{\delta}$ $=u_{k}^{\delta}-\omega_{k}\left(K^{*}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})+K^{*}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})+K_{i_{k}}^{*}\left(F_{i_{k}}(u^{\delta}_{k})-F_{i_{k}}(u^{\dagger})\right.\right.$
-$\left.\left.-K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})-(I-({R^{i_{k}}_{u^{\delta}_{k}}})^{*})(F_{i_{k}}(u^{\delta}_{k})-y^{\delta}_{i_{k}})\right)\right)-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
-Further, from the definition of $v_{k,i}$ as in (4.7), we arrive at
-((4.17)) $u^{\delta}_{k+1}=u^{\delta}_{k}-\omega_{k}K^{*}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})-\omega_{k}K^{*}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})+\omega_{k}K^{*}_{i_{k}}v_{k,i_{k}}-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
-By the definition of $v_{k}$ in (4.6) and the measurability of $u^{\delta}_{k}$ with respect to $\mathcal{F}_{k}$, we obtain
-$\mathbb{E}[u^{\delta}_{k+1}|\mathcal{F}_{k}]=u^{\delta}_{k}-\omega_{k}B(u^{\delta}_{k}-u^{\dagger})-\omega_{k}K^{*}(y^{\dagger}-y^{\delta})+\omega_{k}K^{*}v_{k}-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
-Taking the full conditional yields
-((4.18)) $\mathbb{E}[u^{\delta}_{k+1}]=\mathbb{E}[u^{\delta}_{k}]-\omega_{k}B\mathbb{E}[u^{\delta}_{k}-u^{\dagger}]-\omega_{k}K^{*}(y^{\dagger}-y^{\delta})+\omega_{k}K^{*}\mathbb{E}[v_{k}]-\lambda_{k}\mathbb{E}[u_{k}^{\delta}-u_{0}].$
-Thus, subtracting (4.18) from (4.17) shows that $x_{k}:=u^{\delta}_{k}-\mathbb{E}[u^{\delta}_{k}]$ satisfies
-((4.19)) $x_{k+1}=(I-\lambda_{k}-\omega_{k}B)x_{k}+\omega_{k}G_{k},$
-with $x_{0}=0$ and the iteration noise $G_{k}$ given by $G_{k}=G_{k,1}+G_{k,2}$, where
-$G_{k,1}$ $=(B(u^{\delta}_{k}-u^{\dagger})-K^{*}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger}))+(K^{*}(y^{\dagger}-y^{\delta})-K^{*}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})),$
-$\ G_{k,2}$ $=-(K^{*}\mathbb{E}[v_{k}]-K^{*}_{i_{k}}v_{k,i_{k}}).$
-Repeatedly applying the recursion (4.19) with $x_{0}=0$ leads to
-$x_{k}=\sum_{j=0}^{k-1}\omega_{j}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j}.$
-With the decomposition of $G_{j}=G_{j,1}+G_{j,2}$, we directly obtain
-((4.20)) $\mathbb{E}[\|B^{h}x_{k}\|^{2}]$ $=\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\mathbb{E}[\langle B^{h}\Pi^{k-1}_{i+1}(B_{\lambda})G_{i,1},B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j,1}\rangle]$
-$+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\mathbb{E}[\langle B^{h}\Pi^{k-1}_{i+1}(B_{\lambda})G_{i,1},B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j,2}\rangle]$
-$+\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\mathbb{E}[\langle B^{h}\Pi^{k-1}_{i+1}(B_{\lambda})G_{i,2},B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j,2}\rangle]:=B_{I}+B_{II}+B_{III}.$
-Next, we proceed to simplify the three terms. Noting that $u^{\delta}_{j}$ is measurable with respect to $\mathcal{F}_{j}$, we have $\mathbb{E}[G_{j,1}|\mathcal{F}_{j}]=0.$ Due to this, we claim that
-((4.21)) $\mathbb{E}[\langle B^{h}G_{i,1},B^{h}G_{j,1}\rangle]=0,\quad i\neq j.$
-Indeed, for $i>j$, there holds
-$\mathbb{E}[\langle B^{h}G_{i,1},B^{h}G_{j,1}\rangle|\mathcal{F}_{i}]=\langle B^{h}\mathbb{E}[G_{i,1}|\mathcal{F}_{i}],B^{h}G_{j,1}\rangle=0,$
-and taking the full conditional therefore establishes (4.21). Consequently, $B_{I}$ simplifies to
-((4.22)) $B_{I}=\Omega^{2}\sum_{j=1}^{k}\mathbb{E}[\|B^{h}\Pi^{k}_{j+1}(B_{\lambda})G_{j,1}\|^{2}].$
-Moreover, for $i>j$, a similar reasoning leads to $\mathbb{E}[\langle B^{h}G_{i,1},B^{h}G_{j,2}\rangle]=0,$ and thus
-((4.23)) $B_{II}=2\Omega^{2}\sum_{i=1}^{k}\sum_{j=i}^{k}\mathbb{E}[\langle B^{h}\Pi^{k}_{i+1}(B_{\lambda})G_{i,1},B^{h}\Pi^{k}_{j+1}(B_{\lambda})G_{j,2}\rangle].$
-Next, we proceed to further simplify $G_{j,1}$ and $G_{j,2}$.
-From the definitions of $H_{j,1}$ and $H_{j,2}$, and noting that $(K^{*})^{\dagger}$ represents the pseudoinverse of $K^{*}$, we obtain $(K^{*})^{\dagger}G_{j}=H_{j,1}+H_{j,2}$.
-Thus, by using triangle inequality along with (4.22) and (4.23) in (4.20), we arrive at
-$\mathbb{E}[\|B^{h}x_{k}\|^{2}]$ $\leq\Omega^{2}\sum_{j=0}^{k-1}\mathbb{E}[\|B^{h+1/2}\Pi^{k-1}_{j+1}(B_{\lambda})\|^{2}\|H_{j,1}\|^{2}]$
-$+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\|B^{h+1/2}\Pi^{k-1}_{i+1}(B_{\lambda})\|\|B^{h+1/2}\Pi^{k-1}_{j+1}(B_{\lambda})\|\mathbb{E}[\|H_{i,1}\|\|H_{j,2}\|]$
-$+\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\|B^{h+1/2}\Pi^{k-1}_{i+1}(B_{\lambda})\|\|B^{h+1/2}\Pi^{k-1}_{j+1}(B_{\lambda})\|\mathbb{E}[\|H_{i,2}\|\|H_{j,2}\|].$
-This together with the notation of $\Theta_{j}^{h}$ completes the proof. ∎
-The next result provides the bounds on $H_{j,1}$ and $H_{j,2}$ defined in (4.14).
-Theorem 4.2.
-Let Assumptions 3.1 and Assumption 4.3 be fulfilled. Then for any $h\in[0,\frac{1}{2}]$, there holds
-$\mathbb{E}$ $\left[\|B^{h}(\mathbb{E}[x^{\delta}_{k}]-x^{\delta}_{k})\|^{2}\right]\leq\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{\tilde{h}})^{2}\left(\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+\delta\right)^{2}$
-$+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{\tilde{h}}\Theta_{j}^{\tilde{h}}(\mathbb{E}\left[\|B^{1/2}e_{i}^{\delta}\|^{2}\right]^{1/2}+\delta)\left(\frac{(2+\theta-\eta)c_{R}}{(1+\theta)(1-\eta)}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}$
-$+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\left(\frac{(2+\theta-\eta)c_{R}}{(1+\theta)(1-\eta)}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}\right)^{2}.$
-Proof.
-We recall from [14, Lemma 4.6], with the modification being the exclusion of the scaling factor $n^{1/2}$, that
-((4.24)) $\mathbb{E}[\|H_{j,1}\|^{2}]^{1/2}$ $\leq\left(\mathbb{E}[\|B^{1/2}e^{\delta}_{j}\|^{2}]^{1/2}+\delta\right),$
-$\mathbb{E}[\|H_{j,2}\|^{2}]^{1/2}$ $\leq\left(c_{R}(2+\theta-\eta)\frac{(1+\theta)}{(1-\eta)}\mathbb{E}[\|B^{1/2}e^{\delta}_{j}\|^{2}]^{1/2}+c_{R}\delta\right)\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{\theta/2}.$
-The conclusion directly follows by using (4.24) in Lemma 4.3. ∎
-4.3. Convergence rates using a-posteriori stopping criteria
-We recall that our primary objective is to establish the convergence rates using the heuristic stopping criteria (i.e., Rule 1).
-However, before doing so, in this subsection, we first derive the convergence rates based on the a-posteriori stopping rule (1.7).
-We emphasize that the convergence analysis for the method with the stopping rule (1.7) follows a similar approach to the one we developed for the heuristic rule in Section 3 with some modification being required due to the change in the stopping rule.
-In this subsection, we begin by deriving bounds on the stopping index $k_{\delta}$ and the corresponding error $e_{k}^{\delta}:=u_{k}^{\delta}-u^{\dagger}$.
-Lemma 4.4.
-Suppose that Assumptions 3.1, 4.1 and 4.2 hold with $\|z\|$ sufficiently small, and let $k_{\delta}$ be the integer specified by (1.7).
-Also, suppose $y^{\delta}$ satisfies (1.3) and let the weighted parameter sequence $\{\lambda_{k}\}_{k\geq 0}$ be chosen as $\lambda_{k}=(k+1)^{-n}$ with $n>1$, then
-((4.25)) $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{1/2}\leq C_{\gamma}\|z\|k^{-\gamma}\quad\text{and}\quad\mathbb{E}[\|B^{1/2}e_{k}^{\delta}\|^{2}]^{1/2}\leq C_{\gamma}\|z\|k^{-\gamma-1/2},$
-for $0\leq k<k_{\delta},$ $C_{\gamma}$ is a positive constant depending only on $\gamma,\eta,n,\tau$ and $\theta$. In case of exact data $(\delta=0)$, (4.25) holds for all $k\geq 0.$
-Proof.
-Using Bias-variance decomposition, we know that
-((4.26)) $\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]=\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|^{2}+\mathbb{E}[\|B^{h}(e_{k}^{\delta}-\mathbb{E}[e_{k}^{\delta}])\|^{2}].$
-Using assertions of Theorem 4.1 and Theorem 4.2 in (4.26) with $p_{0}=\frac{(2+\theta-\eta)c_{R}}{(1+\theta)(1-\eta)}$, we have
-$\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]\leq\left(\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}(\delta+\|\mathbb{E}[v_{j}]\|)\right)^{2}\hskip 85.35826pt$
-((4.27)) $+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{\tilde{h}})^{2}\left(\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+\delta\right)^{2}+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{\tilde{h}}\Theta_{j}^{\tilde{h}}\left(\mathbb{E}\left[\|B^{1/2}e_{i}^{\delta}\|^{2}\right]^{1/2}+\delta\right)$
-$\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}\right)^{2}.$
-By plugging $\|\mathbb{E}[v_{j}]\|$ from Lemma 4.2 in (4.27) and the fact that $\frac{(3-\eta)c_{R}}{2(1-\eta)}<p_{0}$, we get
-$\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]\leq\bigg{(}\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}(\delta+p_{0}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{1/2}\mathbb{E}[\|B^{1/2}e^{\delta}_{j}\|^{2}]^{1/2}\hskip 28.45274pt$
-$+c_{R}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{1/2}\delta)\bigg{)}^{2}+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{\tilde{h}})^{2}\left(\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+\delta\right)^{2}+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{\tilde{h}}\Theta_{j}^{\tilde{h}}\left(\mathbb{E}\left[\|B^{1/2}e_{i}^{\delta}\|^{2}\right]^{1/2}+\delta\right)$
-$\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}\right)^{2}.$
-Let $s_{j}=\mathbb{E}[\|e_{j}^{\delta}\|^{2}]$ and $t_{j}=\mathbb{E}[\|B^{1/2}e_{j}^{\delta}\|^{2}]$. Then, by first setting $h=0$ and then $h=1/2$ in the last inequality gives
-((4.28)) $s_{k}$ $\leq\left(\Theta_{-1}^{\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1/2}(\delta+p_{0}s_{j}^{1/2}t_{j}^{1/2}+c_{R}\delta s_{j}^{1/2})\right)^{2}++2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{1/2}\Theta_{j}^{1/2}$
-$\left(t_{j}^{1/2}+\delta\right)\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1/2})^{2}\left(t_{j}^{1/2}+\delta\right)^{2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1/2}\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}\right)^{2},$
-((4.29)) $t_{k}$ $\leq\left(\Theta_{-1}^{1/2+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{1/2+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1}(\delta+p_{0}s_{j}^{1/2}t_{j}^{1/2}+c_{R}s_{j}^{1/2}\delta)\right)^{2}+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{1}\Theta_{j}^{1}$
-$\left(t_{j}^{1/2}+\delta\right)\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1})^{2}\left(t_{j}^{1/2}+\delta\right)^{2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1}\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}\right)^{2}.$
-Because of (3.3) and (1.7), we have
-$\tau\delta<\|y^{\delta}-F(u_{k}^{\delta})\|\leq\delta+\frac{1}{1-\eta}\|Ke_{j}^{\delta}\|.$
-From the measuribility of $e_{j}^{\delta}$ with respect to $\mathcal{F}_{j}$ and the inequality $\|Ke_{j}^{\delta}\|=\|B^{1/2}e_{j}^{\delta}\|,$ we get
-((4.30)) $\delta<\frac{1}{(\tau-1)(1-\eta)}\mathbb{E}[\|B^{1/2}e_{j}^{\delta}\|^{2}]^{1/2}=c_{\tau\eta}t_{j}^{1/2},\quad\text{where}\quad c_{\tau\eta}=\frac{1}{(\tau-1)(1-\eta)}.$
-For the sequence of weighted parameter $\{\lambda_{k}\}_{k\geq 0}$, where $\lambda_{k}=(k+1)^{-n}$ and using (4.30) in (4.28) and (4.29) with the notations $c^{\prime}=p_{0}+c_{R}c_{\tau\eta}$ and $c^{\prime\prime}=1+c_{\tau\eta}$ , we have
-((4.31)) $s_{k}\leq\left(\Theta_{-1}^{\gamma}\|z\|+\sum_{j=0}^{k-1}\Theta_{j}^{\gamma}(j+1)^{-n}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1/2}[c_{\tau\eta}+c^{\prime}s_{j}^{1/2}]t_{j}^{1/2}\right)^{2}\hskip 71.13188pt$
-$+2\Omega^{2}c^{\prime}c^{\prime\prime}\sum_{i=0}^{k-1}\Theta_{i}^{1/2}t_{i}^{1/2}\sum_{j=0}^{k-1}\Theta_{j}^{1/2}t_{j}^{1/2}s_{j}^{\theta/2}+\Omega^{2}(c^{\prime\prime})^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1/2})^{2}t_{j}+\Omega^{2}(c^{\prime})^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1/2}t_{j}^{1/2}s_{j}^{\theta/2}\right)^{2},$
-((4.32)) $t_{k}\leq\left(\Theta_{-1}^{1/2+\gamma}\|z\|+\sum_{j=0}^{k-1}\Theta_{j}^{1/2+\gamma}(j+1)^{-n}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1}[c_{\tau\eta}+c^{\prime}s_{j}^{1/2}]t_{j}^{1/2}\right)^{2}\hskip 42.67912pt$
-$+2\Omega^{2}c^{\prime}c^{\prime\prime}\sum_{i=0}^{k-1}\Theta_{i}^{1}t_{i}^{1/2}\sum_{j=0}^{k-1}\Theta_{j}^{1}t_{j}^{1/2}s_{j}^{\theta/2}+\Omega^{2}(c^{\prime\prime})^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1})^{2}t_{j}+\Omega^{2}(p_{0}+c_{R}c_{\tau\eta})^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1}t_{j}^{1/2}s_{j}^{\theta/2}\right)^{2}.$
-We claim that for all $0\leq j<k_{\delta}$, for a certain constant $C_{\gamma}$, there holds
-((4.33)) $s_{j}^{1/2}\leq C_{\gamma}\|z\|j^{-\gamma}\quad\text{and}\quad t_{j}^{1/2}\leq C_{\gamma}\|z\|j^{-\gamma-1/2}.$
-For $j=0$, (4.33) holds trivially. This is because Assumption 4.2 implies that
-((4.34)) $\|e_{0}\|\leq\|(F^{\prime}(u^{\dagger})^{*}F^{\prime}(u^{\dagger}))^{\gamma}z\|\leq\|F^{\prime}(u^{\dagger})^{*}F^{\prime}(u^{\dagger})\|^{\gamma}\|z\|\leq\|F^{\prime}(u^{\dagger})\|^{2\gamma}\|z\|.$
-Using Assumption 3.1 $(ii)$ in (4.34), we conclude that
-$\|e_{0}\|\leq L^{2\gamma}\|z\|\leq C_{\gamma}\|z\|.$
-for sufficiently large $C_{\gamma}$.
-For $j>0$, the proof proceeds by induction.
-We assume that (4.33) is true for all $0\leq j<k$ for some $k<k_{\delta}$, and then verify it for $j=k$.
-So using the induction hypothesis in (4.31) and (4.32) along with Lemma A1 and A3 with appropriate choices of $a,b$ and $c$, we have
-((4.35)) $s_{k}\leq\bigg{\{}\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}k^{-\gamma}+\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}\left(\frac{1}{1-\gamma}+\frac{1}{1-n}\right)+2^{1-\gamma-n}\|B^{a}\|\right)(k/2)^{1-\gamma-n}\right)\|z\|$
-$+\Omega c_{\tau\eta}\hslash\widetilde{C}_{1/2,1/2,0}(k/2)^{-\gamma}+\Omega c^{\prime}\hslash^{2}\widetilde{C}_{1/2,1/2,1/2}(k/2)^{-2\gamma}\bigg{\}}^{2}+2\Omega^{2}c^{\prime}c^{\prime\prime}\hslash^{2+\theta}\widetilde{C}_{1/2,1/2,0}\widetilde{C}_{1/2,1/2,\theta/2}(k/2)^{-\gamma(2+\theta)}$
-$+\frac{\Omega^{2}}{2}(c^{\prime\prime})^{2}\hslash^{2}\left(\frac{1}{e\omega}(l_{1,0}+2\ln k)+2^{-2\gamma-1}\|B\|\right)(k/2)^{-2\gamma-1}+\Omega^{2}(c^{\prime})^{2}\hslash^{2(1+\theta)}\widetilde{C}_{1/2,1/2,\theta/2}^{2}(k/2)^{-2\gamma(1+\theta)}.$
-Further, using the fact that for any $r>0,$ there holds
-((4.36)) $s^{-r}\ln{s}\leq(er)^{-1}\quad\forall s\geq 0,$
-and thus $k^{-1}\ln{k}\leq e^{-1},$ we further obtain from (4.35) using the notation $\hslash=C_{\gamma}\|z\|$ that
-((4.37)) $s_{k}\leq\hslash^{2}\bigg{\{}\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}2^{-\gamma}+\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}\left(\frac{1}{1-\gamma}+\frac{1}{1-n}\right)+2^{1-\gamma-n}\|B^{a}\|\right)(k/2)^{1-n}\right)C_{\gamma}^{-1}$
-$+\Omega c_{\tau\eta}\widetilde{C}_{1/2,1/2,0}+\Omega c^{\prime}\hslash\widetilde{C}_{1/2,1/2,1/2}(k/2)^{-\gamma}\bigg{\}}^{2}+2\Omega^{2}c^{\prime}c^{\prime\prime}\hslash^{\theta}\widetilde{C}_{1/2,1/2,0}\widetilde{C}_{1/2,1/2,\theta/2}(k/2)^{-\gamma\theta}$
-$+\frac{\Omega^{2}}{2}(c^{\prime\prime})^{2}\bigg{\{}\left(\frac{l_{1,0}}{e\omega}+2^{-2\gamma-1}\|B\|\right)(k/2)^{-1}+\frac{4}{e^{2}\omega}\bigg{\}}+\Omega^{2}(c^{\prime})^{2}\hslash^{2\theta}\widetilde{C}_{1/2,1/2,\theta/2}^{2}(k/2)^{-2\gamma\theta}\Bigg{]}(k/2)^{-2\gamma}$
-$:=\hslash^{2}B_{\gamma,k}(k/2)^{-2\gamma}.\hskip 256.0748pt$
-At this point, let us assume that $B_{\gamma,k}>0$ be such that $B_{\gamma,k}\leq 2^{-2\gamma}$ (for sufficiently small $\|z\|$ and sufficiently large $C_{\gamma}$, this condition is satisfied), which means that
-((4.38)) $s_{k}\leq\hslash^{2}2^{-2\gamma}\left(\frac{k}{2}\right)^{-2\gamma}\leq\hslash^{2}k^{-2\gamma}=C_{\gamma}^{2}\|z\|^{2}k^{-2\gamma}.$
-This completes the induction step and the proof for $s_{k}$. The rest part of the proof is moved to Appendix as it follows on the lines of that of $s_{k}$. ∎
-Theorem 4.3.
-Under the assumptions of Lemma 4.4, we have
-((4.39)) $k_{\delta}\leq C_{\gamma,\eta,n,\tau,\theta}\|z\|^{\frac{2}{2\gamma+1}}\delta^{-\frac{2}{2\gamma+1}}$
-where $C_{\gamma,\eta,n,\tau,\theta}$ is a positive constant depending only on $\gamma,\eta,n,\tau,\theta$. In addition to that
-((4.40)) $\mathbb{E}[\|e_{k_{\delta}}^{\delta}\|^{2}]^{1/2}\leq C_{\gamma}C_{\gamma,\eta,n,\tau,\theta}\|z\|^{1/(2\gamma+1)}\delta^{2\gamma/(2\gamma+1)},$
 […]
-Proof.
-We apply (4.30) along with (4.33) with $k=k_{\delta}$ to obtain
-$k_{\delta}^{\gamma+1/2}\leq c_{\tau\eta}C_{\gamma}\|z\|/\delta$
-with $C_{\gamma}$ as in Lemma 4.4. Hence, we obtain
-$k_{\delta}\leq C_{\gamma,\eta,n,\tau,\theta}\|z\|^{\frac{2}{2\gamma+1}}\delta^{-\frac{2}{2\gamma+1}},$
-where $C_{\gamma,\eta,n,\tau,\theta}=(c_{\tau\eta}C\gamma)^{\frac{2}{2\gamma+1}}.$ The error estimate (4.40) originates from (4.25). ∎
-4.4. Convergence rates using heuristic type stopping criteria
-In this section, we will derive an error estimate for the expression $u^{\delta}_{k_{*}(y^{\delta})}-u^{\dagger}$ in mean squared norm.
-Lemma 4.5.
-Let Assumption 3.1 hold and $y^{\delta}$ be a noisy data satisfying Assumption 3.2, also let $k_{*}:=k_{*}(y^{\delta})$ be the integer defined by Rule 1. Additionally, assume that
-((4.41)) $\|F(u_{0})-y^{\delta}\|>\tau\|y^{\dagger}-y^{\delta}\|.$
-If Assumption 4.1 holds with $20<\gamma<1/$ 2 for sufficiently small $\|z\|,$ then
-((4.42)) $(k_{*}+a)\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}\|z\|^{\frac{2}{2\gamma+1}}\delta^{-\frac{2}{2\gamma+1}},$
-where $C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}=\frac{\tau^{2}(1+a)C_{\gamma,\eta,n,\tau,\theta}}{\varkappa^{2}}$ is a constant depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$.
-Proof.
-By Assumption 3.2 and the definition of $k_{*}$ and $k_{\delta}$, we have
-$\varkappa^{2}(k_{*}+a)\delta^{2}\leq\Psi(k_{*},y^{\delta})\leq\Psi(k_{\delta},y^{\delta})=(k_{\delta}+a)\|F(u_{k_{\delta}}^{\delta})-y^{\delta}\|^{2}\leq\tau^{2}(k_{\delta}+a)\delta^{2}.$
-This implies that
-$\varkappa^{2}(k_{*}+a)\leq\tau^{2}(k_{\delta}+a)\leq\tau^{2}(1+a)k_{\delta},\ \text{as}\ k_{\delta}\geq 1$
-since $\|F(u_{0})-y^{\delta}\|>\tau\delta$. Therefore, by virtue of Theorem 4.3, the proof is complete. ∎
-Lemma 4.6.
-Let Assumption 3.1 hold and let $k_{*}$ be the integer defined by Rule 1 and let $y^{\delta}$ satisfy (1.3). If Assumption 4.1 holds with $20<\gamma<1/$ 2 for sufficiently small $\|z\|,$ then
-((4.43)) $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{1/2}\leq\tilde{C}_{\gamma}\|z\|k^{-\gamma},\ \ \ \text{and}\ \ \ \mathbb{E}[\|B^{1/2}e_{k}^{\delta}\|^{2}]^{1/2}\leq\tilde{C}_{\gamma}\|z\|k^{-\gamma-1/2}$
-for $0\leq k\leq k_{*},$ $\tilde{C}_{\gamma}>0$ is a constant depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$. In case of exact data $(\delta=0)$, (4.25) holds for all $k\geq 0.$
-Proof.
-The proof follows on the similar lines as given in Lemma 4.4 with a small modification in (4.30) due to heuristic type stopping rule $k_{*}$. By virtue of Lemma 4.5, we have
-((4.44)) $\delta\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma+1/2}\|z\|(k_{*}+a)^{-\gamma-1/2}\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma+1/2}\|z\|j^{-\gamma-1/2}.$
-Using (4.44) in (4.28) and (4.29) and following the same procedure as in Lemma 4.4 we get the desired result. ∎
-Remark 4.3.
-In comparison to Lemma 4.4, the only difference in Lemma 4.6 is a constant factor. This variation arises due to (4.44), which is applied in the proof in place of (4.30).
-Theorem 4.4.
-Under the assumptions of Lemma 4.6 we have
-((4.45)) $\mathbb{E}[\|e_{k_{*}}^{\delta}\|^{2}]^{1/2}\leq\frac{\tilde{C}_{\gamma}}{C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}}\|z\|^{\frac{1}{2\gamma+1}}\delta^{\frac{2\gamma}{2\gamma+1}},$
-with some constants $\tilde{C}_{\gamma},C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}>0,$ depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$.
-Proof.
-We apply (4.42) along with (4.43) with $k=k_{*}$ as
-$\mathbb{E}[\|e_{k_{*}}^{\delta}\|^{2}]^{1/2}$ $\leq\tilde{C}_{\gamma}\|z\|C^{-\gamma}_{\gamma,\eta,n,\tau,\theta,\varkappa,a}\left(\frac{\|z\|}{\delta}\right)^{\frac{-2\gamma}{2\gamma+1}}\leq\frac{\tilde{C}_{\gamma}}{C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma}}\|z\|^{\frac{1}{2\gamma+1}}\delta^{\frac{2\gamma}{2\gamma+1}}.$
-Hence we get the desired result.
-4.5. Comparison of convergence rates
-In this subsection, we present a detailed comparison of various iterative methods.
-As shown in Table 1, the convergence rates achieved by our approach are comparable to those reported in the existing literature.
-Furthermore, we selected the step size $\omega_{k}\in[\omega,\Omega]$ in such a way that it can either remain constant or decay monotonically, depending on the specific method used.
-Method Step-size Stopping criteria Convergence Rate
-LM [8] Constant a posteriori (1.7) $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma>0$
-IRLM [16] Constant a-priori $O(\delta^{1/2})$
-SGD [14] Decaying $k^{*}=\left[(\frac{\delta}{\|w\|})^{-\frac{2}{(2\gamma+1)(1-\alpha)}}\right]$ $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma\in(0,\frac{1}{2})$ and $\alpha$ close to $1$
-SGD [12] Decaying a posteriori (1.7) Not available
-IRSGD (This paper) Any a posteriori (1.7) $O(\delta^{\frac{2\gamma}{(2\gamma+1)}}),\gamma\in(0,\frac{1}{2})$
-IRSGD (This paper) Any heuristic type (Rule 1) $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma\in(0,\frac{1}{2})$
-Table 1. Convergence rates of various iterative methods.
-5. Numerical Experiments
-The objective of this section is to discuss the applicability of proposed scheme using Rule 1 on linear as well as nonlinear ill-posed inverse problems.
+4. Numerical Validations and Discussions
+In this section, we present a series of numerical experiments to assess the practical performance of the proposed scheme, implemented in conjunction with Rule 1.
+The experiments are designed to illustrate the effectiveness of the method in solving ill-posed inverse problems, encompassing both linear and nonlinear cases.
 4.1. Linear case
-Consider the following problem of linear ill-posed system (see [15] for details on this problem)
-[…]
-derived from first-kind linear integral equations on $[a,b]$ by sampling at $t_{i}\in[a,b]$ for $i=0,\ldots,P-1,$ where $t_{i}=a+\frac{i(b-a)}{P-1}$ for $i=0,\ldots,P-1,$ and the kernel $k(\cdot,\cdot)$ is continuous on $[a,b]\times[a,b]$.
-In this case, the iterative scheme (3.1) becomes
-[…]
-where $i_{k}$ is the randomly chosen index at the $k$th iteration step from the index set $\{0,\ldots,P-1\}$.
-We consider the system (5.1) with $[a,b]=[-6,6],P=1000,$ and $k(t,s)=\varphi(t-s),$ where $\varphi(t)=(1+\cos(\frac{\pi t}{3}))_{\chi_{\{|t|<3\}}}$ to test the effectiveness of technique (5.2).
-We consider the following solution to be reconstructed
+We consider the following lnear system.
+((4.1)) $F_{i}(u)=\int_{a}^{b}k(t_{i},s)u(s)\,ds=y_{i},\hskip 14.22636pti=0,1,\ldots,P-1,$
+which is derived from a linear integral equation on $[a,b]$, discretized by sampling at points $t_{i}\in[a,b]$ for $i=0,1,\ldots,P-1,$ where $t_{i}=a+\frac{i(b-a)}{P-1}.$ For a comprehensive discussion on such problems, see [20].
+It is assumed that the kernel function $k(\cdot,\cdot)$ maintains continuity throughout the domain $a,b]\times[a,b]$.
+To evaluate the effectiveness of the IRSGD method (1.11), we apply it to system (4.1) with P=1000,[a,b]=[6,6],P=1000,[a,b]=[-6,6], and the kernel function given by $k(t,s)=\varphi(t-s),$ where $\varphi(t)=(1+\cos(\frac{\pi t}{3}))_{\chi_{\{|t|<3\}}}.$ The goal is to reconstruct the target solution given as
 $u^{\dagger}(s)=\sin\left(\frac{\pi s}{12}\right)+\sin\left(\frac{\pi s}{3}\right)+\frac{1}{200}s^{2}(1-s).$
-Instead of the exact data $y_{i}=F_{i}(u^{\dagger})$, we use the noisy data
-((5.3)) $y^{\delta}_{i}=y_{i}+\delta_{\text{rel}}\times|y_{i}|\times{\epsilon_{i}},\quad i=0,\ldots,P-1,$
-where $\epsilon_{i}$ is a random noise that follows a standard Gaussian distribution $N(0,1)$, and $\delta_{\text{rel}}$ is the relative noise level.
-Based on the division of $[-6,6]$ into $P-1$ subintervals of similar length, the trapezoidal rule approximates the integrals involved in the procedure and discretized into a linear system of size $n=1000$ with the forward operator $F_{n\times n}$.
+In place of $y_{i}=F_{i}(u^{\dagger})$ (exact data), we consider noisy observations given by
+((4.2)) $y^{\delta}_{i}=y_{i}+{\epsilon_{i}}\delta_{\text{rel}}|y_{i}|,\quad i=0,1,\ldots,P-1,$
+where $\delta_{\text{rel}}$ denotes the relative noise level and $\epsilon_{i}$ represents random noise drawn from a standard Gaussian distribution $N(0,1)$.
+To discretize the problem, we partition the interval [6,6][-6,6] into P1P-1 equal subintervals and approximate the integrals using the trapezoidal rule.
+This leads to a discretized linear system of dimension $n=1000$ with the corresponding forward operator $F_{n\times n}$.
 We execute the method (1.11) with the numerical settings as follows:
@@
   • (d)
-    The maximum number of iterations is set to $k_{\text{max}}=1000$.
+    The maximum number of iterations, denoted by $k_{\text{max}}$, is set to $1000$.
   • (e)
-    We employ noisy data, where $\delta_{\text{rel}}=10^{-1},10^{-2},10^{-3}$ and $10^{-4}$ represent four different relative noise levels.
+    Noisy data is employed with four noise levels given by $\delta_{\text{rel}}=10^{-1},10^{-2},10^{-3}$ and $10^{-4}$.
 We mainly perform the three experiments as:
@@
 In order to investigate the convergence behavior of $\Psi(k,y^{\delta})$ as $\delta\to 0$, we present the plots of the values of $\Psi(k,y^{\delta})$ versus $k$ in Figure 1 using the fixed constant $a=100$ (as in Rule 1) and for different levels of noise.
-Figure 1. Convergence behavior of $\Psi(k,y^{\delta})$ as defined in Rule 1, alongside the increase in the stopping index $k_{*}$ as the noise level decreases.
+Figure 1. Convergence behavior of $\Psi(k,y^{\delta})$ for varying noise levels, illustrating the sensitivity of the proposed stopping rule to measurement perturbations.
 Figure 1 clearly shows that the noise level in the input data has a significant effect on the convergence behavior of $\Psi(k,y^{\delta})$.
@@
 Additionally, it demonstrates that $\Psi(k_{*},y^{\delta})\to 0$ as $\delta\to 0$, which supports the assertion in Lemma 2.2.
-Furthermore, it is also evident that as the noise level decreases, the stopping index $k_{*}$ increases.
+Furthermore, as $\delta$ decreases, the stopping index $k_{*}$ increases.
 4.1.2. Reconstructed solutions
@@
 $\delta_{\text{rel}}$ $a$ in Rule 1
-[…]
+$1\text{e}1$ $1\text{e}2$ $1\text{e}3$ $1\text{e}4$
 $k_{*}$ $E_{k_{*}}^{\delta}$ $k_{*}$ $E_{k_{*}}^{\delta}$ $k_{*}$ $E_{k_{*}}^{\delta}$ $k_{*}$ $E_{k_{*}}^{\delta}$
@@
 $1\text{e}^{-4}$ 3304 $5.341\text{e}^{-4}$ 3167 $5.134\text{e}^{-4}$ 867 $6.111\text{e}^{-4}$ 4874 $3.185\text{e}^{-4}$
-Table 2. Results of IRSGD with different values of $a$ in Rule 1.
+Table 1.
+Performance of the IRSGD method for various values of the parameter $a$ in Rule 1 and different relative noise levels $\delta_{\text{rel}}$.
+For each combination, the table reports the stopping index $k_{*}$ and the corresponding reconstruction error $E_{k_{*}}^{\delta}$.
 4.1.3. Comparison between SGD and IRSGD
-We compare the results of IRSGD with SGD.
-For this comparison, we set the constant $a$ in the rule (referred to as Rule 1) to $a=100$.
-The numerical results for the comparison are presented in Table 3.
+For this comparison, we set the constant $a$ in the rule (referred to as Rule 1) to $a=100$. The numerical findings are reported in Table 2.
 Methods $\delta_{\text{rel}}$ $k_{*}$ $E_{k_{*}}^{\delta}$ $\Psi(k_{*},y^{\delta})$
@@
 SGD $1e^{-5}$ 3529 $7.71e^{-4}$ 17952.29
-Table 3. Comparison of IRSGD and SGD with Rule 1.
+Table 2.
+Quantitative comparison between IRSGD and SGD methods using Rule 1 for various relative noise levels $\delta_{\text{rel}}$.
+The table reports the stopping index $k_{*}$, reconstruction error $E_{k_{*}}^{\delta}$, and $\Psi(k_{*},y^{\delta})$ for each case.
 We observe that, as noise level decreases (moving down the table), both methods generally require more iterations ($k_{*}$) to achieve an acceptable solution.
@@
 The Schlieren optical system produces a light intensity pattern that is proportional to the square of the line integral of pressure along the light path as it passes through the tank (see [10] for more details).
-For each $i$ between 00 and $P-1$, let $\sigma_{i}\in S^{1}$ be a set of recording directions such that $\zeta_{i}=\zeta(\theta_{i})=(\cos\theta_{i},\sin\theta_{i})$.
-The problem is to reconstruct a function $u$ supported on a bounded domain $D:=[-1,1]\times[-1,1]\subset\mathbb{R}^{2}$ from the system of operator equations
-$F_{i}(u)=y_{i}^{\delta},\quad\text{for }i=0,\ldots,P-1,$
+For each $i$ between 0 and P1P-1, let $\sigma_{i}\in S^{1}$ be a set of recording directions such that $\zeta_{i}=\zeta(\theta_{i})=(\cos\theta_{i},\sin\theta_{i})$.
+The objective is to recover a function $u$ supported within the bounded domain $D:=[-1,1]\times[-1,1]\subset\mathbb{R}^{2}$ from the corresponding system of equations
+$F_{i}(u)=y_{i}^{\delta},\quad\text{for }i=0,1,\ldots,P-1,$
 where the Schlieren transform $F_{i}=R_{i}^{2}$, with the operator $R_{i}$ denoting the Radon transform along the direction $\zeta_{i}$, satisfying $R_{i}:C_{0}^{\infty}(D)\to C_{0}^{\infty}(I)$:
-[…]
-From [10], we know that each operator $F_{i}:H_{0}^{1}(D)\to L^{2}(I)$ with $I=[-1,1]$ is continuous and Fréchet differentiable with
-$F_{i}^{\prime}(u)h=2R_{i}(u)R_{i}(h),\quad\forall h\in H_{0}^{1}(D).$
-In addition, the adjoint $F_{i}^{\prime}(u)^{*}:L^{2}(I)\to H_{1}^{0}(D)$ is given by
+$u\mapsto R_{i}(u):=\left(s\mapsto\int u(s\sigma_{i}+r\sigma_{i}^{\perp})\,dr\right).$
+From [10], we know that each operator $F_{i}:H_{0}^{1}(D)\to L^{2}(I)$ with I=[1,1]I=[-1,1] is continuous and Fréchet differentiable with
+$F_{i}^{\prime}(u)h=2R_{i}(u)R_{i}(h),$
+for all $h\in H_{0}^{1}(D).$ Moreover, the adjoint $F_{i}^{\prime}(u)^{*}:L^{2}(I)\to H_{1}^{0}(D)$ is given by
 $F_{i}^{\prime}(u)^{*}g=(I-\Delta)^{-1}\left(2R_{i}^{*}(R_{i}(u)g)\right),\quad\forall g\in L^{2}(I),$
-where $R_{i}^{*}:L^{2}(I)\to L^{2}(D)$ is the adjoint of $R_{i}$ given by $(R_{i}^{*}g)(u)=g(u\cdot\sigma_{i}),$ $I$ refers to the identity operator and $\Delta$ represents the Laplace operator acting on $H_{0}^{1}(D)$.
-The computation of the adjoint operator requires solving an elliptic boundary value problem, which is approximated using a multigrid method implemented through finite difference discretization.
-In this numerical experiment, we set the parameters as follows: $\delta_{\text{rel}}=10^{-2}$, $k_{\text{max}}=750$, $\omega_{k}=\omega=5\times 10^{-3}$, and $\lambda_{k}=\frac{1}{k^{3}}$.
-The true image and initial guess are shown in Figure LABEL:initial_and_true_figures.
-Using the proposed method, we reconstructed the image with a stopping criterion based on Rule 1.
-We evaluated the reconstruction quality by tracking both the relative errors and the progression of $\Psi(k,y^{\delta})$ for different values of $a$: specifically, $a=10^{3}$ and $10^{4}$.
-The results for each case are displayed in Figures LABEL:RE=1000-LABEL:RE=_10000, respectively.
+where $R_{i}^{*}:L^{2}(I)\to L^{2}(D)$ denotes the adjoint of $R_{i}$, defined by $(R_{i}^{*}g)(u)=g(u\cdot\sigma_{i}),$ II refers to the identity operator and $\Delta$ represents the Laplace operator acting on $H_{0}^{1}(D)$.
+Evaluating the adjoint operator entails the solution of an elliptic boundary value problem, for which a finite difference discretization is employed within a multigrid framework to obtain an efficient numerical approximation.
+In this numerical experiment, the parameters were set as follows: $\delta_{\text{rel}}=10^{-2}$, $k_{\text{max}}=1000$, $\omega_{k}=\omega=5\times 10^{-3}$, and $\lambda_{k}=k^{-3}$.
+The true image and the corresponding initial guess are presented in Figure LABEL:initial_and_true_figures.
+The proposed method was applied to reconstruct the image, with the stopping criterion determined by Rule 1 for various values of the parameter $a$, as illustrated in Figure LABEL:Reconstructed.
+To assess the reconstruction performance, we monitored both the relative error and the evolution of $\Psi(k,y^{\delta})$ for selected values of $a$, specifically $a=10^{2},a=10^{3}$ and $a=10^{4}$.
+The corresponding results are displayed in Figures LABEL:RE=1000, LABEL:RE=_10000 and LABEL:RE=_100000.
+Furthermore, the quality of the reconstruction was evaluated using the peak signal-to-noise ratio (PSNR) and the structural similarity index (SSIM) [34], with the obtained values reported in the Table 3, alongside the stopping index and relative error for each case.
+$\delta_{\text{real}}=0.01$
+$a$ in Rule 1 $k_{*}$ $E_{k_{*}}^{\delta}$ $\Psi(k_{*},y^{\delta})$ PSNR SSIM
+$100$ $719$ $1.62e^{-1}$ $1.29e^{+8}$ $28.11$ $0.8251$
+$1000$ $693$ $1.65e^{-1}$ $2.97e^{+8}$ $27.95$ $0.8215$
+$10000$ $698$ $1.70e^{-1}$ $2.04e^{+9}$ $27.70$ $0.8051$
+Table 3.
+Performance comparison for different values of the parameter $a$ in Rule 1 for $\delta_{\text{real}}=0.01$, showing the stopping index $k_{*}$, reconstruction error $E_{k_{*}}^{\delta}$, $\Psi(k_{*},y^{\delta})$, and image quality metrics (PSNR and SSIM).
 5. Conclusion
-In this paper, we present an Iteratively Regularized Stochastic Gradient Descent (IRSGD) method specifically developed to address systems of ill-posed inverse problems, utilizing the Hanke-Raus heuristic rule.
-Notably, the proposed IRSGD method is fully data-driven, as the Hanke-Raus rule operates independently of any prior knowledge of the noise level.
-Our convergence analysis is based on the tangential cone condition and assumes Lipschitz continuity of the operators involved.
-Furthermore, we have conducted a thorough analysis of the convergence of our method in the mean squared norm.
-Additionally, we have established convergence rates under both a-posteriori and heuristic rules by leveraging the source condition, range invariance condition, and stochastic range invariance condition for arbitrary step sizes.
-To demonstrate the practical effectiveness of our approach, we have provided extensive numerical experiments on both linear and nonlinear inverse problems.
-Lastly, we outline several directions for future research.
-First, it is essential to validate our assumptions and theoretical analysis in the context of specific nonlinear inverse problems to further demonstrate the practical applicability of the IRSGD method.
-Second, following the work in [2, 4, 5, 19], the IRSGD approach can be extended to handle nonsmooth nonlinear problems, thereby broadening its range of applications.
-Finally, it is important to investigate adaptive step size rules to provide practical insights and guidelines for real-world implementation.
-The heuristic rule (based on the modified discrepancy principle), which is used in this analysis can be further modified to study the stochastic analysis of regularization techniques presented in [5, 16, 18, 19, 20].
+In this work, we have introduced an Iteratively Regularized Stochastic Gradient Descent (IRSGD) method for solving systems of nonlinear ill-posed inverse problems, incorporating the Hanke–Raus heuristic rule as a fully data-driven stopping criterion.
+A key advantage of this approach is that it operates without any prior knowledge of the noise level, thereby enhancing its applicability in practical scenarios.
+Our convergence analysis relies on the tangential cone condition, noise assumption and assumes Lipschitz continuity of the underlying operators.
+Moreover, we have established convergence results in the mean-squared norm and validated the method through comprehensive numerical experiments on both linear and nonlinear ill-posed problems.
+The results demonstrate that the proposed IRSGD method, augmented with a damping factor, offers improved efficiency and faster convergence compared to standard stochastic gradient approaches.
+Several promising directions for future research emerge from this study.
+First, it would be valuable to relax Assumption 2.2 by introducing milder conditions, following the ideas in [12, 14].
+Second, in line with [1, 3, 4, 26], the framework could be extended to handle non-smooth problems, thereby broadening its applicability.
+Finally, the development of adaptive step-size strategies could provide more practical guidelines for real-world implementation.
+Additionally, the heuristic stopping rule (based on a modified discrepancy principle) could be further refined to enable a deeper stochastic analysis of regularization methods such as those in [4, 22, 25, 26, 27].
 Acknowledgments
-Harshit Bajpai would like to express gratitude to the Science and Engineering Research Board (SERB), India, for providing his PhD fellowship through grant CRG/2022/005491.
-Ankik Kumar Giri also extends his thanks to SERB for supporting his research under the project “Fast and efficient iterative methods for solving nonlinear inverse problems” funded by grant CRG/2022/005491.
+Harshit Bajpai acknowledges the Anusandhan National Research Foundation (ANRF), India, for supporting his Ph.D.
+fellowship under grant CRG/2022/005491.
+Ankik Kumar Giri gratefully acknowledges ANRF for funding his research through the project CRG/2022/005491.
+Data Availability
+Not applicable for this article.